### v2

# Fourier interpolation in dimensions 3 and 4 and real-variable Kloosterman sums

Abstract.
We give a construction of radial Fourier interpolation formulas in dimensions 3 and 4 using Maass–Poincaré type series. As a corollary we obtain explicit formulas for the basis functions of these interpolation formulas in terms of what we call real-variable Kloosterman sums, which were previously introduced by Stoller. We also improve the bounds on the corresponding basis functions $a_{n,d}(x)$, $d=3,4$, for fixed $x$, in terms of the index $n$.
Key words and phrases:
Fourier interpolation, Maass–Poincaré series, Kloosterman sums, nonholomorphic modular forms, modular integrals
2020 Mathematics Subject Classification:
Primary 42B99, 11F03, 11F37, 11L05
1. Introduction
In [RV19, Theorem 1] it was proved that any even Schwartz function $f\colon{\mathbb{R}}\to{\mathbb{C}}$ is uniquely determined by the values $f(\sqrt{n})$, $\widehat{f}(\sqrt{n})$, $n\geq 0$, where
$\widehat{f}(\xi)=\int_{{\mathbb{R}}}f(x)e^{-2\pi i\xi x}dx$
is the Fourier transform of $f$. More precisely, there is a linear interpolation formula
$f(x)=\sum_{n\geq 0}a_{n}(x)f(\sqrt{n})+\sum_{n\geq 0}\widehat{a_{n}}(x)\widehat{f}(\sqrt{n}),$
where $a_{n}(x)$, $\widehat{a_{n}}(x)$ are certain Schwartz functions, explicitly defined as contour integrals of weakly holomorphic modular forms of weight $3/2$. Similar interpolation formulas exist also for radial Schwartz functions in higher dimensions: from [BRS23, Theorem 3.1] it follows that there exist radial Schwartz functions $a_{d,n}$, $\widetilde{a}_{d,n}$ such that
$f(x)=\sum_{n\geq 0}a_{d,n}(x)f(\sqrt{n})+\sum_{n\geq 0}\widetilde{a}_{d,n}(x)\widehat{f}(\sqrt{n}),$ (1.1)
holds for all $f\in\mathcal{S}_{\operatorname{rad}}({\mathbb{R}}^{d})$, and we abuse notation, denoting $g(r)=g(r,0,\dots,0)$ for any radial function $g$ on ${\mathbb{R}}^{d}$ and $r\in{\mathbb{R}}$. Interpolation formulas of similar kind, with $\sqrt{n}$ replaced by $\sqrt{2n}$ and including the values of derivatives of $f$ and $\widehat{f}$, have played an important role in sphere packing and energy minimization problems [Via17], [CKM+17], [CKM+22], where they were used in conjunction with linear programming techniques to prove optimality of the $E_{8}$ lattice and the Leech lattice. For more context and recent developments on the analytic aspects of similar interpolation formulas see [RS22b, RS23, RS22a, Adv23, KNS25].
The coefficients $a_{d,n}(x)$ and $\widetilde{a}_{d,n}(x)$ [^1] The choice of coefficients for which (1.1) holds is not unique. For a fixed $x$, the space of all interpolation formulas of the type (1.1) forms a coset for the $\mathbb{C}$}-vector space of all linear relations between $f(\sqrt{n})$ and $\widehat{f}(\sqrt{n})$, $n\geq 0$, which is finite-dimensional and isomorphic to $M_{d/2}(\Gamma(2),\nu_{\Theta}^{d})$. In [BRS23] a particular choice of $a_{d,n}(x)$, $\widetilde{a}_{d,n}(x)$ is given, uniquely specified by certain vanishing conditions. Henceforth, the notation $a_{d,n}(x)$, $\widetilde{a}_{d,n}(x)$ will refer to this particular choice (for a more precise statement, see Theorem 1.1). from (1.1) can be represented as contour integrals involving weakly holomorphic modular forms of weight 2d/22-d/2. For instance, in dimension 1 we have
$a_{1,0}(x)=\frac{1}{4}\int_{-1}^{1}\Theta^{3}(z)e^{\pi izx^{2}}dz,$
where the integral is taken over a semicircle in the upper half plane, and $\Theta(z)=\sum_{n\in{\mathbb{Z}}}e^{\pi in^{2}z}$ is the usual theta function. Although this construction is explicit, obtaining precise estimates for $a_{d,n}(x)$ as a function of $n$ is not easy. Having such estimates is of interest, for example, since they allow one to extend the class of functions $f$ for which (1.1) holds. In [RV19] it was proved that $a_{1,n}(x),\widetilde{a}_{1,n}(x)$ are bounded by $O(n^{2})$, uniformly in $x\in{\mathbb{R}}$, which was used to show that (1.1) holds if $f(x)$ and $\widehat{f}(x)$ are both $O(|x|^{-13})$ at infinity. In [BRS23] the estimates were improved to
$a_{d,n}(x),\widetilde{a}_{d,n}(x)=\begin{cases}O_{x}(n^{d/4+o(1)}),\qquad d\leq 4,\\ O_{x}(n^{d/2-1}),\;\;\;\qquad d>4,\end{cases}$ (1.2)
which roughly correspond to the classical Hecke bound for the coefficients of cusp forms and estimates for the coefficients of Eisenstein series, respectively.
Getting anything more precise than the bounds in (1.2) appears to be challenging. This motivates the search for different formulas of the basis functions $a_{d,n}(x)$, $\widetilde{a}_{d,n}(x)$. In [Sto21] Stoller proposed an approach based on generalized Poincaré series. His construction works for $d\geq 5$ and produces coefficients $a_{d,n}^{\mathrm{Sto}}(x)$ and $\widetilde{a}_{d,n}^{\mathrm{Sto}}(x)$ that are easier to analyze for $n\to\infty$, but they end up not being Schwartz functions of $x$, and in general they also differ from the basis functions in (1.1).
The main goal of this paper is to give a construction of interpolation formulas of type (1.1) in dimensions $d=3$ and $d=4$ based on (generalized) Maass–Poincaré series, and to explicitly relate these to the formulas from [BRS23]. As one of the applications we obtain an explicit infinite sum representation for the functions $a_{d,n}$ and $\widetilde{a}_{d,n}$, in dimensions $d=3,4$, involving what we call real-variable Kloosterman sums, as well as Hurwitz class numbers (for $d=3$). As another application, we are able to go beyond the bounds (1.2) for $d=3,4$.
As is explained in [RV19, §6], the interpolation formula (1.1) is equivalent to a functional equation for the generating functions of the sequences $a_{d,n}(x)$, $\widetilde{a}_{d,n}(x)$, $n\geq 0$. Our starting point is the following precise version of this claim.
Theorem 1.1 (Corollary of [BRS23, Theorem 3.1]).
Let $\mathbb{H}$} be the complex upper-half plane.
  • (1)
    For any $r\geq 0$ there exists a unique pair of $2$-periodic analytic functions
    $\mathcal{F}_{2}(\,\cdot\,;r),\widetilde{\mathcal{F}_{2}}(\,\cdot\,;r)\colon{\mathbb{H}}\rightarrow{\mathbb{C}}$
    of moderate growth with Fourier expansions of the form
    $\mathcal{F}_{2}(\tau;r)=\sum_{n\geq 1}a_{4,n}(r)e^{\pi in\tau},\quad\widetilde{\mathcal{F}}_{2}(\tau;r)=\sum_{n\geq 1}\widetilde{a}_{4,n}(r)e^{\pi in\tau},\quad\tau\in{\mathbb{H}},$
    such that
    $\mathcal{F}_{2}(\tau;r)+(\tau/i)^{-2}\mathcal{F}_{2}(-1/\tau;r)=e^{\pi ir^{2}\tau}.$
  • (2)
    For any $r\geq 0$ there exists a unique pair of $2$-periodic analytic function
    $\mathcal{F}_{3/2}(\,\cdot\,;r),\widetilde{\mathcal{F}}_{3/2}(\,\cdot\,;r):{\mathbb{H}}\rightarrow{\mathbb{C}}$
    of moderate growth with Fourier expansions of the form
    $\mathcal{F}_{3/2}(\tau;r)=\sum_{n\geq 0}a_{3,n}(r)e^{\pi in\tau},\quad\widetilde{\mathcal{F}}_{3/2}(\tau;r)=\sum_{n\geq 0}\widetilde{a}_{3,n}(r)e^{\pi in\tau},\quad\tau\in{\mathbb{H}},$
    and $a_{3,0}(r)=\widetilde{a}_{3,0}(r)$, such that
    $\mathcal{F}_{3/2}(\tau;r)+(\tau/i)^{-3/2}\widetilde{\mathcal{F}}_{3/2}(-1/\tau;r)=e^{\pi ir^{2}\tau}.$
Here the condition of moderate growth simply means polynomial growth (in $n$) of the coefficients $a_{4,n}(r),\widetilde{a}_{4,n}(r)$ and $a_{3,n}(r),\widetilde{a}_{3,n}(r)$. The functions appearing in (1.1) for $d=3,4$ are precisely the coefficients $a_{d,n}(r),\widetilde{a}_{d,n}(r)$ that are uniquely determined by the conditions of Theorem 1.1 (by convention we set $a_{4,0}=\widetilde{a}_{4,0}=0$ and $a_{3,0}=\widetilde{a}_{3,0}$).
In this paper we will construct the analytic functions $\mathcal{F}_{2}$ and $\mathcal{F}_{3/2}$ from Theorem 1.1 via Maass–Poincaré-type series. Specifically, we will construct a Maass–Poincaré-type series with a spectral parameter $s\in{\mathbb{C}}$, weight $k\in\{\frac{3}{2},2\}$ and a real parameter $r\in{\mathbb{R}}$, and prove that its Fourier expansion converges for $s=\frac{k}{2}$. Moreover, we will show that these series can be adjusted to satisfy the conditions of Theorem 1.1. This will give us an explicit identity for $a_{d,n}$, $d=3,4$, in terms of sums of real-variable Kloosterman sums. We now formulate these results in more detail.
In what follows we denote by $J_{\alpha}(x)$ the $J$-Bessel function [DLMF, (10.2.2)]. In dimension $d=4$, we have the following result.
Theorem 1.2.
Let two sequences of entire functions $b_{4,n},\widetilde{b}_{4,n}$, $n\geq 1$, be given by
$b_{4,n}(r)$ $=B_{4,n}(r)+\frac{8\sin(\pi r^{2})}{\pi r^{2}}\left(\sigma_{1}(\tfrac{n}{2})-(-1)^{n}\sigma_{1}(n)\right),$
$\widetilde{b}_{4,n}(r)$ $=\widetilde{B}_{4,n}(r)+\frac{8\sin(\pi r^{2})}{\pi r^{2}}\left(\sigma_{1}(\tfrac{n}{2})-(-1)^{n}\sigma_{1}(n)\right),$
where $\sigma_{1}(n)=\sum_{d|n}d$ is the sum of divisors of $n$, $\sigma_{1}(x)=0$ if $x\notin{\mathbb{Z}}$. Here $B_{4,n}(r)$ and $\widetilde{B}_{4,n}(r)$ are defined as
$B_{4,n}(r)=\pi\left(\frac{n}{r^{2}}\right)^{\frac{1}{2}}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right),$
[…]
where the real-variable Kloosterman sums $K$ and $\widetilde{K}$ are defined at (3.1) and (3.2), respectively. Then we have the following.
  • (1)
    For every $f\in{\mathcal{S}}_{\operatorname{rad}}({\mathbb{R}}^{4})$ and every $x\in{\mathbb{R}}^{4}$, we have
    $f(x)=\sum_{n=1}^{\infty}f(\sqrt{n})b_{4,n}(|x|)+\sum_{n=1}^{\infty}\widehat{f}(\sqrt{n})\widetilde{b}_{4,n}(|x|).$
  • (2)
    We have the following estimates for $B_{4,n}(r)$ and $\widetilde{B}_{4,n}(r)$:
    $|B_{4,n}(r)|,|\widetilde{B}_{4,n}(r)|\ll_{\varepsilon}\left\{\begin{array}[]{ll}n^{1+\varepsilon}&\text{if }r^{2}\leq 1/n,\\ (|r|^{-1}+|r|^{-\frac{1}{2}+\varepsilon})n^{\frac{3}{4}+\varepsilon}&\text{if }r^{2}\geq 1/n.\end{array}\right$.
  • (3)
    Let $\nu_{2}(n)$ be the 2-adic valuation of $n$. The values $B_{4,n}(0)$ and $\widetilde{B}_{4,n}(0)$ are given by
    $B_{4,n}(0)=8\sigma_{1}(n)\cdot\left\{\begin{array}[]{ll}\displaystyle\frac{2^{\nu_{2}(n)}-3}{2^{\nu_{2}(n)+1}-1},&2|n,\vskip 2.00749pt\\ 0,&2\nmid n,\end{array}\right.\quad\widetilde{B}_{4,n}(0)=8\sigma_{1}(n)\cdot\frac{2^{\nu_{2}(n)}}{2^{\nu_{2}(n)+1}-1}.$
Remark.
Note that the coefficients $A_{4}(n)\vcentcolon=\sigma_{1}(\frac{n}{2})-(-1)^{n}\sigma_{1}(n)$ satisfy
$2\sum_{n=1}^{\infty}A_{4}(n)q^{n}=\sum_{n\in{\mathbb{Z}}}n^{2}q^{n^{2}}\Big/\sum_{n\in{\mathbb{Z}}}q^{n^{2}}.$
In dimension $d=3$, we have the following interpolation function.
Theorem 1.3.
Let two sequences of entire functions $b_{3,n},\widetilde{b}_{3,n}$, $n\geq 0$, be given by
$b_{3,0}(r)$ $=\widetilde{b}_{3,0}(r)=\frac{\sin(\pi r^{2})}{2r\sinh(\pi r)},$
$b_{3,n}(r)$ $=B_{3,n}(r)-\frac{\sin(\pi r^{2})}{r\sinh(\pi r)}\left(8H(n)+\frac{r_{3}(n)}{6}\right),\quad n\geq 1,$
$\widetilde{b}_{3,n}(r)$ $=\widetilde{B}_{3,n}(r)-\frac{\sin(\pi r^{2})}{r\sinh(\pi r)}\left(8H(n)+\frac{r_{3}(n)}{6}\right),\quad n\geq 1,$
where $H(n)$ is the Hurwitz class number and $r_{3}(n)$ is the number of ways to write $n$ as a sum of three squares (also the coefficient of $e^{\pi in\tau}$ in $\Theta(\tau)^{3}$ (2.6)). Here $B_{3,n}(r)$ and $\widetilde{B}_{3,n}(r)$ are defined as
$B_{3,n}(r)$ $=\frac{{\mathbf{e}}(\tfrac{1}{8})}{|r|}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{\sqrt{c}}\sin\left(\frac{2\pi|r|\sqrt{n}}{c}\right),$
[…]
where the real-variable Kloosterman sums $K$ and $\widetilde{K}$ are defined at (3.1) and (3.2), respectively. Then
  • (1)
    For every $f\in{\mathcal{S}}_{\operatorname{rad}}({\mathbb{R}}^{3})$ and every $x\in{\mathbb{R}}^{3}$ we have
    $f(x)=\sum_{n=0}^{\infty}f(\sqrt{n})b_{3,n}(|x|)+\sum_{n=0}^{\infty}\widehat{f}(\sqrt{n})\widetilde{b}_{3,n}(|x|).$
  • (2)
    We have the following estimates for $B_{3,n}(r)$ and $\widetilde{B}_{3,n}(r)$:
    $|B_{3,n}(r)|,|\widetilde{B}_{3,n}(r)|\ll_{\varepsilon}\left\{\begin{array}[]{ll}\max\big(n^{\frac{437}{588}},n^{\frac{11+\kappa}{16}}\big)n^{\varepsilon}&\text{if }r^{2}\leq 1/n,\vskip 2.00749pt\\ (|r|^{-\frac{1}{2}}+|r|^{\varepsilon})n^{\frac{1}{2}+\varepsilon}&\text{if }r^{2}\geq 1/n.\end{array}\right$.
    Here $\kappa\in[0,1]$ is chosen so that $2^{\nu_{2}(n)}\ll n^{\kappa}$. Note that $\frac{437}{588}\approx 0.743\ldots<0.75$.
  • (3)
    For $n\geq 1$, we have
    $\begin{array}[]{rll}&B_{3,n}(0)=8H(n)-r_{3}(n)/3,&\widetilde{B}_{3,n}(0)=8H(n)+2r_{3}(n)/3,\\ &\>\,b_{3,n}(0)=-{r_{3}(n)}/2,&\;\widetilde{b}_{3,n}(0)={r_{3}(n)}/2.\end{array}$
Remark.
(1) Note that $B_{3,n}(0),\widetilde{B}_{3,n}(0)\ll_{\varepsilon}n^{\frac{1}{2}+\varepsilon}$, while the exponent $\frac{437}{588}$, appearing in the estimate for general $r>0$, is equal to $\frac{3}{4}-\frac{1}{147}$.
(2) In §9 we prove [^2] It is likely that (1.3) is well-known, but we could not find it in the literature the following identity between Zagier’s weight $3/2$ non-holomorphic modular form $\mathcal{H}(\tau)$ (see (9.1) for definition) and $\Theta(\tau)$:
$\mathcal{H}(\tau)+(2\tau/i)^{-3/2}\mathcal{H}(-1/4\tau)=-\Theta(2\tau)^{3}/24.$ (1.3)
The quantity $A_{3}(n)\vcentcolon=8H(n)+\frac{1}{6}r_{3}(n)$ comes from the Fourier coefficients of the function
$\widetilde{\mathcal{H}}(\tau)\vcentcolon=8\mathcal{H}\left(\frac{\tau}{2}\right)+\frac{\Theta(\tau)^{3}}{6}=\sum_{n=1}^{\infty}A_{3}(n)e^{\pi in\tau}+\text{non-holomorphic terms}$
that satisfies $\widetilde{\mathcal{H}}(\tau)+(\tau/i)^{-3/2}\widetilde{\mathcal{H}}(-1/\tau)=0$.
1.1. Relation with Theorem 1.1
For $d\in\{3,4\}$, the formulas for the functions $b_{d,n}(r)$ and $\widetilde{b}_{d,n}(r)$ will follow from a construction of the generating series
$G_{\frac{d}{2}}(\tau;r)=\sum_{n=0}^{\infty}b_{d,n}(r)e^{\pi in\tau}\quad\text{and}\quad\widetilde{G}_{\frac{d}{2}}(\tau;r)=\sum_{n=0}^{\infty}\widetilde{b}_{d,n}(r)e^{\pi in\tau}.$ (1.4)
In the proofs of Theorem 1.2 and Theorem 1.3, we will show that these series satisfy the functional equation
$G_{\frac{d}{2}}(\tau;r)+(-i\tau)^{-d/2}\widetilde{G}_{\frac{d}{2}}(-\tfrac{1}{\tau};r)=e^{\pi ir^{2}\tau},$ (1.5)
which implies (and by [RV19, §6] is equivalent to) the corresponding Fourier interpolation formula. For $d=4$ we have $b_{4,0}=\widetilde{b}_{4,0}=0$, and for $d=3$ we have $b_{3,0}=\widetilde{b}_{3,0}$. Therefore, by the uniqueness part of Theorem 1.1, we get the following result.
Corollary 1.4.
Let $d\in\{3,4\}$. Then for $\mathcal{F}_{\frac{d}{2}}(\tau;r)$ and $\widetilde{\mathcal{F}}_{\frac{d}{2}}(\tau;r)$ defined in Theorem 1.1 and $\mathcal{G}_{\frac{d}{2}}^{\varepsilon}(\tau;r)$ defined at (1.4), we have
$\mathcal{G}_{\frac{d}{2}}(\tau;r)=\mathcal{F}_{\frac{d}{2}}(\tau;r),\qquad\widetilde{\mathcal{G}}_{\frac{d}{2}}(\tau;r)=\widetilde{\mathcal{F}}_{\frac{d}{2}}(\tau;r).$
In particular, we have $a_{3,n}=b_{3,n}$, $\widetilde{a}_{3,n}=\widetilde{b}_{3,n}$, $a_{4,n}=b_{4,n}$, and $\widetilde{a}_{4,n}=\widetilde{b}_{4,n}$ for all $n\geq 0$.
In [RV19, §7], the first author together with Viazovska constructed an interpolation basis for one-dimensional odd Schwartz functions.
Theorem 1.5 (Corollary of [RV19, Prop. 3, Thm. 7]).
There exists a collection of odd Schwartz functions $d_{n}^{\varepsilon}$, $n\geq 0$, $\varepsilon\in\{\pm\}$, satisfying
$\widehat{d_{n}^{\varepsilon}}(x)=\varepsilon(-i)d_{n}^{\varepsilon}(x)\quad\text{and}\quad d_{n}^{\varepsilon}(\sqrt{m})=\delta_{n,m}\sqrt{m},\quad m\geq 0,$
together with ${d_{n}^{+}}^{\prime}(0)=\delta_{n,0}$ and ${d_{n}^{-}}^{\prime}(0)=0$, $n\geq 0$. The Schwartz functions $d_{n}^{\varepsilon}$ are uniquely determined by the above properties.
From Theorem 1.3, Corollary 1.4, and [RV19, Eq. (39)] we get that
$\begin{split}d_{n}^{+}(x)&=xB_{3,n}(|x|)+x\widetilde{B}_{3,n}(|x|)-\frac{2\sin(\pi x^{2})}{\sinh(\pi x)}\left(8H(n)+\frac{r_{3}(n)}{6}\right),\\ d_{n}^{-}(x)&=xB_{3,n}(|x|)-x\widetilde{B}_{3,n}(|x|).\end{split}$ (1.6)
Remark.
The above equation for $n=0$ implies (using that $H(0)=-\tfrac{1}{12}$) that $d_{0}^{+}(x)=\frac{\sin(\pi x^{2})}{\sinh(\pi x)}$, as noted in [RV19, §8].
When rewritten in terms of real-variable Kloosterman sums, we get the following corollary.
Corollary 1.6.
For $n\geq 1$, let us define
$K_{n}(r)$ $\vcentcolon=\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{\sqrt{c}}\sin\left(\frac{2\pi r\sqrt{n}}{c}\right)={\mathbf{e}}(-\tfrac{1}{8})rB_{3,n}(|r|),$ (1.7)
$\widetilde{K}_{n}(r)$ $\vcentcolon=\sum_{2\nmid{\widetilde{c}}>0}\frac{\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})}{\sqrt{{\widetilde{c}}}}\sin\left(\frac{2\pi r\sqrt{n}}{{\widetilde{c}}}\right)={\mathbf{e}}(\tfrac{3}{8})r\widetilde{B}_{3,n}(|r|).$ (1.8)
Then
  • (1)
    $K_{n}(r),\widetilde{K}_{n}(r)$ are odd Schwartz functions,
  • (2)
    $\widehat{K_{n}}(r)=i\,\widetilde{K}_{n}(r)$, $\widehat{\widetilde{K}_{n}}(r)=i\,K_{n}(r),$ and
  • (3)
    for $m\geq 1$,
    $K_{n}(\sqrt{m})={\mathbf{e}}(-\tfrac{1}{8})\delta_{m,n}\,\sqrt{m},\quad\widetilde{K}_{n}(\sqrt{m})=0.$ (1.9)
Remark.
The property (1.9) provides another proof of the following identity on half-integral weight Kloosterman sums on $\Gamma_{0}(4)$ (for definitions see (2.5) and (2.14)):
$2\pi{\mathbf{e}}(-\tfrac{1}{8})\left(\frac{n}{m}\right)^{\frac{1}{4}}\sum_{4|c>0}\frac{S(-m,-n,c,\nu_{\theta})}{c}J_{\frac{1}{2}}\left(\frac{4\pi\sqrt{mn}}{c}\right)=\delta_{m,n},\quad\text{for }m,n\geq 1.$ (1.10)
We also significantly improve the growth estimates for $d_{n}^{\varepsilon}(r)$ in [RV19, Theorem 6], and improve the bound (1.2) from [BRS23].
Corollary 1.7.
For the function $d_{n}^{\varepsilon}(x)$ in Theorem 1.5 where $\varepsilon\in\{\pm\}$, we have
$d_{n}^{\varepsilon}(x)=-(1+\varepsilon)\frac{\sin(\pi x^{2})}{\sinh(\pi x)}\left(8H(n)+\frac{r_{3}(n)}{6}\right)+B_{n}^{\varepsilon}(x),$
where $B_{n}^{\varepsilon}(x)=xB_{3,n}(|x|)+\varepsilon x\widetilde{B}_{3,n}(|x|)$ is an odd Schwartz function of $x$, given by a sum of real-variable Kloosterman sums, and can be estimated by
$B_{n}^{\varepsilon}(x)\ll_{\delta}\left\{\begin{array}[]{ll}|x|\max\big(n^{\frac{3}{4}-\frac{1}{147}},n^{\frac{11+\kappa}{16}}\big)n^{\delta}&\text{ if }x^{2}\leq 1/n,\vskip 2.00749pt\\ (|x|^{\frac{1}{2}}+|x|^{1+\delta})n^{\frac{1}{2}+\delta}&\text{ if }x^{2}\geq 1/n,\end{array}\right$.
for any $\delta>0$. As Theorem 1.3, $\kappa\in[0,1]$ is chosen so that $2^{\nu_{2}(n)}\ll n^{\kappa}$.
The paper is organized as follows. In Section 2 we setup notation and recall some preliminary results on Kloosterman sums and Maass forms. In Section 3 we construct (following the ideas of [Sto21]) modular integrals via Maass–Poincaré series by a special choice of coset representatives modulo parabolic subgroup for $\Gamma(2)$ and compute their Fourier expansions. Section 4 reviews properties of weight $2$ Kloosterman sums, which we then use to prove properties of the corresponding weight $2$ real-variable Kloosterman sums in Section 5. We prove Theorem 1.2 in Section 6. Sections 7, 8 and 9 treat, respectively, the properties of weight $3/2$ Kloosterman sums, the properties of weight $3/2$ real-variable Kloosterman sums, and the proof of Theorem 1.3. We conclude with a brief discussion of open questions in Section 10.
Acknowledgement
The authors acknowledge funding by the European Union (ERC, FourIntExP, 101078782). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or European Research Council (ERC). Neither the European Union nor ERC can be held responsible for them.
Conflict of interest statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data availability statement
No data were generated or analyzed in this study.
2. Preliminaries and notation
We use $\mathbb{Z}$$_{+}$ and $\mathbb{R}$$_{+}$ to denote the sets of positive integers and positive real numbers, respectively. Let $I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$ denote the $2\times 2$ identity matrix, $T=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ and $S=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. For $n\in{\mathbb{Z}}_{+}$, let $\varphi(n)$ denote Euler’s totient function and define the divisor sum function by
$\sigma_{s}(n)\vcentcolon=\sum_{d|n}d^{s}.$
As a special case, $\sigma_{0}(n)$ is the number of divisors of $n$. Let $\operatorname{sgn}n\in\{-1,0,1\}$ denote the sign function for $n\in{\mathbb{Z}}$. For any prime $p$, let $\nu_{p}(n)$ denote the $p$-adic valuation of $n$, i.e., $p^{\nu_{p}(n)}\|n$, where $p^{t}\|b$ means $p^{t}|b$ and $p^{t+1}\nmid b$. We also denote
${\mathbf{e}}(\tau)\vcentcolon=e^{2\pi i\tau}$
and let $\Gamma(a,\tau)$ be the incomplete Gamma function [DLMF, (8.2.2)].
Let $\zeta(s)$ be the Riemann zeta function. If $\chi$ is a Dirichlet character, let $L(s,\chi)$ be the Dirichlet $L$-function for $s\in{\mathbb{C}}$ where
$L(s,\chi)=\sum_{n=1}^{\infty}\frac{\chi(n)}{n^{s}}\qquad\text{for }\operatorname{Re}(s)>1.$
For $m,n\in{\mathbb{Z}}$, let $(\frac{m}{n})$ be the Kronecker symbol where $(\frac{m}{-1})=1$ if $m\geq 0$ and $(\frac{m}{-1})=-1$ if $m<0$. In particular, for any $m\in{\mathbb{Z}}$, $(\frac{-4m}{\cdot})$ is a Dirichlet character modulo $|4m|$.
We recall the following standard congruence subgroups of $\operatorname{SL}_{2}({\mathbb{Z}})$. For $N\in{\mathbb{Z}}_{+}$, define
$\Gamma_{0}(N)=\{\begin{pmatrix}*&*\\ c&*\end{pmatrix}\in\operatorname{SL}_{2}({\mathbb{Z}}):c\equiv 0\ (\mathrm{mod}\ N)\},\quad\Gamma(N)=\{\gamma\in\operatorname{SL}_{2}({\mathbb{Z}}):\gamma\equiv I\ (\mathrm{mod}\ N)\}.$
Whether $\Gamma(\cdot)$ refers the congruence subgroup (we only use $\Gamma(2)$) or the Gamma function (like $\Gamma(s\pm\frac{k}{2})$) shall be clear among the context.
Let $\Gamma_{\Theta}$ denote another congruence subgroup of $\operatorname{SL}_{2}({\mathbb{Z}})$:
$\Gamma_{\Theta}=\{\gamma\in\operatorname{SL}_{2}({\mathbb{Z}}):\gamma\equiv I\text{ or }S\ (\mathrm{mod}\ 2)\}.$ (2.1)
It is known that $\Gamma_{\Theta}$ is freely generated by $T^{2}$ and $S$ subject to $S^{2}=-1$.
For $\tau\in{\mathbb{C}}^{\times}$, we define the argument $\arg(\tau)$ to lie in $(-\pi,\pi]$. Let ${\mathbb{H}}\vcentcolon=\{\tau\in{\mathbb{C}}:\operatorname{Im}\tau>0\}$ be the complex upper half-plane. For any $\gamma=\begin{pmatrix}*&*\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}({\mathbb{R}})$ and $\tau\in{\mathbb{H}}$, we define the automorphic factor
$j(\gamma,\tau)\vcentcolon=c\tau+d.$
For $k\in\frac{1}{2}{\mathbb{Z}}$ and $\tau\in{\mathbb{C}}$, we define
$\tau^{k}=|\tau|^{k}\exp(ik\arg(\tau))$
and the weight $k$ slash operator
$(f|_{k}\gamma)(\tau)\vcentcolon=j(\gamma,\tau)^{-k}f(\gamma\tau).$ (2.2)
Definition 2.1.
We say that $\nu:\Gamma\to{\mathbb{C}}^{\times}$ is a multiplier system of weight $k\in\frac{1}{2}{\mathbb{Z}}$ if
  1. (i)
    $|\nu|=1$,
  2. (ii)
    $\nu(-I)=e^{-\pi ik}$, and
  3. (iii)
    $\nu(\gamma_{1}\gamma_{2})=w_{k}(\gamma_{1},\gamma_{2})\nu(\gamma_{1})\nu(\gamma_{2})$ for all $\gamma_{1},\gamma_{2}\in\Gamma$, where
    $w_{k}(\gamma_{1},\gamma_{2})\vcentcolon=j(\gamma_{2},\tau)^{k}j(\gamma_{1},\gamma_{2}\tau)^{k}j(\gamma_{1}\gamma_{2},\tau)^{-k}.$
If $\nu$ is a multiplier system of weight $k$, then it is also a multiplier system of weight $k^{\prime}$ for any $k^{\prime}\equiv k\pmod{2}$, and its conjugate $\overline{\nu}$ is a multiplier system of weight k-k. One can also easily check that
$\nu(\gamma)\nu(\gamma^{-1})=1\quad\text{and}\quad\nu(\gamma\begin{pmatrix}1&bt\\ 0&1\end{pmatrix})=\nu(\gamma)\nu(\begin{pmatrix}1&b\\ 0&1\end{pmatrix})^{t}\quad\text{for }b,t\in{\mathbb{Z}}.$ (2.3)
From the definition it follows that if $\nu$ is a multiplier system of weight $\frac{1}{2}$, then
$\nu(-\gamma)=i\,\nu(\gamma)\quad\text{if }\gamma=\begin{pmatrix}*&*\\ c&*\end{pmatrix}\text{ and }c>0.$ (2.4)
For any congruence subgroup $\Gamma$ and any cusp $\mathfrak{a}\in\mathbb{P}^{1}({\mathbb{Q}})/\Gamma$, let $\Gamma_{\mathfrak{a}}$ be the stabilizer of $\mathfrak{a}$ in $\Gamma$. For example, $\Gamma_{\infty}=\{\pm\begin{pmatrix}1&b\\ 0&1\end{pmatrix}:b\in{\mathbb{Z}}\}\cap\Gamma$. Let $\sigma_{\mathfrak{a}}\in\operatorname{SL}_{2}({\mathbb{R}})$ denote a scaling matrix satisfying $\sigma_{\mathfrak{a}}\infty=\mathfrak{a}$ and $\sigma_{\mathfrak{a}}^{-1}\Gamma_{\mathfrak{a}}\sigma_{\mathfrak{a}}=\Gamma_{\infty}$.
Let $\overline{\Gamma}\vcentcolon=\Gamma/\{\pm I\}$ for any congruence subgroup $\Gamma$. For an element $\gamma\in\operatorname{SL}_{2}({\mathbb{Z}})$, we let $\gamma]$ denote its class modulo $\{\pm I\}$. In this subsection we focus on $\Gamma(2)=\{\gamma\in\operatorname{SL}_{2}({\mathbb{Z}}):\gamma\equiv I\ (\mathrm{mod}\ 2)\}$ as a congruence subgroup of $\Gamma_{\Theta}$ (2.1). We know that $\overline{\Gamma(2)}$ is freely generated by $A=[T^{2}]$ and $B=[ST^{2}S]$, and $\Gamma_{\Theta}=\Gamma(2)\sqcup\Gamma(2)S$, where $\sqcup$ means the disjoint union.
2.1. Kloosterman sums
For any congruence subgroup $\Gamma$, we write the stabilizer of the cusp at $\infty$ as $\Gamma_{\infty}=\{\pm\begin{pmatrix}1&b\\ 0&1\end{pmatrix}:b\in w{\mathbb{Z}}\}$ for some $w\in{\mathbb{Z}}_{+}$ which denotes the width of the cusp. For $c>0$, we define the Kloosterman sums with multiplier system $\nu$ on $\Gamma$ as
$S(m,n,c,\nu)=\sum_{\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{\infty}\setminus\Gamma/\Gamma_{\infty}}\nu(\gamma)^{-1}{\mathbf{e}}\left(\frac{ma+nd}{wc}\right).$ (2.5)
Let $S(m,n,c)\vcentcolon=S(m,n,c,{\text{id}}_{\operatorname{SL}_{2}({\mathbb{Z}})})$ denote the standard Kloosterman sums as [ST09, (1)].
We define the theta function as
$\Theta(\tau)=\sum_{n\in{\mathbb{Z}}}e^{\pi in^{2}\tau}.$ (2.6)
It satisfies
$\Theta(\gamma\tau)=\nu_{\Theta}(\gamma)(c\tau+d)^{\frac{1}{2}}\Theta(\tau),\quad\gamma\in\Gamma_{\Theta},$ (2.7)
where $\nu_{\Theta}$ is a weight $\frac{1}{2}$ multiplier system on $\Gamma_{\Theta}$. For $c>0$, we have
$\nu_{\Theta}\begin{pmatrix}a&b\\ c&d\end{pmatrix}=\left\{\begin{array}[]{ll}\varepsilon_{d}^{-1}\left(\frac{2c}{d}\right),&c\equiv 0\ (\mathrm{mod}\ 2),\\ {\mathbf{e}}(-\frac{1}{8})\varepsilon_{c}\left(\frac{2d}{c}\right),&c\equiv 1\ (\mathrm{mod}\ 2),\end{array}\right.\quad\text{where }\varepsilon_{d}=\left\{\begin{array}[]{ll}1,&d\equiv 1\ (\mathrm{mod}\ 4),\\ i,&d\equiv 3\ (\mathrm{mod}\ 4),\\ 0,&\text{otherwise.}\end{array}\right$. (2.8)
Equivalently, by [Mum07, Theorem 7.1], for $c>0$ we have
$\nu_{\Theta}\begin{pmatrix}a&b\\ c&d\end{pmatrix}=\left\{\begin{array}[]{ll}i^{\frac{d-1}{2}}(\frac{c}{|d|}),&c\equiv 0\ (\mathrm{mod}\ 2),\\ {\mathbf{e}}(-\frac{c}{8})(\frac{d}{c}),&c\equiv 1\ (\mathrm{mod}\ 2).\end{array}\right$. (2.9)
Specifically, we have $\nu_{\Theta}\begin{pmatrix}*&*\\ 0&*\end{pmatrix}=1$, $\nu_{\Theta}(S)={\mathbf{e}}(-\frac{1}{8})$, and
$\nu_{\Theta}(\gamma S)=(\operatorname{sgn}d)\cdot{\mathbf{e}}(-\tfrac{1}{8})\,\nu_{\Theta}(\gamma)\quad\text{for }c>0.$ (2.10)
For $c<0$ we recall (2.4).
Another normalization for Jacobi’s theta function is
$\theta(\tau)\vcentcolon=\sum_{n\in{\mathbb{Z}}}{\mathbf{e}}(n^{2}\tau)=\Theta(2\tau).$ (2.11)
It is a modular form of weight $\frac{1}{2}$ on $\Gamma_{0}(4)$ with a multiplier system $\nu_{\theta}$ that is given by
$\theta(\gamma\tau)=\nu_{\theta}(\gamma)(c\tau+d)^{\frac{1}{2}}\theta(\tau),\quad\nu_{\theta}\begin{pmatrix}a&b\\ c&d\end{pmatrix}=\varepsilon_{d}^{-1}(\tfrac{c}{d}),\quad\text{for }\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}(4).$ (2.12)
Since the stabilizer in $\Gamma_{\Theta}$ of the cusp at $\infty$ is $(\Gamma_{\Theta})_{\infty}=\{\pm T^{2n}:n\in{\mathbb{Z}}\}$, the width of the cusp at $\infty$ is 2. For $k\in\frac{1}{2}{\mathbb{Z}}$ and $c\in{\mathbb{Z}}_{+}$, we can explicitly write down the weight $k$ Kloosterman sums defined on $\Gamma_{\Theta}$ with multiplier system $\nu_{\Theta}^{2k}$, where we shorten a(modc)a\ ($\mathrm${mod}\ c) to $a(c)$ for simplicity:
[…]
For $m,n\in{\mathbb{Z}}$ and $c\in 2{\mathbb{Z}}_{+}$, we have $4|2c$ and the following relations are clear by definition:
$S(m,n,c,\nu_{\Theta}^{2k})=S(m,n,2c,\nu_{\theta}^{2k}),\quad\text{in particular, \ }S(m,n,c,\nu_{\Theta}^{4})=S(m,n,2c).$ (2.14)
We also need the conjugation property of Kloosterman sums. For any multiplier system $\nu$ on a congruence subgroup $\Gamma$, if $\nu(\gamma)=1$ for all $\gamma\in\Gamma_{\infty}$, then
$\overline{S(m,n,c,\nu)}=S(-m,-n,c,\overline{\nu}).$ (2.15)
Specifically, we have
$\overline{S(m,n,c,\nu_{\Theta}^{2k})}=S(-m,-n,c,\nu_{\Theta}^{-2k})\quad\text{and}\quad\overline{S(m,n,c,\nu_{\theta}^{2k})}=S(-m,-n,c,\nu_{\theta}^{-2k}).$ (2.16)
We also have
$S(m,n,c,\nu_{\Theta}^{2k})=S(n,m,c,\nu_{\Theta}^{2k})\quad\text{for }2|c,$ (2.17)
because $ad\equiv 1\ (\mathrm{mod}\ 2c)$, for even $c$, implies $\varepsilon_{a}=\varepsilon_{d}$ and $(\frac{2c}{a})=(\frac{2c}{d})$.
2.2. Other relations between Kloosterman sums
In this subsection we prove the following proposition.
Proposition 2.2.
For $m,n\in{\mathbb{Z}}$ and $c=2\widetilde{c}$ where $\widetilde{c}\in{\mathbb{Z}}_{+}$ is odd, we have:
[…]
$S(2m,2n,c,\nu_{\Theta}^{4})=2(-1)^{m+n+1}S(m,n,\widetilde{c},\nu_{\Theta}^{4}).$ (2.20)
These relations will help us in estimating sums of $S(m,n,\widetilde{c},\nu_{\Theta}^{2k})$ for $k=\frac{3}{2}$ and $2$.
The proof goes along similar lines as Biró’s work in [Bir00, Appendix §A.3], which is about Kloosterman sums $S(m,n,c,\nu_{\theta})$ on $\Gamma_{0}(4)$ with cusp pairs $(\infty,\infty)$, $(\infty,0)$ and $(\infty,\frac{1}{2})$. One can show that the Kloosterman sum $S(m,n,\widetilde{c},\nu_{\Theta}^{2k})$ for odd $\widetilde{c}$ is essentially the Kloosterman sum with cusp pair $(\infty,0)$ on $\Gamma(2)$. The relation between $\Gamma_{0}(4)$ and $\Gamma(2)$ allows us to translate Biró’s computations to our setting.
The following lemma is direct and we omit the proof.
Lemma 2.3.
For odd $\widetilde{c}\in{\mathbb{Z}}_{+}$, the following integer sets are the same modulo $2\widetilde{c}$:
  • (1)
    $\{d:\ \ 1\leq d\leq 2\widetilde{c},\ 2|d,\ (d,\widetilde{c})=1\}$;
  • (2)
    $\{2d:\ 1\leq d\leq 4\widetilde{c},\ (d,2\widetilde{c})=1,\ d\equiv 1\ (\mathrm{mod}\ 4)\}$;
  • (3)
    $\{\beta d:\ 1\leq d\leq 2\widetilde{c},\ (d,2\widetilde{c})=1\}$, for any even $\beta$ with $(\beta,\widetilde{c})=1$.
Now we suppose $\widetilde{c}\in{\mathbb{Z}}_{+}$ is odd and $c=2\widetilde{c}$. Let $\overline{x_{n}}$ denote the inverse of $x$ modulo $n$. For every $d$ modulo $2c$ such that $(d,2c)=1$, we use the pairing $d\leftrightarrow d+c$, the fact that $d\equiv(-1)^{n}\ (\mathrm{mod}\ 4)$ if and only if $d+c\equiv(-1)^{n+1}\ (\mathrm{mod}\ 4)$, and the fact that $\overline{(d+c)_{2c}}\equiv\overline{d_{2c}}+c$ modulo $2c$. Then we have
$\begin{split}S(m,4n,c,\nu_{\Theta}^{2k})&=\sum_{\begin{subarray}{c}d\ (\mathrm{mod}\ 2c)\\ ad\equiv 1(2c)\end{subarray}}\varepsilon_{d}^{2k}\left(\frac{2c}{d}\right)^{2k}{\mathbf{e}}\left(\frac{ma+4nd}{2c}\right)\\ &=\left(1+i^{2k}(-1)^{\frac{\widetilde{c}-1}{2}\cdot 2k+m}\right)\sum_{\begin{subarray}{c}d\ (\mathrm{mod}\ 2c)\\ ad\equiv 1(2c)\\ d\equiv 1\ (\mathrm{mod}\ 4)\end{subarray}}\left(\frac{d}{c/2}\right)^{2k}{\mathbf{e}}\left(\frac{ma+4nd}{2c}\right).\end{split}$ (2.21)
Since $4\cdot\overline{4_{\widetilde{c}}}=1+\beta\widetilde{c}$ implies $\beta+\widetilde{c}\equiv 0\ (\mathrm{mod}\ 4)$, for $a\equiv 1\ (\mathrm{mod}\ 4)$ we have
${\mathbf{e}}\left(\frac{ma}{2c}\right)={\mathbf{e}}\left(\frac{ma\cdot\overline{4_{\widetilde{c}}}}{\widetilde{c}}\right){\mathbf{e}}\left(\frac{m\widetilde{c}}{4}\right).$ (2.22)
We obtain that
[…]
Taking the following facts into account:
[…]
the equations (2.18) and (2.19) in Proposition 2.2 are then proved by combining (2.13), (2.21), (2.22), (2.23), and Lemma 2.3.
Note that the left hand side of (2.23) is always 0 when $k=2$ and $m\equiv 1,3\ (\mathrm{mod}\ 4)$. Hence from there we cannot conclude the relation between $S(m,4n,c,\nu_{\Theta}^{4})$ and $S(m,n,\widetilde{c},\nu_{\Theta}^{4})$. In fact, for $k=2$ we need to modify the coefficients in (2.20). By (2.22) we have
$S(2m,2n,c,\nu_{\Theta}^{4})$ $=\sum_{\begin{subarray}{c}d\ (\mathrm{mod}\ 2c)\\ ad\equiv 1(2c)\end{subarray}}{\mathbf{e}}\left(\frac{2ma}{2c}\right){\mathbf{e}}\left(\frac{2nd}{2c}\right)$
[…]
$=2\sum_{\begin{subarray}{c}d\ (\mathrm{mod}\ 2\widetilde{c})\\ ad\equiv 1(\widetilde{c})\end{subarray}}{\mathbf{e}}\left(\frac{m(4\cdot\overline{4_{\widetilde{c}}})a}{2\widetilde{c}}\right){\mathbf{e}}\left(\frac{n(4\cdot\overline{4_{\widetilde{c}}})d}{2\widetilde{c}}\right){\mathbf{e}}\left(\frac{2(m+n)\widetilde{c}}{4}\right)$
We finish the proof of (2.20) by Lemma 2.3 and ${\mathbf{e}}(\frac{k}{4})=-1$.
2.3. Maass forms
Details in this subsection can be find in various references [Pro05, DFI02, DFI12, AA18, AD19]; we are mainly following [Sun24b, §3].
We call a function $f:{\mathbb{H}}\rightarrow{\mathbb{C}}$ automorphic of weight $k$ with a multiplier $\nu$ on a congruence subgroup $\Gamma$ if
$\left(\frac{cz+d}{|cz+d|}\right)^{-k}f(\gamma\tau)=\nu(\gamma)f(\tau)\quad\text{for all }\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma.$
Let $\mathcal{A}_{k}(\Gamma,\nu)$ denote the linear space consisting of all such functions, and let ${\mathcal{L}}_{k}(\Gamma,\nu)$ denote the space of square-integrable functions on $\Gamma\setminus{\mathbb{H}}$ with respect to the hyperbolic measure
$d\mu(\tau)=\frac{dxdy}{y^{2}}\,.$
The space ${\mathcal{L}}_{k}(\Gamma,\nu)$ is equipped with the Petersson inner product
$\langle f,g\rangle_{\Gamma}\vcentcolon=\int_{\Gamma\setminus{\mathbb{H}}}f(\tau)\overline{g(\tau)}\frac{dxdy}{y^{2}}\quad\text{for }f,g\in{\mathcal{L}}_{k}(\Gamma,\nu).$ (2.24)
The weight $k$ hyperbolic Laplacian is defined as
$\Delta_{k}\vcentcolon=y^{2}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)-iky\frac{\partial}{\partial x}.$
It is known that $\Delta_{k}$ is a self-adjoint operator on the Hilbert space ${\mathcal{L}}_{k}(\Gamma,\nu)$. The spectrum of $\Delta_{k}$ contains two parts: the continuous spectrum $\lambda\in[\frac{1}{4},\infty)$ and the discrete spectrum of finite multiplicity $\lambda_{0}=\frac{|k|}{2}(1-\frac{|k|}{2})<\lambda_{1}<\cdots\rightarrow\infty$.
Let $M_{k}(\Gamma,\nu)$ denote the space of weight $k$ holomorphic modular forms on $(\Gamma,\nu)$. We recall the Serre-Stark basis theorem.
Theorem 2.4 ([SS76, Theorem A]).
The basis of $M_{\frac{1}{2}}(\Gamma_{0}(N),\nu_{\theta})$ consists of the theta series
$\theta_{\psi,t}(\tau)=\sum_{n\in{\mathbb{Z}}}\psi(n){\mathbf{e}}(tn^{2}z)$
where $t\in{\mathbb{Z}}_{+}$ and $\psi$ is an even primitive character with conductor $r(\psi)$, satisfying both $4r(\psi)^{2}t|N$ and $\psi=(\frac{D}{\cdot})$ where $D$ is the discriminant of ${\mathbb{Q}}(\sqrt{t})/{\mathbb{Q}}$.
We have the following simple corollary.
Corollary 2.5.
The two spaces are identical: $M_{\frac{1}{2}}(\Gamma_{0}(4),\nu_{\theta})=M_{\frac{1}{2}}(\Gamma_{0}(8),\nu_{\theta})$. Both of them are one-dimensional and generated by $\theta$, defined in (2.11).
Let ${\widetilde{\mathcal{L}}}_{k}(\Gamma,\nu,\lambda)\subset{\mathcal{L}}_{k}(\Gamma,\nu)$ be the subspace spanned by eigenfunctions of $\Delta_{k}$ on $(\Gamma,\nu)$ with eigenvalue $\lambda$. There is a one-to-one correspondence between all $f\in{\widetilde{\mathcal{L}}}_{k}(\Gamma,\nu,\lambda_{0})$ and $F\in M_{k}(\Gamma,\nu)$ by
$f(\tau)=\left\{\begin{array}[]{ll}y^{\frac{k}{2}}F(\tau),&\quad k\geq 0,\ F\in M_{k}(\Gamma,\nu);\\ y^{-\frac{k}{2}}\overline{F(\tau)},&\quad k<0,\ F\in M_{-k}(\Gamma,\overline{\nu}).\end{array}\right$. (2.25)
For example, every $f\in{\mathcal{L}}_{\frac{1}{2}}(\Gamma_{0}(4),\nu_{\theta},\frac{3}{16})$ and every $g\in{\mathcal{L}}_{-\frac{1}{2}}(\Gamma_{0}(4),\overline{\nu_{\theta}},\frac{3}{16})$ are of the form
$f(\tau)=C_{f}y^{\frac{1}{4}}\theta(\tau),\quad g(\tau)=C_{g}y^{\frac{1}{4}}\overline{\theta(\tau)}\quad\text{for some }C_{f},C_{g}\in{\mathbb{C}},$ (2.26)
Moreover, such $f$ and $g$ are eigenfunctions of corresponding hyperbolic Laplacian with eigenvalue $\lambda_{0}=\frac{3}{16}$. By [Sun25, (2.18)] (or also [AA16, (6.2)]) and (2.26) above, their Fourier expansions satisfy:
$\begin{split}f(\tau)&=C_{f}y^{\frac{1}{4}}+\sum_{n=1}^{\infty}\rho_{0}^{(f)}(n)W_{\frac{1}{4},\frac{1}{4}}(4\pi ny){\mathbf{e}}(nx)=C_{f}y^{\frac{1}{4}}+\sum_{m=1}^{\infty}2C_{f}y^{\frac{1}{4}}{\mathbf{e}}(m^{2}\tau),\\ g(\tau)&=C_{g}y^{\frac{1}{4}}+\sum_{n=-1}^{-\infty}\rho_{0}^{(g)}(n)W_{\frac{1}{4},\frac{1}{4}}(4\pi|n|y){\mathbf{e}}(nx)=C_{g}y^{\frac{1}{4}}+\sum_{m=1}^{\infty}2C_{g}y^{\frac{1}{4}}{\mathbf{e}}(-m^{2}\overline{\tau}).\end{split}$ (2.27)
We need to know $\rho_{0}^{(f)}(n)$ and $\rho_{0}^{(g)}(n)$ when $f$ and $g$ are normalized eigenforms, i.e. when $\langle f,f\rangle_{\Gamma_{0}(4)}=\langle g,g\rangle_{\Gamma_{0}(4)}=1$, in order to apply theorems like [GS83, Theorem 2] to estimate sums of Kloosterman sums. If $\langle f,f\rangle_{\Gamma_{0}(4)}=\langle g,g\rangle_{\Gamma_{0}(4)}=1$, then
$C_{f}=C_{g}=(2\pi)^{-\frac{1}{2}}$
due to the regularized Petersson inner product on the theta function:
$\int_{\Gamma_{0}(4)\setminus{\mathbb{H}}}y^{\frac{1}{2}}|\theta(\tau)|^{2}\frac{dxdy}{y^{2}}=2\pi.$
This result can be found in [Chi07, Theorem 2.2], noting that $\operatorname{SL}_{2}({\mathbb{Z}}):\Gamma_{0}(4)]=6$. Similarly, identities
$\int_{\Gamma_{0}(8)\setminus{\mathbb{H}}}y^{\frac{1}{2}}|\theta(\tau)|^{2}\frac{dxdy}{y^{2}}=4\pi\quad\text{and}\quad[\operatorname{SL}_{2}({\mathbb{Z}}):\Gamma_{0}(8)]=12$
help us deal with the case on $\Gamma_{0}(8)$.
By [DLMF, (13.18.2)], $W_{\frac{1}{4},\frac{1}{4}}(4\pi|n|y)=e^{-2\pi|n|y}(4\pi|n|y)^{\frac{1}{4}}$. Therefore, we get $\rho_{0}^{(f)}(n)$ and $\rho_{0}^{(g)}(n)$ by comparing coefficients in (2.27).
Lemma 2.6.
Suppose $f\in{\widetilde{\mathcal{L}}}_{\frac{1}{2}}(\Gamma_{0}(4),\nu_{\theta},\frac{3}{16})$, $g\in{\widetilde{\mathcal{L}}}_{-\frac{1}{2}}(\Gamma_{0}(4),\overline{\nu_{\theta}},\frac{3}{16})$, $\langle f,f\rangle_{\Gamma_{0}(4)}=1$, and $\langle g,g\rangle_{\Gamma_{0}(4)}=1$. Then $f$ and $g$ have the Fourier expansion as in (2.27) with coefficients
$\rho_{0}^{(f)}(n)=\left\{\begin{array}[]{ll}(\pi^{3}n)^{-\frac{1}{4}},&\ n=m^{2}>0;\\ 0,&\ \text{other }n\neq 0,\end{array}\right.\quad\rho_{0}^{(g)}(n)=\left\{\begin{array}[]{ll}|\pi^{3}n|^{-\frac{1}{4}},&\ n=-m^{2}<0;\\ 0,&\ \text{other }n\neq 0.\end{array}\right$. (2.28)
Similarly, if $f\in{\widetilde{\mathcal{L}}}_{\frac{1}{2}}(\Gamma_{0}(8),\nu_{\theta},\frac{3}{16})$, $g\in{\widetilde{\mathcal{L}}}_{-\frac{1}{2}}(\Gamma_{0}(8),\overline{\nu_{\theta}},\frac{3}{16})$, and $\langle f,f\rangle_{\Gamma_{0}(8)}=\langle g,g\rangle_{\Gamma_{0}(8)}=1$, then
$\rho_{0}^{(f)}(n)=\left\{\begin{array}[]{ll}(4\pi^{3}n)^{-\frac{1}{4}},&\ n=m^{2}>0;\\ 0,&\ \text{other }n\neq 0,\end{array}\right.\quad\rho_{0}^{(g)}(n)=\left\{\begin{array}[]{ll}|4\pi^{3}n|^{-\frac{1}{4}},&\ n=-m^{2}<0;\\ 0,&\ \text{other }n\neq 0.\end{array}\right$. (2.29)
For the discrete spectrum of hyperbolic Laplacian, other than $\lambda_{0}$, Selberg conjectured [Sel65] that $\lambda_{1}\geq\frac{1}{4}$ for $\Delta_{0}$ on $(\Gamma,\text{id})$ for all congruence subgroups $\Gamma$ of $\operatorname{SL}_{2}({\mathbb{Z}})$ and showed that $\lambda_{1}\geq\frac{3}{16}$. We call any $\lambda\in(\lambda_{0},\frac{1}{4})$ as an exceptional eigenvalue. The best progress is known today is $\lambda_{1}\geq\frac{1}{4}-\left(\frac{7}{64}\right)^{2}$ for $(\Gamma_{0}(N),\text{id})$, for all $N\in{\mathbb{Z}}_{+}$ by [Kim03]. For small $N$, the following fact is also known.
Lemma 2.7.
[Hux85, BS07] There is no exceptional eigenvalue on $\Gamma_{0}(N)$ for $N\leq 18$, i.e. Selberg’s eigenvalue conjecture is known to be true for $N\leq 18$.
2.4. Some useful elementary inequalities
For $r\in{\mathbb{R}}$, let $\left\lceil r\right\rceil$ denote the smallest integer larger than or equal to $r$, and $\left\lfloor r\right\rfloor$ denote the largest integer smaller than or equal to $r$. Let $\|r\|=\min(r-\left\lfloor r\right\rfloor,\left\lceil r\right\rceil-r)$ denote the distance from $r$ to its closest integer.
The following lemmas will be helpful in the proofs.
Lemma 2.8.
For $r\in{\mathbb{R}}\setminus{\mathbb{Z}}$, $a,c\in{\mathbb{Z}}$, $c\leq a<c$, we have
${\mathbf{e}}\left(\frac{ra}{2c}\right)=\frac{1}{2c}\sum_{k\ (\mathrm{mod}\ 2c)}\frac{2i\sin(\pi(r-k))}{{\mathbf{e}}(\frac{r-k}{2c})-1}{\mathbf{e}}\left(\frac{ka}{2c}\right).$ (2.30)
Proof.
On the right hand side, note that the summand (as a function of $k$) is periodic modulo $2c$, so the sum is well-defined. We define and compute the discrete Fourier transform:
$x_{c}(k)\vcentcolon=\sum_{a=-c}^{c-1}{\mathbf{e}}\left(\frac{ra}{2c}\right){\mathbf{e}}\left(-\frac{ka}{2c}\right)=\frac{2i\sin(\pi(r-k))}{{\mathbf{e}}(\frac{r-k}{2c})-1}.$
It is then straighforward to verify that for any $d\in{\mathbb{Z}}$ and $d\equiv a\ (\mathrm{mod}\ 2c)$,
$\frac{1}{2c}\sum_{k\ (\mathrm{mod}\ 2c)}x_{c}(k){\mathbf{e}}\left(\frac{kd}{2c}\right)={\mathbf{e}}\left(\frac{ra}{2c}\right).\qed$
Remark.
When applying the lemma above, we will usually require the range |rk|<c|r-k|<c for k(mod 2c)k\ ($\mathrm${mod}\ 2c), because $z=\frac{r-k}{2c}$ will satisfy $|z|<\frac{1}{2}$ and we would be able to use the expansion from the lemma that follows.
Lemma 2.9.
For $z\in{\mathbb{C}}$ and $0<|z|<1$, we have
$\frac{1}{{\mathbf{e}}(z)-1}=\frac{1}{2\pi iz}+\sum_{\ell=1}^{\infty}B_{\ell}\frac{(2\pi iz)^{\ell-1}}{\ell!},$
where $B_{\ell}$ is the $\ell$-th Bernoulli number, $B_{0}=1$, $B_{1}=-\frac{1}{2}$.
We will also need the basic inequality
$(X+Y)^{\alpha}\leq X^{\alpha}+Y^{\alpha}\quad\text{for }X,Y>0\text{ and }\alpha\in[0,1]$ (2.31)
and the following lemma.
Lemma 2.10.
Let $r\in{\mathbb{R}}_{+}\setminus{\mathbb{Z}}$ and $x\geq 1$. Then for $\alpha\in(0,1]$ we have
$\left.\begin{array}[]{r}\displaystyle\sum_{k=\left\lceil r-x\right\rceil}^{\left\lfloor r\right\rfloor}\frac{k^{\alpha}}{r-k}\\ \displaystyle\sum_{k=\left\lceil r\right\rceil}^{\left\lfloor r+x\right\rfloor}\frac{k^{\alpha}}{k-r}\end{array}\right\}\ll_{\varepsilon}\frac{(r^{\alpha}+1)x^{\varepsilon}}{\|r\|}+\frac{x^{\alpha}}{\alpha}\quad\text{for any }\varepsilon>0.$
Proof.
These inequalities are helpful: $\sum_{u=1}^{n}u^{-1}\leq\log n+1$, and for $\alpha\in(0,1]$,
$\sum_{u=1}^{n}u^{\alpha-1}\leq 1+\int_{1}^{n}x^{\alpha-1}dx\leq 1+\frac{n^{\alpha}}{\alpha}.$
First we consider $\sum_{k=\left\lceil r-x\right\rceil}^{\left\lfloor r\right\rfloor}\frac{k^{\alpha}}{r-k}$. For $k=\left\lfloor r\right\rfloor$, the summand is bounded by $r^{\alpha}/\|r\|$ (note that $0<r<1$ implies $\left\lfloor r\right\rfloor=0$). For the other terms we have
$\sum_{k=\left\lceil r-x\right\rceil}^{\left\lfloor r\right\rfloor-1}\frac{k^{\alpha}}{r-k}\leq\sum_{u=1}^{\left\lfloor x\right\rfloor}\frac{r^{\alpha}+u^{\alpha}}{u}\leq r^{\alpha}(1+\log x)+\frac{x^{\alpha}}{\alpha}+1.$
Next we prove the bound for $\sum_{k=\left\lceil r\right\rceil}^{\left\lfloor r+x\right\rfloor}\frac{k^{\alpha}}{k-r}$. The first term $k=\left\lceil r\right\rceil$ gives
$\frac{\left\lceil r\right\rceil^{\alpha}}{\left\lceil r\right\rceil-r}\leq\frac{(r+1)^{\alpha}}{\|r\|}\leq\frac{r^{\alpha}+1}{\|r\|}$
by (2.31). The remaining terms are bounded by
$\sum_{k=\left\lceil r\right\rceil+1}^{\left\lfloor r+x\right\rfloor}\frac{k^{\alpha}}{k-r}\leq\sum_{u=1}^{\left\lfloor x\right\rfloor}\frac{u^{\alpha}+r^{\alpha}+1}{u}\leq(r^{\alpha}+1)(\log x+1)+\frac{x^{\alpha}}{\alpha}+1.$
Combining the calculations above we get
$\left.\begin{array}[]{r}\displaystyle\sum_{k=\left\lceil r-x\right\rceil}^{\left\lfloor r\right\rfloor}\frac{k^{\alpha}}{r-k}\\ \displaystyle\sum_{k=\left\lceil r\right\rceil}^{\left\lfloor r+x\right\rfloor}\frac{k^{\alpha}}{k-r}\end{array}\right\}\leq\frac{r^{\alpha}+1}{\|r\|}+(r^{\alpha}+1)(\log x+1)+\frac{2x^{\alpha}}{\alpha}+1.$ (2.32)
The concluded bound in the lemma is clear. ∎
3. Construction of modular integrals via Poincaré-type sums
In this section we recall Stoller’s construction of modular integrals for the group $\Gamma(2)$ via Poincaré-type series. First, we recall a few basic notions on cocycles and modular integrals and explain the connection to Theorem 1.1.
Given a Fuchsian group $\Gamma\subset\operatorname{PSL}_{2}({\mathbb{R}})$ acting on functions on the upper half-plane via slash operators
$|\gamma=|_{k,\nu}\gamma:\quad(f|\gamma)(\tau)=\overline{\nu(\gamma)}j(\gamma,\tau)^{-k}f(\gamma\tau)$
(for simplicity, we will omit the weight and the multiplier system from notation), we say that a collection of functions $\{f_{\gamma}\}_{\gamma\in\Gamma}$ is a cocycle if
$f_{\gamma_{1}\gamma_{2}}=f_{\gamma_{1}}|\gamma_{2}+f_{\gamma_{2}}\,,\qquad\gamma_{1},\gamma_{2}\in\Gamma.$
A cocycle $\{f_{\gamma}\}_{\gamma}$ is called trivial if there exists a function $F$ such that $f_{\gamma}=F-F|\gamma$ for all $\gamma\in\Gamma$. If this is the case, we say that $F$ is a modular integral for the cocycle $\{f_{\gamma}\}_{\gamma}$. We are usually interested in cocycles restricted to a certain class of functions on $\mathbb{H}$}, for example, with all $f_{\gamma}$ holomorphic and of moderate growth, in which case we require $F$ to belong to the same class. (The terminology modular integral comes from Eichler integrals, which are modular integrals for polynomial-valued cocycles).
For the group $\Gamma(2)$, which is free on two generators $T^{2}$ and $ST^{2}S$, any cocycle is uniquely determined by an arbitrary choice of two functions $\varphi_{1}=f_{T^{2}}$ and $\varphi_{2}=f_{ST^{2}S}$. It follows from the results of Knopp [Kno74] (for weights $k>2$), and, e.g., from [BRS23, Theorem 3.1] (for all $k\geq 0$), that any cocycle (with values in holomorphic functions of moderate growth) for the group $\Gamma(2)$ in non-negative weight is trivial. Hence for any holomorphic, moderately growing $\varphi_{i}$, $i=1,2,$ there exists a holomorphic function of moderate growth $F$ such that
$\begin{cases}F-F|T^{2}=\varphi_{1},\\ F-F|ST^{2}S=\varphi_{2}.\end{cases}$
Choosing $\varphi_{1}=0$ and $\varphi_{2}(\tau)=\varphi-\varphi|ST^{2}S$, where $\varphi(\tau)=e^{\pi ir^{2}\tau}$ then leads to functions satisfying the conditions of Theorem 1.1. In [RV19] and [BRS23] modular integrals are constructed using contour integrals against modular Green functions, which is convenient for proving analytic properties, but is hard to use to estimates the size of the Fourier coefficients of $F$. In [Sto21] it is shown that $F$ can be constructed as $F=\sum_{\gamma\in\mathcal{B}}\varphi|\gamma$, for a suitably chosen set $\mathcal{B}$ of coset representatives for $\Gamma(2)_{\infty}\setminus\Gamma(2)$ (since $\varphi$ is not assumed to be periodic, the choice of representatives matters). We call these Poincaré-type series since they becomes the usual Poincaré series in the case when $\varphi$ is a 2-periodic function.
Unfortunately, the above Poincaré-type series only converges when the weight $k$ is larger than $2$ and the resulting modular integral does not agree with the one obtained by contour integrals. We overcome these difficulties by considering $F_{s}=\sum_{\gamma\in\mathcal{B}}\varphi_{s}|\gamma$ (a Maass–Poincaré-type series) with $\varphi_{s}$ an eigenfunction of the weight $k$ hyperbolic Laplace operator. We show that the series converges for $\operatorname{Re}(s)>1$, and via its Fourier expansion $F_{s}$ admits an analytic continuation (in the variable $s$) to the point $s=k/2$. For this value of $s$ the function $\varphi_{s}$ specializes to $e^{\pi ir^{2}\tau}$ and $F_{s}$ becomes harmonic. By analyzing the specialization more carefully we are then able to precisely identify the holomorphic part of $F_{k/2}$ and relate it to the solutions from Theorem 1.1.
3.1. A special choice of coset representatives
Following [Sto21], we make a special choice of coset representatives for $\Gamma(2)_{\infty}\setminus\Gamma(2)$ and $(\Gamma(2)_{\infty}\setminus\Gamma(2))S$. Recall the notation from §2: $\overline{\Gamma}=\Gamma/\{\pm I\}$, $A=[T^{2}]$ and $B=[ST^{2}S]$ are the classes of $T^{2}$ and $ST^{2}S$, respectively. Also recall that $\Gamma_{\Theta}=\Gamma(2)\sqcup\Gamma(2)S$.
Definition 3.1 ([Sto21, Definition 5.1]).
The subset ${\mathcal{B}}\subset\overline{\Gamma(2)}$ is defined as the set of all nonempty finite reduced words in $A$ and $B$ that start with a nonzero power of $B$. More formally, an element $\gamma\in\overline{\Gamma(2)}$ belongs to $\mathcal{B}$}, if and only if there are integers $m\geq 1$, and $e_{1},\cdots,e_{m},f_{1},\cdots,f_{m}$, all non-zero, except possibly $e_{m}$, such that $\gamma=B^{f_{1}}A^{e_{1}}\cdots B^{f_{m}}A^{e_{m}}$. We also define
$\widetilde{{\mathcal{B}}}\vcentcolon={\mathcal{B}}[S]\sqcup\{[S]\}=\{\gamma[S]:\gamma\in{\mathcal{B}}\}\sqcup\{[S]\}\subset\overline{\Gamma(2)}.$
Proposition 3.2 ([Sto21, Lemma 5.1, Lemma 5.2]).
We have the following properties.
  • (1)
    For each $\gamma\in{\mathcal{B}}$, each $\widetilde{\gamma}\in\widetilde{{\mathcal{B}}}$, and each $\ell\in{\mathbb{Z}}$, one has $\gamma A^{\ell}\in{\mathcal{B}}$ and $\widetilde{\gamma}A^{\ell}\in\widetilde{{\mathcal{B}}}$.
  • (2)
    Define
    $\mathcal{P}$ $=\{(c,d)\in{\mathbb{Z}}^{2}:\ \gcd(c,d)=1,\ c\equiv 0,\ d\equiv 1\ (\mathrm{mod}\ 2),\ c\neq 0\},$
    $\mathcal{P}_{I}$ $=\mathcal{P}\sqcup\{(0,1),\ (0,-1)\},$
    $\widetilde{\mathcal{P}}$ $=\{(c,d)\in{\mathbb{Z}}^{2}:\ \gcd(c,d)=1,\ c\equiv 1,\ d\equiv 0\ (\mathrm{mod}\ 2)\}.$
    Then
    $\left[\begin{pmatrix}a&b\\ c&d\end{pmatrix}\right]\rightarrow[(c,d)]$
    defines $\mathbb{Z}$}-equivariant bijections ${\mathcal{B}}\cong\mathcal{P}/\{\pm 1\}$, ${\mathcal{B}}\sqcup\{[I]\}\cong\mathcal{P}_{I}/\{\pm 1\}$ and $\widetilde{{\mathcal{B}}}\cong\widetilde{\mathcal{P}}/\{\pm 1\}$.
  • (3)
[…]
Remark.
If we denote by $\mathcal{D}=\{z\in{\mathbb{H}}\colon|\operatorname{Re}(z)|<1,|z\pm 1/2|>1/2\}$ the standard fundamental domain for $\Gamma(2)$, then from the figure below we can see that this choice of coset representatives ensures that $\{\gamma\mathcal{D}:\gamma\in\mathcal{B}\}$ fill the region $\mathcal{D}_{\infty}=\{z\in{\mathbb{H}}:|\operatorname{Re}(z)|<1\}$ up to a measure-zero set. In other words, when doing the “unfolding trick” to compute the Fourier expansion of Maass–Poincaré series in §3.4 below, we end up with an integral over $\mathcal{D}_{\infty}$, same as in the classical case.
[…]
$ST^{-4}$
$ST^{-2}S$
$ST^{2}ST^{2}$
$ST^{-2}ST^{-2}$
[…]
3.2. Real-variable Kloosterman sums
In the proof we need to define real-variable Kloosterman sums analogous to (2.5). Let $r\in{\mathbb{R}}$, $n\in{\mathbb{Z}}$, and $k\in\frac{1}{2}{\mathbb{Z}}$. For $2|c>0$, we define
$\begin{split}K(r,n,c,\nu_{\Theta}^{2k})&\vcentcolon=\sum_{\begin{subarray}{c}-c<a<c\\ (a,2c)=1\\ ad\equiv 1\ (\mathrm{mod}\ 2c)\end{subarray}}\nu_{\Theta}\begin{pmatrix}a&*\\ c&d\end{pmatrix}^{-2k}{\mathbf{e}}\left(\frac{ra+nd}{2c}\right)\\ &=\sum_{\begin{subarray}{c}-c<a<c\\ (a,2c)=1\\ ad\equiv 1\ (\mathrm{mod}\ 2c)\end{subarray}}\varepsilon_{d}^{2k}\left(\frac{2c}{d}\right)^{-2k}{\mathbf{e}}\left(\frac{ra+nd}{2c}\right);\end{split}$ (3.1)
for $2\nmid d>0$, we define
$\begin{split}\widetilde{K}(r,n,d,\nu_{\Theta}^{2k})&\vcentcolon=\sum_{\begin{subarray}{c}-d<b<d\\ 2|b,\ (b,d)=1\\ 2|c,\ bc\equiv-1\ (\mathrm{mod}\ d)\end{subarray}}\nu_{\Theta}\begin{pmatrix}b&*\\ d&-c\end{pmatrix}^{-2k}{\mathbf{e}}\left(\frac{rb-nc}{2d}\right)\\ &=\sum_{\begin{subarray}{c}-d<b<d\\ 2|b,\ (b,d)=1\\ 2|c,\ bc\equiv-1\ (\mathrm{mod}\ d)\end{subarray}}{\mathbf{e}}(\tfrac{k}{4})\varepsilon_{d}^{-2k}\left(\frac{-2c}{d}\right)^{-2k}{\mathbf{e}}\left(\frac{rb-nc}{2d}\right).\end{split}$ (3.2)
As a special case, we set $\widetilde{K}(r,n,1,\nu_{\Theta}^{2k})={\mathbf{e}}(\frac{k}{4})$.
Remark.
Here are a few remarks about real-variable Kloosterman sums.
  • (1)
    Note that we specify c<a<c-c<a<c in (3.1) and d<b<d-d<b<d in (3.2). This restriction is important because $r$ is a real number and changing $a\rightarrow a+2c$ (or $b\rightarrow b+2d$) will change the value of the sum. This aligns with part (3) of Proposition 3.2.
  • (2)
    The function $K(r,n,c,\nu_{\Theta}^{2k})$ is the real-variable analogue of $S(m,n,c,\nu_{\Theta}^{2k})$ for $2|c$, in the sense that setting $r=m$ specializes $K$ to $S$, and the function $\widetilde{K}(r,n,d,\nu_{\Theta}^{2k})$ is the real-variable analogue of $S(m,n,d,\nu_{\Theta}^{2k})$ for $2\nmid d$.
  • (3)
    This definition of real-variable Kloosterman sums originated from [Sto21, (8.4)]. Stoller’s estimate focused on the properties of Poincaré series. In this paper, we prove properties of these Kloosterman sums directly with more precise results.
  • (4)
    To avoid ambiguity about the meaning of $d$, we will use $\widetilde{K}(m,n,\widetilde{c},\nu_{\Theta}^{2k})$ for $2\nmid\widetilde{c}>0$ in the rest of the paper.
3.3. Maass–Poincaré type series
Let $M_{\beta,\mu}$ and $W_{\beta,\mu}$ denote the $M$- and $W$-Whittaker functions (for definition see [DLMF, (13.14.2-3)]). For $s\in{\mathbb{C}}$, $x,y\in{\mathbb{R}}$, $z=x+iy$ and $k\in\frac{1}{2}{\mathbb{Z}}$, we define
${\mathcal{M}}_{s}(y)\vcentcolon=|y|^{-\frac{k}{2}}M_{\frac{k}{2}\operatorname{sgn}y,\,s-\frac{1}{2}}(|y|)\quad\text{and}\quad\mathcal{W}_{s}(y)\vcentcolon=|y|^{-\frac{k}{2}}W_{\frac{k}{2}\operatorname{sgn}y,\,s-\frac{1}{2}}(|y|).$ (3.3)
We also define
$\varphi_{s,k}(x+iy)\vcentcolon={\mathcal{M}}_{s}(4\pi y){\mathbf{e}}(x).$ (3.4)
These functions have the following properties. For $y>0$, by [DLMF, (13.18.2)], we have
${\mathcal{M}}_{\frac{k}{2}}(y)=y^{-\frac{k}{2}}M_{\frac{k}{2},\,\frac{k}{2}-\frac{1}{2}}(y)=e^{-\frac{y}{2}};$ (3.5)
by [DLMF, (13.14.31), (13.18.2)], we have
$W_{\kappa,\mu}(z)=W_{\kappa,-\mu}(z),\quad W_{-\frac{k}{2},\,\frac{k}{2}-\frac{1}{2}}(y)=y^{\frac{k}{2}}e^{\frac{y}{2}}\Gamma(1-k,y)\ \text{and}\ W_{\frac{k}{2},\,\frac{k}{2}-\frac{1}{2}}(y)=y^{\frac{k}{2}}e^{-\frac{y}{2}};$ (3.6)
moreover, by [DLMF, (13.14.14)], we have
$\varphi_{s,k}(z)=O\left(y^{\operatorname{Re}s-\frac{k}{2}}\right)\quad\text{for }z\rightarrow 0.$ (3.7)
Recall Definition 3.1 for $\mathcal{B}$} and $\widetilde{{\mathcal{B}}}$. For $\tau=x+iy\in{\mathbb{H}}$, $r\in{\mathbb{R}}\setminus\{0\}$, $k\in\frac{1}{2}{\mathbb{Z}}$, and $\operatorname{Re}(s)>1$, we define the series
$\begin{split}&F_{k}(\tau;r,s)\vcentcolon=-\sum_{\gamma=\begin{pmatrix}*&*\\ c&d\end{pmatrix}\in{\mathcal{B}}}\nu_{\Theta}(\gamma)^{-2k}(c\tau+d)^{-k}\varphi_{s,k}(\tfrac{r^{2}}{2}\gamma\tau),\\ &\widetilde{F}_{k}(\tau;r,s)\vcentcolon=\sum_{\widetilde{\gamma}=\begin{pmatrix}*&*\\ c&d\end{pmatrix}\in\widetilde{{\mathcal{B}}}}\nu_{\Theta}(\gamma)^{-2k}(c\tau+d)^{-k}\varphi_{s,k}(\tfrac{r^{2}}{2}\widetilde{\gamma}\tau).\end{split}$ (3.8)
These two series converge uniformly and absolutely for $s$ in any compact subset of $\{s\in{\mathbb{C}}:\operatorname{Re}(s)>1\}$ due to (3.7). Note that $\mathcal{B}$} and $\widetilde{{\mathcal{B}}}$ are defined up to the equivalence relation modulo $\{\pm I\}$. The series in (3.8) are well-defined because
$\nu_{\Theta}(-\gamma)^{-1}(-cz-d)^{-\frac{1}{2}}=-i\nu_{\Theta}(\gamma)^{-1}\cdot i(cz+d)^{-\frac{1}{2}}=\nu_{\Theta}(\gamma)^{-1}(cz+d)^{-\frac{1}{2}}$
by (2.4) for $c>0$ and $\nu_{\Theta}\begin{pmatrix}*&*\\ 0&*\end{pmatrix}=1$.
For all $r\in{\mathbb{R}}$, we have the following functional equation.
Proposition 3.3.
For $r\in{\mathbb{R}}$, $k\in\tfrac{1}{2}{\mathbb{Z}}$, $\tau=x+iy\in{\mathbb{H}}$, and for $F_{k}(\tau,r;s)$, $\widetilde{F}_{k}(\tau,r;s)$ defined at (3.8), we have
$F_{k}(\tau;r,s)+(-i\tau)^{-k}\widetilde{F}_{k}(-\tfrac{1}{\tau};r,s)=\varphi_{s,k}(\tfrac{r^{2}}{2}\tau)=(2\pi r^{2}y)^{-\frac{k}{2}}M_{\frac{k}{2},s-\frac{1}{2}}(2\pi r^{2}y)e^{\pi ir^{2}x}.$ (3.9)
Proof.
We start with $\widetilde{F}_{k}(-\frac{1}{\tau};r,s)$. By Proposition 3.2, we choose representatives of $\mathcal{P}/{\mathbb{Z}}^{\times}$ and $\widetilde{\mathcal{P}}$$\mathbb{Z}$$^{\times}$ as
$\{(c,d)\in{\mathbb{Z}}\times{\mathbb{Z}}:c=2,4,6,\cdots,\ (c,d)=1\},\quad\text{and}$
$\{(1,0)\}\cup$ $\{(c,-d)\in{\mathbb{Z}}\times{\mathbb{Z}}:d=2,4,6,\cdots,\ (c,d)=1\},\quad\text{respectively.}$
The corresponding representatives of $\mathcal{B}$} and $\widetilde{{\mathcal{B}}}$ are chosen by
$\mathfrak{B}\vcentcolon=\{\begin{pmatrix}a&b+2ta\\ c&d+2tc\end{pmatrix}\in\Gamma_{\Theta}:c\in 2{\mathbb{Z}}_{+},\ d\ (\mathrm{mod}\ 2c)^{*},\ ad\equiv 1\ (\mathrm{mod}\ 2c),\ t\in{\mathbb{Z}}\},$ (3.10)
$\text{and}\quad\{\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\}\cup\mathfrak{B}S,\quad\text{respectively}.$ (3.11)
For $\tau\in{\mathbb{H}}$ and $\gamma=\begin{pmatrix}*&*\\ c&d\end{pmatrix}\in\Gamma_{\Theta}$ with $c>0$, by (2.10) we get
$\nu_{\Theta}(\gamma S)\left(d(-\tfrac{1}{\tau})-c\right)^{\frac{1}{2}}$ $=(\operatorname{sgn}d){\mathbf{e}}(-\tfrac{1}{8})\nu_{\Theta}(\gamma)\cdot(\operatorname{sgn}d)(-\tfrac{1}{\tau})^{\frac{1}{2}}(c\tau+d)^{\frac{1}{2}}$
$={\mathbf{e}}(\tfrac{1}{8})\tau^{-\frac{1}{2}}\nu_{\Theta}(\gamma)(c\tau+d)^{\frac{1}{2}}$
Hence we have
$\widetilde{F}_{k}(-\tfrac{1}{\tau};r,s)=\nu_{\Theta}(S)^{-2k}(-\tau^{-1})^{-k}\varphi_{s,k}(\tfrac{r^{2}}{2}\tau)$
$+\sum_{2|c>0}\sum_{t\in{\mathbb{Z}}}\sum_{\begin{subarray}{c}d\ (\mathrm{mod}\ 2c)^{*}\\ ad\equiv 1(2c)\end{subarray}}\nu_{\Theta}\left(\begin{pmatrix}a&b+2ta\\ c&d+2tc\end{pmatrix}S\right)^{-2k}\frac{\varphi_{s,k}\left(\frac{r^{2}}{2}\begin{pmatrix}a&b+2ta\\ c&d+2tc\end{pmatrix}S(-\frac{1}{\tau})\right)}{\left((d+2tc)(-\frac{1}{\tau})-c\right)^{k}}$
$={\mathbf{e}}(\tfrac{k}{4})e^{-\pi ik}\tau^{k}\varphi_{s,k}(\tfrac{r^{2}}{2}\tau)+\sum_{\gamma\in{\mathcal{B}}}{\mathbf{e}}(-\tfrac{k}{4})\tau^{k}\nu_{\Theta}(\gamma)^{-2k}\frac{\varphi_{s,k}(\tfrac{r^{2}}{2}\gamma\tau)}{(c\tau+d)^{k}}$
$={\mathbf{e}}(-\tfrac{k}{4})\tau^{k}\left(\varphi_{s,k}(\tfrac{r^{2}}{2}\tau)-F_{k}(\tau;r,s)\right).$
We finish the proof by noticing $(-i\tau)^{k}={\mathbf{e}}(-\frac{k}{4})\tau^{k}$ for $\tau\in{\mathbb{H}}$. ∎
In Proposition 3.3, if we take $s=\frac{k}{2}$, by (3.5) we have
$F_{k}(\tau;r,\tfrac{k}{2})+(-i\tau)^{-k}\widetilde{F}_{k}(-\tfrac{1}{\tau};r,\tfrac{k}{2})=e^{\pi ir^{2}\tau}.$ (3.12)
This is clear if the weight $k>2$ because $s=\frac{k}{2}>1$. Since we are focusing on the case $k=2$ and $k=\frac{3}{2}$ in this paper, by computing their Fourier expansions in Lemma 3.4, we are able to prove the following two theorems, Theorem 3.5 and Theorem 3.6, which show that $F_{2}(\tau;r,s)$ and $\widetilde{F}_{2}(\tau;r,s)$ can be analytically continued to $s=1$ and $F_{\frac{3}{2}}(\tau;r,s)$ and $\widetilde{F}_{\frac{3}{2}}(\tau;r,s)$ can be analytically continued to $s=\frac{3}{4}$, respectively. After proving these theorems, we conclude that (3.12) still holds for $k=2$ and $\frac{3}{2}$ by analytic continuation.
3.4. Computation of Fourier expansions
Lemma 3.4.
Let $\tau=x+iy\in{\mathbb{H}}$. For $n\in{\mathbb{Z}}$, we define
$B_{2k,n}(r;s,y)\vcentcolon=-\Gamma(2s)\pi{\mathbf{e}}(-\tfrac{k}{4})\cdot$
$\left\{\begin{array}[]{ll}\displaystyle\frac{W_{\frac{k}{2},s-\frac{1}{2}}(2\pi ny)}{\Gamma(s+\frac{k}{2})(2\pi r^{2}y)^{\frac{k}{2}}}\left|\frac{r^{2}}{n}\right|^{\frac{1}{2}}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{2k})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right),&n>0,\\ \displaystyle\frac{(\pi r^{2})^{s-\frac{k}{2}}2^{-s-\frac{k}{2}}y^{1-s-\frac{k}{2}}}{\Gamma(s+\frac{k}{2})\Gamma(s-\frac{k}{2})(s-\frac{1}{2})}\sum_{2|c>0}\frac{K(r^{2},0,c,\nu_{\Theta}^{2k})}{c^{2s}},&n=0,\\ \displaystyle\frac{W_{-\frac{k}{2},s-\frac{1}{2}}(2\pi|n|y)}{\Gamma(s-\frac{k}{2})(2\pi r^{2}y)^{\frac{k}{2}}}\left|\frac{r^{2}}{n}\right|^{\frac{1}{2}}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{2k})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right),&n<0,\end{array}\right$.
and
$\widetilde{B}_{2k,n}(r;s,y)\vcentcolon=\Gamma(2s)\pi{\mathbf{e}}(-\tfrac{k}{4})\cdot$
$\left\{\begin{array}[]{ll}\displaystyle\frac{W_{\frac{k}{2},s-\frac{1}{2}}(2\pi ny)}{\Gamma(s+\frac{k}{2})(2\pi r^{2}y)^{\frac{k}{2}}}\left|\frac{r^{2}}{n}\right|^{\frac{1}{2}}\sum_{2\nmid d>0}\frac{\widetilde{K}(r^{2},n,d,\nu_{\Theta}^{2k})}{d}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{d}\right),&n>0,\\ \displaystyle\frac{(\pi r^{2})^{s-\frac{k}{2}}2^{-s-\frac{k}{2}}y^{1-s-\frac{k}{2}}}{\Gamma(s+\frac{k}{2})\Gamma(s-\frac{k}{2})(s-\frac{1}{2})}\sum_{2\nmid d>0}\frac{\widetilde{K}(r^{2},0,d,\nu_{\Theta}^{2k})}{d^{2s}},&n=0,\\ \displaystyle\frac{W_{-\frac{k}{2},s-\frac{1}{2}}(2\pi|n|y)}{\Gamma(s-\frac{k}{2})(2\pi r^{2}y)^{\frac{k}{2}}}\left|\frac{r^{2}}{n}\right|^{\frac{1}{2}}\sum_{2\nmid d>0}\frac{\widetilde{K}(r^{2},n,d,\nu_{\Theta}^{2k})}{d}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{d}\right),&n<0.\end{array}\right$.
Then the Fourier expansions of $F_{k}(\tau,r;s)$ and $\widetilde{F}_{k}(\tau,r;s)$ at the cusp at $\infty$ are
$F_{k}(\tau;r,s)$ $=\sum_{n\in{\mathbb{Z}}}B_{2k,n}(r;s,y)\;e^{\pi inx},$ (3.13)
$\widetilde{F}_{k}(\tau;r,s)$ $=\sum_{n\in{\mathbb{Z}}}\widetilde{B}_{2k,n}(r;s,y)\;e^{\pi inx}.$ (3.14)
Proof.
When $r^{2}$ is an integer, this is a standard computation in the theory of harmonic Maass forms, see [Bru02, Proof of Theorem 1.9], [BO06, Proof of Theorem 3.2], [BO12, Proof of Theorem 3.2], [DJ13, Theorem 3.2] and [Sun24a, §5.2] for examples. However, since we require $r$ in $\varphi_{s,k}(\frac{r^{2}}{2}\tau)$ to be real and use real-variable Kloosterman sums (3.1), (3.2) here, we provide a detailed proof.
We first define an auxiliary function
$f(\tau;c)\vcentcolon=\sum_{t\in{\mathbb{Z}}}\frac{|\tau+2t|^{k}}{(\tau+2t)^{k}}M_{\frac{k}{2},\,s-\frac{1}{2}}\left(\frac{2\pi r^{2}y}{c^{2}|\tau+2t|^{2}}\right){\mathbf{e}}\left(\frac{r^{2}/2}{-c^{2}}\operatorname{Re}\left(\frac{1}{\tau+2t}\right)\right).$
This function is invariant under $\tau\rightarrow\tau+2$ and has the Fourier expansion
$f(\tau;c)=:\sum_{n\in{\mathbb{Z}}}f_{n}(y)e^{\pi inx}\quad\text{for }\tau=x+iy,$
where the Fourier coefficient $f_{n}(y)$ can be computed by
$f_{n}(y)$ $=\frac{1}{2}\int_{-1}^{1}f(\tau;c)e^{-\pi inx}dx$
$=\frac{1}{2}\int_{{\mathbb{R}}}\frac{|x+iy|^{k}}{(x+iy)^{k}}M_{\frac{k}{2},\,s-\frac{1}{2}}\left(\frac{2\pi r^{2}y}{c^{2}(x^{2}+y^{2})}\right){\mathbf{e}}\left(\frac{r^{2}/2}{-c^{2}}\operatorname{Re}\left(\frac{1}{x+iy}\right)-\frac{nx}{2}\right)dx.$
By letting x=yux=-yu where $y>0$, we have
$\frac{|x+iy|^{k}}{(x+iy)^{k}}={\mathbf{e}}(-\tfrac{k}{4})\left(\frac{1-iu}{1+iu}\right)^{\frac{k}{2}}.$ (3.15)
Then we continue and apply [DJ13, Proposition 3.6] to get
[…]
Now we compute the Fourier expansions with the help of $f(\tau;c)$. By (3.10) and (2.3)
$F_{k}(\tau;r,s)$ $=-\sum_{2|c>0}\sum_{\begin{subarray}{c}d\ (\mathrm{mod}\ 2c)^{*}\\ -c<a<c\\ ad\equiv 1\ (\mathrm{mod}\ 2c)\end{subarray}}\frac{\nu_{\Theta}\begin{pmatrix}*&*\\ c&d\end{pmatrix}^{-2k}}{|2\pi r^{2}y|^{\frac{k}{2}}}$
$\cdot\sum_{t\in{\mathbb{Z}}}\frac{|c\tau+d+2t|^{k}}{(c\tau+d+2t)^{k}}M_{\frac{k}{2},s-\frac{1}{2}}\left(\frac{2\pi r^{2}y}{|c\tau+d+2t|^{2}}\right){\mathbf{e}}\left(\frac{r^{2}}{2}\left(\frac{a}{c}-\frac{1}{c(c\tau+d+2t)}\right)\right)$
$=-\sum_{2|c>0}\sum_{\begin{subarray}{c}d\ (\mathrm{mod}\ 2c)^{*}\\ -c<a<c\\ ad\equiv 1\ (\mathrm{mod}\ 2c)\end{subarray}}\frac{\nu_{\Theta}\begin{pmatrix}*&*\\ c&d\end{pmatrix}^{-2k}}{|2\pi r^{2}y|^{\frac{k}{2}}}{\mathbf{e}}\left(\frac{r^{2}a}{2c}\right)f\left(\tau+\frac{d}{c};c\right).$
The choice of $a$ in ${\mathbf{e}}(\frac{r^{2}a}{2c})$ is important and due to Proposition 3.2. The real-variable Kloosterman sums defined in (3.1) are therefore involved. We have
$f\left(\tau+\frac{d}{c};c\right)=\sum_{n\in{\mathbb{Z}}}f_{n}(y){\mathbf{e}}\left(\frac{nd}{2c}\right)e^{\pi inx}$
and (3.13) is proved by (3.16).
The proof of (3.14) follows in a similar way as the proof of (3.13), by recalling Proposition 3.2, noticing that every $\widetilde{\gamma}\in\widetilde{{\mathcal{B}}}$ has the form
$\widetilde{\gamma}=\left[\begin{pmatrix}b&-a\\ d&-c\end{pmatrix}\right]\quad\text{for }\left[\begin{pmatrix}a&b\\ c&d\end{pmatrix}\right]\in{\mathcal{B}}\text{ where we choose }d>0$
and d<b<d-d<b<d, as well as the Fourier expansion of $f(\tau-\frac{c}{d};d)$ with (3.16). ∎
The following two theorems play a key role in the proofs of Theorem 1.2 and Theorem 1.3. We prove them in Section 6 and Section 9, respectively.
Theorem 3.5.
Let $\tau=x+iy\in{\mathbb{H}}$ and $r\in{\mathbb{R}}\setminus\{0\}$. For $s\in[1,1.001]$, both $B_{4,n}(r;s,y)$ and $\widetilde{B}_{4,n}(r;s,y)$ are convergent and bounded by
$\begin{array}[]{r}|B_{4,n}(r;s,y)|\vskip 5.01874pt\\ |\widetilde{B}_{4,n}(r;s,y)|\end{array}\ll_{\varepsilon}\left\{\begin{array}[]{ll}n^{1+\varepsilon}e^{-\pi yn}&\text{if }n\geq 1,\ r^{2}n\leq 1,\\ (|r|^{-1}+|r|^{-\frac{1}{2}+\varepsilon})n^{\frac{3}{4}+\varepsilon}e^{-\pi yn}&\text{if }n\geq 1,\ r^{2}n\geq 1,\\ (|r|^{-1}+|r|^{1+\varepsilon})|n|^{2}e^{-\pi y|n|+2\pi|r||n|^{1/2}}/\Gamma(s-1)&\text{if }n\leq-1.\end{array}\right$.
The limits $\lim_{s\rightarrow 1^{+}}B_{4,0}(r;s,y)$ and $\lim_{s\rightarrow 1^{+}}\widetilde{B}_{4,0}(r;s,y)$ both exists. Therefore, both expansions of $F_{2}(\tau;r,s)$ (3.13) and $\widetilde{F}_{2}(\tau;r,s)$ (3.14) are absolutely convergent for $s\in[1,1.001]$.
Moreover, $F_{2}(\tau;r,1)$ and $\widetilde{F}_{2}(\tau;r,1)$ are explicitly given by
$F_{2}(\tau;r,1)=\frac{\sin(\pi r^{2})}{y\pi^{2}r^{2}}+\pi\sum_{n=1}^{\infty}e^{\pi in\tau}\left(\frac{n}{r^{2}}\right)^{\frac{1}{2}}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)$ (3.17)
and
[…]
They satisfy
$F_{2}(\tau;r,1)+(-i\tau)^{-2}\widetilde{F}_{2}(-1/\tau;r,1)=e^{\pi ir^{2}\tau}.$ (3.19)
Theorem 3.6.
Let $\tau=x+iy\in{\mathbb{H}}$, and $r\in{\mathbb{R}}\setminus\{0\}$. For $s\in[\frac{3}{4},1.001]$, both $|B_{3,n}(r;s,y)|$ and $|\widetilde{B}_{3,n}(r;s,y)|$ are convergent and bounded by
[…]
Then both $F_{\frac{3}{2}}(\tau;r,s)$ (3.13) and $\widetilde{F}_{\frac{3}{2}}(\tau;r,s)$ (3.14) are absolutely convergent for $s\in[\frac{3}{4},1.001]$. Explicitly, we have
$\begin{split}F_{\frac{3}{2}}(\tau;r,\tfrac{3}{4})&=\pi{\mathbf{e}}(\tfrac{1}{8})\sum_{n=1}^{\infty}e^{\pi in\tau}\left(\frac{n}{r^{2}}\right)^{\frac{1}{4}}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}J_{\frac{1}{2}}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ &+\frac{\sin(\pi r^{2})}{\pi r\sinh(\pi r)}\left(\sqrt{\frac{2}{y}}+2\pi^{\frac{1}{2}}\sum_{m=1}^{\infty}m\Gamma(-\tfrac{1}{2},2\pi m^{2}y)e^{-\pi im^{2}\tau}\right)\end{split}$ (3.20)
and
$\begin{split}\widetilde{F}_{\frac{3}{2}}(\tau;r,\tfrac{3}{4})&=\pi{\mathbf{e}}(-\tfrac{3}{8})\sum_{n=1}^{\infty}e^{\pi in\tau}\left(\frac{n}{r^{2}}\right)^{\frac{1}{4}}\sum_{2\nmid{\widetilde{c}}>0}\frac{\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})}{{\widetilde{c}}}J_{\frac{1}{2}}\left(\frac{2\pi|r|\sqrt{n}}{\widetilde{c}}\right)\\ &+\frac{\sin(\pi r^{2})}{\pi r\sinh(\pi r)}\left(\sqrt{\frac{2}{y}}+2\pi^{\frac{1}{2}}\sum_{m=1}^{\infty}m\Gamma(-\tfrac{1}{2},2\pi m^{2}y)e^{-\pi im^{2}\tau}\right).\end{split}$ (3.21)
They satisfy
$F_{\frac{3}{2}}(\tau;r,\tfrac{3}{4})+(-i\tau)^{-\frac{3}{2}}\widetilde{F}_{\frac{3}{2}}(-1/\tau;r,\tfrac{3}{4})=e^{\pi ir^{2}\tau}.$ (3.22)
4. Kloosterman sums in weight 2
In this section we assume $m,n,c\in{\mathbb{Z}}$ and $c>0$. Recall the notation from §2.1.
For the standard Kloosterman sum $S(m,n,c)$, by [ST09, (2)] we have the Weil bound
$|S(m,n,c)|\leq\sigma_{0}(c)\sqrt{\gcd(m,n,c)}\sqrt{c}.$ (4.1)
By Proposition 2.2, for positive integers $2|c$ and $2\nmid\widetilde{c}$ we have
[…]
When $mn=0$, the Kloosterman sums are Ramanujan sums. For $n\neq 0$,
$S(n,0,c)=S(0,n,c)=\sum_{d\ (\mathrm{mod}\ c)^{*}}{\mathbf{e}}\left(\frac{nd}{c}\right)=\sum_{d|(n,c)}d\mu\left(\frac{c}{d}\right).$
Then we have
$\sum_{c\leq x}\frac{S(0,n,c)}{c}=\sum_{\begin{subarray}{c}d|n\\ d\leq x\end{subarray}}\sum_{t\leq\frac{x}{d}}\frac{\mu(t)}{t}\ll\min\left(x,\sigma_{0}(n)\right)\ll_{\varepsilon}n^{\varepsilon}.$ (4.3)
The same bound holds for $\sum_{2|c}S(m,n,c,\nu_{\Theta}^{4})/c$ and $\sum_{2\nmid\widetilde{c}}S(m,n,\widetilde{c},\nu_{\Theta}^{4})/\widetilde{c}$. Moreover, we have $S(0,0,c)=\varphi(c)$ and
$\left.\begin{array}[]{rcl}\displaystyle\sum_{1\leq c\leq x}\frac{S(0,0,c)}{c}&=&\displaystyle\sum_{c\leq x}\frac{\varphi(c)}{c}\\ \displaystyle\sum_{2|c\leq x}\frac{S(0,0,c,\nu_{\Theta}^{4})}{c}&=&\displaystyle\sum_{2|c\leq x}\frac{\varphi(2c)}{c}\\ \displaystyle\sum_{2\nmid\widetilde{c}\leq x}\frac{S(0,0,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}&=&\displaystyle\sum_{2\nmid\widetilde{c}\leq x}\frac{-\varphi(\widetilde{c})}{\widetilde{c}}\end{array}\right\}=\left\{\begin{array}[]{l}\displaystyle\frac{6}{\pi^{2}}x\vskip 10.03749pt\\ \displaystyle\frac{4}{\pi^{2}}x\vskip 10.03749pt\\ \displaystyle-\frac{4}{\pi^{2}}x\end{array}\right.+O(\log x).$ (4.4)
5. Real-variable Kloosterman sums in weight 2
In this section we suppose $k=2$, $n\in{\mathbb{Z}}$ and $r\in{\mathbb{R}}_{+}$. Recall the real-variable Kloosterman sums $K(r,n,c,\nu_{\Theta}^{4})$ and $\widetilde{K}(r,n,c,\nu_{\Theta}^{4})$ from (3.1) and (3.2). Our goal is to estimate
$\sum_{2|c\leq x}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}\quad\text{and}\quad\sum_{2\nmid\widetilde{c}\leq x}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}$
and conclude the properties of the expansion in Lemma 3.4.
5.1. Case $n\neq 0$
We first prove the Weil bound for individual real-variable Kloosterman sums.
Lemma 5.1.
For $r\in{\mathbb{R}}\setminus{\mathbb{Z}}$ and $n\in{\mathbb{Z}}\setminus\{0\}$, we have
$K(r,n,c,\nu_{\Theta}^{4})\ll_{\varepsilon}(n,c)^{\frac{1}{2}+\varepsilon}c^{\frac{1}{2}+\varepsilon}\quad\text{and}\quad\widetilde{K}(r,n,\widetilde{c},\nu_{\Theta}^{4})\ll_{\varepsilon}(n,\widetilde{c})^{\frac{1}{2}+\varepsilon}\widetilde{c}^{\frac{1}{2}+\varepsilon}$
for $2|c>0$ and $2\nmid\widetilde{c}>0$.
Proof.
For $k\in{\mathbb{Z}}$, recall the Weil-type bound of Kloosterman sums (4.2) and $\sigma_{0}(n)\ll_{\varepsilon}n^{\varepsilon}$. Then by Lemma 2.8,
$K(r,n,c,\nu_{\Theta}^{4})$ $=\frac{i}{c}\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ |r-k|<c\end{subarray}}\frac{\sin(\pi(r-k))}{{\mathbf{e}}(\frac{r-k}{2c})-1}S(k,n,c,\nu_{\Theta}^{4}).$
Since
$|{\mathbf{e}}(t)-1|=\sqrt{2-2\cos(2\pi t)}\geq\pi|t|\quad\text{for }t\in[-\tfrac{1}{2},\tfrac{1}{2}],$ (5.1)
we get
$|K(r,n,c,\nu_{\Theta}^{4})|$ $\leq 2\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ r-c<k<r+c\end{subarray}}\left|\frac{\sin(\pi(r-k))}{\pi(r-k)}\right|\cdot|S(k,n,c,\nu_{\Theta}^{4})|$
$\leq 2\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ r-c<k\leq\left\lfloor r\right\rfloor-1\end{subarray}}\frac{|S(k,n,2c)|}{\pi(r-k)}+2\big|S(\left\lfloor r\right\rfloor,n,2c)\big|$
$+2\big|S(\left\lceil r\right\rceil,n,2c)\big|+2\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ \left\lceil r\right\rceil+1\leq k<r+c\end{subarray}}\frac{|S(k,n,2c)|}{\pi(k-r)}.$
By the fact that $\sigma_{s}(n)\ll_{s,\varepsilon}n^{s+\varepsilon}$ for $s\geq 0$, we reorder the summation and get the desired bound: the sum over $r-c<k\leq\left\lfloor r\right\rfloor-1$ is bounded in the following way:
$\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ r-c<k\leq\left\lfloor r\right\rfloor-1\end{subarray}}\frac{|S(k,n,2c)|}{r-k}$ $\ll_{\varepsilon}c^{\frac{1}{2}+\varepsilon}\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ r-c<k\leq\left\lfloor r\right\rfloor-1\end{subarray}}\frac{\sqrt{(k,n,c)}}{r-k}$
$\ll_{\varepsilon}c^{\frac{1}{2}+\varepsilon}\sum_{\delta|(n,c)}\delta^{\frac{1}{2}}\sum_{\begin{subarray}{c}t\in{\mathbb{Z}}\\ \frac{r-c}{\delta}<t\leq\frac{\left\lfloor r\right\rfloor-1}{\delta}\end{subarray}}\frac{1}{r-\delta t}$
$\ll_{\varepsilon}(n,c)^{\frac{1}{2}+\varepsilon}c^{\frac{1}{2}+\varepsilon};$
the bound over $\left\lceil r\right\rceil+1\leq k<r+c$ follows the same way.
The proof for $\widetilde{K}(r,n,\widetilde{c},\nu_{\Theta}^{4})$ follows from (2.20) and the same process. ∎
Proposition 5.2.
For $x\geq 1,r\in{\mathbb{R}}_{+}$ and $n\neq 0$, we have the following estimates:
$\sum_{2|c\leq x}\frac{|K(r^{2},n,c,\nu_{\Theta}^{4})|}{c}\ll_{\varepsilon}|n|^{\varepsilon}x^{\frac{1}{2}+\varepsilon},\quad\sum_{2\nmid\widetilde{c}\leq x}\frac{|K(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})|}{\widetilde{c}}\ll_{\varepsilon}|n|^{\varepsilon}x^{\frac{1}{2}+\varepsilon}.$
Proof.
The case $r^{2}\in{\mathbb{Z}}$ follows by directly applying the Weil bound of Kloosterman sums, and the case $r^{2}\notin{\mathbb{Z}}$ follows by directly applying Lemma 5.1. ∎
Proposition 5.3.
Let $s\in[1,1.001]$ and $r\in{\mathbb{R}}_{+}$. If $n>0$, both the following sums are absolute convergent and have the same estimate depending on $r^{2}n\leq 1$ or $r^{2}n\geq 1$:
$\text{both}\quad\left.\begin{array}[]{r}\displaystyle\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}>0}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}\left\{\begin{array}[]{ll}|r|n^{\frac{1}{2}+\varepsilon},&\text{if }r^{2}\leq\frac{1}{n},\vskip 5.01874pt\\ (1+|r|^{\frac{1}{2}+\varepsilon})n^{\frac{1}{4}+\varepsilon},&\text{if }r^{2}\geq\frac{1}{n}.\end{array}\right$.
If $n<0$, the following sums are absolute convergent and have estimates
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c>2\pi|r^{2}n|^{1/2}}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}>2\pi|r^{2}n|^{1/2}}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}(1+|r|^{1+\varepsilon})|n|^{\frac{1}{2}+\varepsilon}.$
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c\leq 2\pi|r^{2}n|^{1/2}}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}\leq 2\pi|r^{2}n|^{1/2}}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}(1+|r|^{2+\varepsilon})|n|^{1+\varepsilon}e^{2\pi|r||n|^{1/2}}.$
Proof.
The proof requires the following bounds on $J$- and $I$-Bessel functions. By [DLMF, (10.14.1)], $|J_{\nu}(x)|\leq 1$ for $\nu\geq 0$; by [DLMF, (10.7.3), (10.7.8)],
$J_{\nu}(x)\ll_{\nu}\min\left(x^{-\frac{1}{2}},x^{\nu}\right),\quad\text{hence }J_{2s-1}(x)\ll\min\left(x^{-\frac{1}{2}},x\right)\text{ for }s\in[1,1.001].$ (5.2)
By Lemma 5.1 and (5.2), we have the following two cases:
(1) when $r^{2}\leq\frac{1}{n}$, the following sum converges absolutely and has estimate
$\begin{split}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)&\ll_{\varepsilon}|r|n^{\frac{1}{2}}\sum_{2|c>0}(n,c)^{\frac{1}{2}+\varepsilon}c^{-\frac{3}{2}+\varepsilon}\\ &\ll_{\varepsilon}|r|n^{\frac{1}{2}+\varepsilon};\end{split}$ (5.3)
(2) when $r^{2}\geq\frac{1}{n}$, we break the sum at $|r^{2}n|^{1/2}$ and still get the absolute convergence
$\begin{split}&\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ &\ll_{\varepsilon}|r^{2}n|^{-\frac{1}{4}}\sum_{2|c,\,c\leq|r^{2}n|^{1/2}}(n,c)^{\frac{1}{2}+\varepsilon}c^{\varepsilon}+|r^{2}n|^{\frac{1}{2}}\sum_{2|c,\,c\geq|r^{2}n|^{1/2}}(n,c)^{\frac{1}{2}+\varepsilon}c^{-\frac{3}{2}+\varepsilon}\\ &\ll_{\varepsilon}|r^{2}n|^{\frac{1}{4}+\varepsilon}.\end{split}$ (5.4)
The case $2\nmid\widetilde{c}$ follows similarly.
For $n<0$ we need following bounds for the $I$-Bessel function: by [DLMF, (10.30.1)], for fixed $\nu\geq 0$, $I_{\nu}(x)\ll_{\nu}x^{\nu}$ for $x\leq 1$, hence $I_{2s-1}(x)\ll x$ for $x\leq 1$; by [DLMF, (10.30.4)], $I_{\nu}(x)\ll e^{x}$ for $x\geq 1$.
Those series are absolute convergent following from Lemma 5.1, Proposition 5.2 and the above inequalities. The proof for case $2\nmid\widetilde{c}$ follows by the same argument. We have finished the proof of the proposition. ∎
5.2. Case $n=0$
For $2|c>0$, we define and compute
$\begin{split}\varphi_{c}(r)&\vcentcolon=\sum_{\begin{subarray}{c}1\leq a<c\\ (a,c)=1\end{subarray}}e^{\pi ira/c}=\sum_{1\leq a\leq c}\sum_{\delta|(a,c)}\mu(\delta)e^{\pi ira/c}=\sum_{\delta|c}\mu(\delta)\sum_{\begin{subarray}{c}a=m\delta\\ 1\leq m\leq\frac{c}{\delta}\end{subarray}}e^{\pi irm\delta/c}\\ &=\sum_{\delta|c}\mu(\delta)\frac{e^{\pi ir}-1}{1-e^{-\pi ir\delta/c}}.\end{split}$ (5.5)
Recall the real-variable Kloosterman sum defined in (3.1). We have
$2\operatorname{Re}\varphi_{c}(r^{2})=K(r^{2},0,c,\nu_{\Theta}^{4}).$
Let $Z(s)$ denote the zeta function
$Z(s)\vcentcolon=\sum_{2|c>0}\frac{\varphi_{c}(r^{2})}{c^{2s}},\quad\operatorname{Re}s>1.$ (5.6)
We want to understand the behavior of $Z(s)$ when $s\rightarrow 1^{+}$. By (5.5), for $\operatorname{Re}s>1$ we have
$Z(s)$ $=(e^{\pi ir^{2}}-1)\sum_{2|c>0}\sum_{\delta|c}\frac{\mu(\delta)}{1-e^{-\pi ir^{2}\delta/c}}c^{-2s}$
$(\delta\text{ odd, }c=2m\delta)$ $=(e^{\pi ir^{2}}-1)\sum_{2\nmid\delta>0}\frac{\mu(\delta)}{(2\delta)^{2s}}\sum_{m=1}^{\infty}\frac{m^{-2s}}{(1-{\mathbf{e}}\left(\frac{-r^{2}}{4m}\right))}$
$(\delta\text{ even, }c=m\delta)$ $+(e^{\pi ir^{2}}-1)\sum_{2|\delta>0}\frac{\mu(\delta)}{\delta^{2s}}\sum_{m=1}^{\infty}\frac{m^{-2s}}{(1-{\mathbf{e}}\left(\frac{-r^{2}}{2m}\right))}$
$=:\Sigma_{1}+\Sigma_{2}.$
To apply Lemma 2.9 we require $m\geq r^{2}$, which will only affect finitely many of $m$. Then for $\Sigma_{1}$ we have
$\sum_{m=1}^{\infty}\frac{m^{-2s}}{(1-{\mathbf{e}}\left(\frac{-r^{2}}{4m}\right))}$ $=O_{r}(1)+\sum_{m>r^{2}}\frac{m^{-2s}}{(1-{\mathbf{e}}\left(\frac{-r^{2}}{4m}\right))}$
$=O_{r}(1)+\sum_{m>r^{2}}\frac{m^{-2s+1}}{\pi ir^{2}/2}-\sum_{\ell=1}^{\infty}B_{\ell}\frac{(-2\pi ir^{2})^{\ell-1}}{\ell!4^{\ell-1}}\sum_{m>r^{2}}\frac{1}{m^{2s+\ell-1}}$
$=\sum_{m>r^{2}}\frac{m^{-2s+1}}{\pi ir^{2}/2}+O_{r}(1),$
where the last step is the result of
$\sum_{m>r^{2}}\frac{1}{m^{2s+\ell-1}}=O((r^{2})^{-\ell})\quad\text{for }\operatorname{Re}s\geq 1,\quad\text{when }\ell\geq 1.$
We have similar estimates for $\Sigma_{2}$. Then the main contribution to $Z(s)$ is given by
$Z(s)$ $=\frac{e^{\pi ir^{2}}-1}{\pi ir^{2}/2}\sum_{2\nmid\delta>0}\frac{\mu(\delta)}{(2\delta)^{2s}}\zeta(2s-1)$
$+\frac{e^{\pi ir^{2}}-1}{\pi ir^{2}}\sum_{2|\delta>0}\frac{\mu(\delta)}{\delta^{2s}}\zeta(2s-1)+O_{r}(1).$
Since
$\lim_{s\rightarrow 1^{+}}\frac{\zeta(2s-1)}{\Gamma(s-1)}=\frac{1}{2}\quad\text{and}\quad\sum_{2\nmid\delta>0}\frac{\mu(\delta)}{\delta^{2}}=\frac{8}{\pi^{2}},$
we have
$\lim_{s\rightarrow 1^{+}}\frac{1}{\Gamma(s-1)}\sum_{2|c>0}\frac{K(r^{2},0,c,\nu_{\Theta}^{4})}{c^{2s}}=\lim_{s\rightarrow 1^{+}}\frac{2\operatorname{Re}Z(s)}{\Gamma(s-1)}=\frac{2\sin(\pi r^{2})}{\pi^{3}r^{2}}.$ (5.7)
The similar process applies to $\widetilde{K}(r^{2},0,d,\nu_{\Theta}^{4})$. We have $\widetilde{K}(r^{2},0,1,\nu_{\Theta}^{4})=-1$ and
$\widetilde{K}(r^{2},0,d,\nu_{\Theta}^{4})=-\sum_{\begin{subarray}{c}-d<b<d\\ 2|b,\ (b,d)=1\end{subarray}}e^{\pi ir^{2}b/d}=(e^{\pi ir^{2}}-e^{-\pi ir^{2}})\sum_{\delta|2d}\frac{\mu(\delta)}{{\mathbf{e}}(-\frac{r\delta}{2d})-1}$
for odd $d>1$. Finally, we find
$\lim_{s\rightarrow 1^{+}}\frac{1}{\Gamma(s-1)}\sum_{2\nmid d>0}\frac{\widetilde{K}(r^{2},0,d,\nu_{\Theta}^{4})}{d^{2s}}=\frac{2\sin(\pi r^{2})}{\pi^{3}r^{2}}.$ (5.8)
6. Proof of Theorem 3.5 and Theorem 1.2
We are now ready to prove Theorem 3.5.
Proof of Theorem 3.5.
We first cite the following bounds of Whittaker function [DLMF, (13.14.21)]: for $n\neq 0$,
$\frac{|W_{\pm\frac{k}{2},s-\frac{1}{2}}(2\pi|n|y)|}{|2\pi r^{2}y|^{\frac{k}{2}}}\left|\frac{r^{2}}{n}\right|^{\frac{1}{2}}\ll e^{-\pi|n|y}\left|\frac{n}{r^{2}}\right|^{\frac{k-1}{2}}$ (6.1)
Combining Proposition 5.3 and §5.2, for $r\in{\mathbb{R}}_{+}$ and $n\in{\mathbb{Z}}$, we conclude that for $s\in[1,1.001]$, $B_{4,n}(r;y,s)$ and $\widetilde{B}_{4,n}(r;y,s)$ are absolute convergent and have the claimed bound.
By dominated convergence theorem, the factor $e^{-\pi y|n|}$ ensures that the Fourier expansions of $F_{2}(\tau,r;s)$ and $\widetilde{F}_{2}(\tau,r;s)$ for $s\in[1,1.001]$ in Lemma 3.4 are both absolutely convergent. By analytic continuation, $F_{2}(\tau,r;1)$ and $\widetilde{F}_{2}(\tau,r;1)$ has the corresponding Fourier expansion given by $B_{4,n}(r;1,y)$ and $\widetilde{B}_{4,n}(r;1,y)$. The explicit formula in (3.17) and (3.18) follows from (3.6), (5.7) and (5.8). Note that when $n\leq-1$, $B_{4,n}(r;y,s)$ and $\widetilde{B}_{4,n}(r;y,s)$ are absolute convergent and converge to 0 when $s\rightarrow 1^{+}$ because of the Gamma factor.
The functional equation (3.19) follows from Proposition 3.3, analytic continuation, and (3.12). ∎
To remove the $n=0$ term $\frac{\sin(\pi r^{2})}{y\pi^{2}r^{2}}$ in Theorem 3.5 such that the resulting functions fulfill the conditions of Theorem 1.1, we need to use the weight $2$ Eisenstein series.
Definition 6.1 ([DS05, (1.5)]).
For $\tau=x+iy\in{\mathbb{H}}$ and $x,y\in{\mathbb{R}}$, the weight $2$ Eisenstein series $E_{2}(\tau)$ is defined by
$E_{2}(\tau)=1-24\sum_{n=1}^{\infty}\sigma_{1}(n)e^{2\pi in\tau},\quad\text{where }\sigma_{1}(n)=\sum_{d|n}d.$
The weight 2 Eisenstein series is not modular, but it satisfies the following identity
$E_{2}(\tau)-\tau^{-2}E_{2}(-1/\tau)=-\frac{12}{2\pi i\tau}.$ (6.2)
Let $\sigma_{1}(x)=0$ for $x\notin{\mathbb{Z}}$ and define
$\mathcal{E}_{2}(\tau)\vcentcolon=-\frac{3}{\pi y}+E_{2}\left(\frac{\tau+1}{2}\right)-E_{2}(\tau)=-\frac{3}{\pi y}+24\sum_{n=1}^{\infty}\left(\sigma_{1}\left(\frac{n}{2}\right)-(-1)^{n}\sigma_{1}(n)\right)e^{\pi in\tau}.$ (6.3)
It is straightforward by (6.2) that
$\mathcal{E}_{2}(\tau+2)=\mathcal{E}_{2}(\tau)\quad\text{and}\quad\mathcal{E}_{2}(\tau)+(\tau/i)^{-2}\mathcal{E}_{2}(-1/\tau)=0.$
Therefore, by defining the series
$\begin{split}G_{2}(\tau;r)&\vcentcolon=F_{2}(\tau;r,1)+\frac{\sin(\pi r^{2})}{6\pi r^{2}}\mathcal{E}_{2}(\tau)\\ &=\sum_{n=1}^{\infty}e^{\pi in\tau}\cdot\pi\left(\frac{n}{r^{2}}\right)^{\frac{1}{2}}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ &+\sum_{n=1}^{\infty}e^{\pi in\tau}\cdot\frac{8\sin(\pi r^{2})}{\pi r^{2}}\left(\sigma_{1}\left(\frac{n}{2}\right)-(-1)^{n}\sigma_{1}(n)\right),\end{split}$ (6.4)
and
$\begin{split}\widetilde{G}_{2}(\tau;r)&\vcentcolon=\widetilde{F}_{2}(\tau;r,1)+\frac{\sin(\pi r^{2})}{6\pi r^{2}}\left(E_{2}^{*}(\tfrac{\tau+1}{2})-\Theta(\tau)^{4}\right)\\ &=\sum_{n=1}^{\infty}e^{\pi in\tau}\cdot(-\pi)\left(\frac{n}{r^{2}}\right)^{\frac{1}{2}}\sum_{2\nmid{\widetilde{c}}>0}\frac{\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{4})}{\widetilde{c}}J_{1}\left(\frac{2\pi|r|\sqrt{n}}{{\widetilde{c}}}\right)\\ &+\sum_{n=1}^{\infty}e^{\pi in\tau}\cdot\frac{8\sin(\pi r^{2})}{\pi r^{2}}\left(\sigma_{1}\left(\frac{n}{2}\right)-(-1)^{n}\sigma_{1}(n)\right),\end{split}$ (6.5)
we get that they satisfy
$G_{2}(\tau;r)+(\tau/i)^{-2}\widetilde{G}_{2}(\tau;r)=e^{\pi ir^{2}\tau}.$ (6.6)
Proof of Theorem 1.2.
The coefficients $b_{4,n}(r)$ and $\widetilde{b}_{4,n}(r)$ can be read from (6.4) and (6.5). The bound for them in $n$ is clear by Proposition 5.3. It remains to determine the values
$B_{4,n}(0)=\lim_{r\rightarrow 0^{+}}B_{4,n}(r)\quad\text{and}\quad\widetilde{B}_{4,n}(0)=\lim_{r\rightarrow 0^{+}}\widetilde{B}_{4,n}(r)\quad\text{for }n\geq 1.$ (6.7)
By [DLMF, (10.2.2)]:
$\frac{\pi\sqrt{n}}{|r|}J_{1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)=\frac{\pi^{2}n}{c}\left(1+\sum_{j=1}^{\infty}\frac{(-1)^{j}\left(\pi|r|\sqrt{n}/c\right)^{2j}}{j!(j+1)!}\right),$ (6.8)
we claim that the limits in (6.7) are given by
$\begin{split}B_{4,n}(0)=\pi^{2}n\sum_{2|c>0}\frac{S(0,n,c,\nu_{\Theta}^{4})}{c^{2}}\quad\text{and}\quad\widetilde{B}_{4,n}(0)=\pi^{2}n\sum_{2\nmid\widetilde{c}>0}\frac{S(0,n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}^{2}}.\end{split}$ (6.9)
This can be proved by absolute convergence with Proposition 5.2.
To compute $B_{4,n}(0)$ and $\widetilde{B}_{4,n}(0)$ explicitly, we define the Ramanujan sum
$R_{c}(n)\vcentcolon=\sum_{d\ (\mathrm{mod}\ c)}{\mathbf{e}}\left(\frac{nd}{c}\right).$
It is well-known that $R_{c}(n)$ is a multiplicative function of $c$,
$\sum_{c=1}^{\infty}\frac{R_{c}(n)}{c^{s}}=\frac{\sigma_{s-1}(|n|)}{|n|^{s-1}\zeta(s)},\quad\text{and}\quad R_{p^{j}}(n)=\left\{\begin{array}[]{ll}0,&j\geq\nu_{p}(n)+2,\\ -p^{j-1},&j=\nu_{p}(n)+1,\\ \varphi(p^{j}),&j\leq\nu_{p}(n).\end{array}\right$. (6.10)
By defining $h=\nu_{2}(n)$ and computing the $2$-factor of the $L$-function, we conclude that
$B_{4,n}(0)$ $=\frac{\pi^{2}n}{4}\sum_{c=1}^{\infty}\frac{R_{4c}(n)}{c^{2}}=8\sigma_{1}(n)\cdot\left\{\begin{array}[]{ll}\displaystyle\frac{2^{h}-3}{2^{h+1}-1},&2|n,\vskip 2.00749pt\\ 0,&2\nmid n,\end{array}\right$.
$\widetilde{B}_{4,n}(0)$ $=\pi^{2}n\sum_{2\nmid\widetilde{c}>0}\frac{R_{\widetilde{c}}(n)}{\widetilde{c}^{2}}=8\sigma_{1}(n)\cdot\frac{2^{h}}{2^{h+1}-1}.$
This concludes the proof. ∎
7. Kloosterman sums in weight $3/2$
Recall the notation from §2. In this section we prove the properties of $S(m,n,c,\nu_{\Theta}^{3})$ for $2|c>0$ and of $S(m,n,\widetilde{c},\nu_{\Theta}^{3})$ for $2\nmid\widetilde{c}>0$, where $m,n\in{\mathbb{Z}}$.
7.1. Case $mn\neq 0$
For half-integer weight $k\in{\mathbb{Z}}+\frac{1}{2}$, for Kloosterman sums with multiplier $\nu_{\theta}$ we have the Weil bound
$|S(m,n,c,\nu_{\theta}^{2k})|\leq\sigma_{0}(c)\sqrt{\gcd(m,n,c)}\sqrt{c}$ (7.1)
by [Blo08, (2.15)]. According to the relation (2.14), we get the Weil bound for $\nu_{\Theta}$ when $c\in 2{\mathbb{Z}}_{+}$:
$|S(m,n,c,\nu_{\Theta}^{2k})|=|S(m,n,2c,\nu_{\theta}^{2k})|\leq 2\sigma_{0}(2c)\sqrt{\gcd(m,n,c)}\sqrt{c}.$ (7.2)
Proposition 7.1.
For $mn\neq 0$, we have the following estimate:
[…]
Here $A_{u}(m,n)$ is defined as in [Sun24b, Theorem 1.4] and [Sun25, Theorem 1.3]: writing $m=t_{m}u_{m}^{2}w_{m}^{2}$ and $n=t_{n}u_{n}^{2}w_{n}^{2}$, where $t_{m},t_{n}$ are square-free, $u_{m},u_{n}$ are even, $w_{m},w_{n}$ are odd, and $\delta=\frac{1}{147}$, we define
$A_{u}(m,n)\vcentcolon=\left(|m|^{\frac{1}{2}-8\delta}+u_{m}\right)^{\frac{1}{8}}\left(|n|^{\frac{1}{2}-8\delta}+u_{n}\right)^{\frac{1}{8}}|mn|^{\frac{3}{16}}.$ (7.4)
We have $A_{u}(m,n)\ll|mn|^{\frac{1}{4}}$, and if $2^{\nu_{2}(m)}$ and $2^{\nu_{2}(n)}$ are bounded, $A_{u}(m,n)\ll|mn|^{\frac{1}{4}-\delta}$.
Remark.
In other words, $u_{n}=2^{\left\lfloor\nu_{2}(n)/2\right\rfloor}$.
Proof.
By (2.14), we have
$\begin{split}\sum_{2|c\leq x}\frac{S(m,n,c,\nu_{\Theta}^{3})}{c}&=\sum_{4|2c\leq 2x}\frac{S(m,n,2c,\nu_{\theta}^{3})}{c}\\ \quad&=2\sum_{4|c\leq 2x}\frac{S(m,n,c,\nu_{\theta}^{3})}{c}=2\overline{\sum_{4|c\leq 2x}\frac{S(-m,-n,c,\nu_{\theta})}{c}}.\end{split}$ (7.5)
So it suffices to prove the asymptotics for the sum of $S(m,n,c,\nu_{\theta}^{3})$ and $S(m,n,c,\nu_{\theta})$. Note that $\nu_{\theta}^{3}=\nu_{\theta}^{-1}$ is a weight $\frac{1}{2}$ multiplier system on $\Gamma_{0}(4)$.
Recall §2.3 for the theory of Maass forms. We apply [Sun24b, Theorem 1.4] (respectively [Sun25, Theorem 1.3]) to estimate when $m>0$ and $n>0$ (respectively $m>0$ and $n<0$). For $k=\pm\frac{1}{2}$, the multiplier $\nu_{\theta}$ is on $\Gamma_{0}(4)$ and satisfies [Sun24b, Sun25, Definition 1.1] by (7.1) and taking $D=4$. Therefore, for $k=\pm\frac{1}{2}$,
$\sum_{4|c\leq x}\frac{S(m,n,c,\nu_{\theta}^{2k})}{c}-\sum_{r_{j}\in i(0,\frac{1}{4}]}\tau_{j}^{\pm}(m,n)\frac{x^{2s_{j}-1}}{2s_{j}-1}\ll_{\varepsilon}\left(A_{u}(m,n)+x^{\frac{1}{6}}\right)|xmn|^{\varepsilon}$
where $s_{j}=\operatorname{Im}r_{j}+\frac{1}{2}$ runs over eigenvalues $\lambda_{j}=\frac{1}{4}+r_{j}^{2}<\frac{1}{4}$ of weight $k$ hyperbolic Laplacian on $\Gamma_{0}(4)$. Here we write $\tau_{j}^{\pm}$ to mention that $\tau_{j}(m,n)$ depends on the weight $k=\pm\frac{1}{2}$. By Lemma 2.7, the Shimura correspondence of eigenvalues of $\Delta_{0}$ and $\Delta_{\frac{1}{2}}$ ([Sun24b, Theorem 5.6], [Sar84, p. 304]), and (2.25), we only need to consider $\tau_{0}^{\pm}(m,n)$, where $r_{0}=\frac{i}{4}$ and $s_{0}=\frac{3}{4}$.
By [GS83] (corrected in [AA16, (6.3)]), or by combining [Sun24b, Theorem 1.4] and [Sun25, Theorem 1.3], when $m>0$ and $k=\pm\frac{1}{2}$, the main contribution to the estimate above is
$2\tau_{0}^{\pm}(m,n)x^{\frac{1}{2}}=4\sqrt{2}\,{\mathbf{e}}(\tfrac{k}{4})\overline{\rho_{0}(m)}\rho_{0}(n)|mn|^{\frac{1}{4}}\frac{\Gamma(\frac{3}{4}+\operatorname{sgn}(n)\frac{k}{2})}{\Gamma(\frac{3}{4}-\frac{k}{2})}x^{\frac{1}{2}},$
where $\rho_{0}(n)$ for $n\in{\mathbb{Z}}\setminus\{0\}$ is given by (2.28). We get
$2\tau_{0}^{+}(m,n)=\left\{\begin{array}[]{ll}{\mathbf{e}}(\tfrac{1}{8})\frac{4\sqrt{2}}{\pi^{2}},&\ \text{if }m,n\text{ are both positive squares;,}\\ 0,&\ \text{other }m>0,\ n\neq 0,\end{array}\right$. (7.6)
and
$2\tau_{0}^{-}(m,n)=0\quad\text{if }m>0,\ n\neq 0.$ (7.7)
Thus, the estimate of (7.3) is given by (7.5) and the following three cases: (1) when $m>0$, (7.7) suggests the coefficient for $x^{\frac{1}{2}}$ as 0; (2) when $m<0$, (7.6) suggests main term consisting of $x^{\frac{1}{2}}$ as
$2{\mathbf{e}}(-\tfrac{1}{8})\frac{4\sqrt{2}}{\pi^{2}}(2x)^{\frac{1}{2}}={\mathbf{e}}(-\tfrac{1}{8})\frac{16}{\pi^{2}}x^{\frac{1}{2}}\quad\text{when }m\text{ and }n\text{ are negative squares};$
(3) in the other cases when $m<0$, (7.6) and (7.7) combined show that the coefficient of $x^{\frac{1}{2}}$ is 0. We have finished the proof. ∎
A similar method applies to the sum over $4|c$ by Lemma 2.6, hence we state the following proposition without proof.
Proposition 7.2.
For $mn\neq 0$, we have
$\begin{split}\sum_{4|c\leq x}\frac{S(m,n,c,\nu_{\Theta}^{3})}{c}&-\left\{\begin{array}[]{ll}{\mathbf{e}}(-\frac{1}{8})\frac{8}{\pi^{2}}x^{\frac{1}{2}}&\text{both }m\text{ and }n\text{ are negative squares}\\ 0&\text{other }mn\neq 0\end{array}\right.\\ &\ll_{\varepsilon}\left(x^{\frac{1}{6}}+A_{u}(m,n)\right)|xmn|^{\varepsilon},\end{split}$ (7.8)
where $A_{u}(m,n)$ is defined at (7.4).
We also have the following proposition for the odd case.
Proposition 7.3.
For $mn\neq 0$, we have the following estimate:
$\begin{split}\sum_{2\nmid\widetilde{c}\leq x}\frac{S(m,n,\widetilde{c},\nu_{\Theta}^{3})}{\widetilde{c}}&+\left\{\begin{array}[]{ll}{\mathbf{e}}(-\frac{1}{8})\frac{16}{\pi^{2}}x^{\frac{1}{2}}&\text{both }m\text{ and }n\text{ are negative squares}\\ 0&\text{other }mn\neq 0\end{array}\right.\\ &\ll_{\varepsilon}\left(x^{\frac{1}{6}}+A_{u}(m,n)\right)|xmn|^{\varepsilon}.\end{split}$ (7.9)
Where $A_{u}(m,n)$ is defined as in (7.4).
Proof.
Define $\xi\in\{\pm 1\}$ by $\xi=-1$ if $m\equiv 0,3\ (\mathrm{mod}\ 4)$ and $\xi=+1$ if $m\equiv 1,2\ (\mathrm{mod}\ 4)$. By (2.19) we have
$\begin{split}\sum_{2\nmid\widetilde{c}\leq x}\frac{S(m,n,\widetilde{c},\nu_{\Theta}^{3})}{\widetilde{c}}&=\frac{\xi}{\sqrt{2}}\sum_{2\|c\leq 2x}\frac{S(m,4n,c,\nu_{\Theta}^{3})}{c/2}\\ &=\xi\sqrt{2}\left(\sum_{2|c\leq 2x}-\sum_{4|c\leq 2x}\right)\frac{S(m,4n,c,\nu_{\Theta}^{3})}{c/2}.\end{split}$ (7.10)
The main term involving $x^{\frac{1}{2}}$ is then given by Proposition 7.1 and Proposition 7.2. Note that when $m$ is a negative square, we always have $m\equiv 0,3\ (\mathrm{mod}\ 4)$ and hence $\xi=-1$. ∎
7.2. Case $mn=0$
This case does not satisfy the conditions of [Sun24b, Theorem 1.4] and [Sun25, Theorem 1.3], so we deal with it separately in this subsection.
Recall our notations in §2: $\varepsilon_{d}$ at (2.8); for any prime $p$ and $n\in{\mathbb{Z}}\setminus\{0\}$, let $p^{\nu}\|n$ denote that $p^{\nu}|n$ while $p^{\nu+1}\nmid n$, and we write $\nu_{p}(n)=\nu$ for the integer $\nu$. First we need the following property: for positive odd integers $a,b$,
$\left(\frac{a}{b}\right)=\left(\frac{b}{a}\right)\varepsilon_{a}\varepsilon_{b}\varepsilon_{ab}^{-1}=\left(\frac{b}{a}\right)\varepsilon_{a}^{-1}\varepsilon_{b}^{-1}\varepsilon_{ab}.$ (7.11)
The following construction is from [WP12, Chapter 1] and can be traced back to [Maa37, Bat51]. We record the explicit construction and results here for our application. For $n\in{\mathbb{Z}}$, $k\in{\mathbb{Z}}+\frac{1}{2}$ and $p$ as a prime, we define
$\alpha_{2k}(2^{\nu},n)$ $\vcentcolon=\sum_{a=1}^{2^{\nu}}\left(\frac{2^{\nu}}{a}\right)\varepsilon_{a}^{2k}{\mathbf{e}}\left(\frac{na}{2^{\nu}}\right),\quad\nu\geq 2;$ (7.12)
$\alpha_{2k}(p^{\nu},n)$ $\vcentcolon=\varepsilon_{p^{\nu}}^{-2k}\sum_{a=1}^{p^{\nu}}\left(\frac{a}{p^{\nu}}\right){\mathbf{e}}\left(\frac{na}{p^{\nu}}\right),\quad\nu\geq 0,\ p>2.$ (7.13)
Then for $2|c>0$, the condition $ad\equiv 1\ (\mathrm{mod}\ 2c)$ implies $(\frac{2c}{d})=(\frac{2c}{a})$ and $\varepsilon_{d}=\varepsilon_{a}$, hence we have
$S(0,n,c,\nu_{\Theta}^{2k})=S(n,0,c,\nu_{\Theta}^{2k})=\sum_{a\ (\mathrm{mod}\ 2c)^{*}}\varepsilon_{a}^{2k}\left(\frac{2c}{a}\right){\mathbf{e}}\left(\frac{na}{2c}\right)=\prod_{p^{\nu}\|2c}\alpha_{2k}(p^{\nu},n).$ (7.14)
[…]
$\begin{split}S(0,n,\widetilde{c},\nu_{\Theta}^{2k})&=S(n,0,\widetilde{c},\nu_{\Theta}^{2k})={\mathbf{e}}(\tfrac{k}{4})\sum_{\begin{subarray}{c}a\ (\mathrm{mod}\ 2\widetilde{c})\\ 2|a,\,(a,\widetilde{c})=1\end{subarray}}\varepsilon_{\widetilde{c}}^{-2k}\left(\frac{2a}{\widetilde{c}}\right){\mathbf{e}}\left(\frac{na}{2\widetilde{c}}\right)\\ &={\mathbf{e}}(\tfrac{k}{4})\varepsilon_{\widetilde{c}}^{-2k}\sum_{a\ (\mathrm{mod}\ \widetilde{c})^{*}}\left(\frac{a}{\widetilde{c}}\right){\mathbf{e}}\left(\frac{na}{\widetilde{c}}\right)={\mathbf{e}}(\tfrac{k}{4})\prod_{p^{\nu}\|\widetilde{c}}\alpha_{2k}(p^{\nu},n).\end{split}$ (7.15)
For $p$ as a prime, we also define
$A_{2k}(2,n,s)\vcentcolon=\sum_{\nu=2}^{\infty}\frac{\alpha_{2k}(2^{\nu},n)}{2^{2s\nu}}\quad\text{and}\quad A_{2k}(p,n,s)\vcentcolon=\sum_{\nu=0}^{\infty}\frac{\alpha_{2k}(p^{\nu},n)}{p^{2s\nu}}\quad\text{for }p>2.$
Then (7.14) and (7.15) imply
$\sum_{2|c>0}\frac{S(n,0,c,\nu_{\Theta}^{2k})}{(2c)^{2s}}=\prod_{p}A_{2k}(p,n,s)$ (7.16)
and
$\sum_{2\nmid\widetilde{c}>0}\frac{S(n,0,\widetilde{c},\nu_{\Theta}^{2k})}{\widetilde{c}^{2s}}=\prod_{p>2}A_{2k}(p,n,s).$ (7.17)
In the following two subsections we compute $\alpha_{2k}$ and $A_{2k}$ in the cases $p=2$ or $p>2$, respectively.
7.2.1. $p=2$
It is direct to calculate the results for $\nu=0,1$. For $\nu\geq 2$ we divide into cases for $\nu$ is even or odd.
If $\nu\geq 2$ is even, then for odd $d$, $1\leq d<2^{\nu}$, we can write $d=4d^{\prime}+1$ or $d=4d^{\prime}+3$ for $d^{\prime}$ from $1$ to $2^{\nu-2}$. Since $2|\nu$, $(\frac{2^{\nu}}{d})=1$ and we have
[…]
If $\nu\geq 3$ is odd, then $(\frac{2^{\nu}}{d})=(\frac{2}{d})$ and we get
[…]
Here $\mathbf{1}_{{\mathbb{Z}}}(x)=1$ if $x\in{\mathbb{Z}}$ and 0 otherwise.
According to [WP12, Chapter 2], or by direct calculation, we have the following lemma. The facts that ${\mathbf{e}}(\frac{u-1}{8})=(\frac{u}{2})$ for $u\equiv 1\ (\mathrm{mod}\ 4)$ and ${\mathbf{e}}(\frac{u+1}{8})=(\frac{-u}{2})$ for $u\equiv 3\ (\mathrm{mod}\ 4)$ are helpful.
Lemma 7.4.
For $k=\lambda+\frac{1}{2}=\frac{1}{2}$ or $\frac{3}{2}$, let $h=\nu_{2}(n)$ and we have
$A_{2k}(2,n,s)=2^{-4s}(1+i^{2k})\times$
[…]
Particularly, when $n=m^{2}\geq 1$ is a square, we can write $m_{o}$ as the odd part of $m$. We get $2|h$, $\frac{n}{2^{h}}=-m_{o}^{2}\equiv 3\ (\mathrm{mod}\ 4)$, $(\frac{-n/2^{h}}{2})=(\frac{m_{o}^{2}}{2})=1$,
[…]
and
$\left.\frac{A_{3}(2,-m^{2},s)}{1+2^{-(2s-\frac{1}{2})}}\right|_{s=\frac{3}{4}}=\frac{2^{-3}(1-i)(2-2^{-h/2}+(\frac{m_{o}^{2}}{2})2^{-h/2})}{3/2}=\frac{{\mathbf{e}}(-\tfrac{1}{8})}{3\sqrt{2}}.$ (7.21)
When $n\neq 0$ is not a negative square, we still have the rough bound
$0\leq|A_{3}(2,n,s)|\leq\frac{\sqrt{2}}{4}\left(\frac{\nu_{2}(n)}{2}+1\right)\quad\text{ for }\operatorname{Re}s\in[\tfrac{1}{2},1].$ (7.22)
7.2.2. $p>2$
For $\nu=0$ we have $\alpha_{2k}(1,n)=1$. For $\nu\geq 1$, we write $a=a^{\prime}+pb$ for $a^{\prime}$ from $1$ to p1p-1 and $b$ from $1$ to $p^{\nu-1}$. Then $(\frac{a}{p})=(\frac{a^{\prime}}{p})$ and we have
$\alpha_{2k}(p^{\nu},n)=\varepsilon_{p^{\nu}}^{-2k}\sum_{a^{\prime}=1}^{p-1}\left(\frac{a^{\prime}}{p^{\nu}}\right){\mathbf{e}}\left(\frac{na^{\prime}}{p^{\nu}}\right)\sum_{b=1}^{p^{\nu-1}}{\mathbf{e}}\left(\frac{nb}{p^{\nu-1}}\right).$
If $\nu-1>\nu_{p}(n)$, i.e. $\nu\geq\nu_{p}(n)+2$, the latter sum on $b$ is zero, otherwise the latter sum is $p^{\nu-1}$. The remaining calculation is under the assumption $\nu\leq\nu_{p}(n)+1$.
If $\nu$ is odd, then
$\alpha_{2k}(p^{\nu},n)$ $=\varepsilon_{p^{\nu}}^{-2k}p^{\nu-1}\sum_{a=1}^{p-1}\left(\frac{a}{p}\right){\mathbf{e}}\left(\frac{an/p^{\nu-1}}{p}\right)$
$=\left\{\begin{array}[]{ll}0,&\ p|\frac{n}{p^{\nu-1}},\text{ i.e. }\nu\leq\nu_{p}(n);\\ p^{\nu-\frac{1}{2}}\left(\frac{(-1)^{k-1/2}\,n/p^{\nu-1}}{p}\right),&\ \nu=\nu_{p}(n)+1.\end{array}\right$.
Here in the last case we used $\varepsilon_{p^{\nu}}=\varepsilon_{p}$ for odd $\nu$ and $\varepsilon_{p}^{2}=(\frac{-1}{p})$.
If $\nu$ is even, then $\varepsilon_{p^{\nu}}=1$ and $(\frac{a^{\prime}}{p^{\nu}})=1$ for $\gcd(a^{\prime},p)=1$. We have
$\alpha_{2k}(p^{\nu},n)=p^{\nu-1}\sum_{a=1}^{p-1}{\mathbf{e}}\left(\frac{an/p^{\nu-1}}{p}\right)=p^{\nu-1}\cdot\left\{\begin{array}[]{ll}p-1,&\ \nu\leq\nu_{p}(n);\\ -1,&\ \nu=\nu_{p}(n)+1.\end{array}\right$.
Finally, we conclude
[…]
Then we have the following lemma:
Lemma 7.5.
For prime $p>2$, $k\in{\mathbb{Z}}+\frac{1}{2}$, $n\neq 0$ and $h\vcentcolon=\nu_{p}(n)$, we have
$A_{2k}(p,n,s)=1+\left\{\begin{array}[]{ll}\displaystyle\frac{p^{2(1-2s)}-p^{(h+1)(1-2s)}}{1-p^{2(1-2s)}}(1-p^{-1})-p^{(h+1)(1-2s)-1},&\text{if }2\nmid h;\vskip 10.03749pt\\ \displaystyle\frac{p^{2(1-2s)}-p^{(h+2)(1-2s)}}{1-p^{2(1-2s)}}(1-p^{-1})&\vskip 10.03749pt\\ \displaystyle\qquad\qquad+\left(\frac{(-1)^{k-1/2}\,n/p^{h}}{p}\right)p^{(h+1)(1-2s)-\frac{1}{2}},&\text{if }2|h.\end{array}\right$. (7.33)
When $n=m^{2}\geq 1$ is a square, by writing $h=\nu_{p}(n)$, we have $2|h$, $(\frac{-n/p^{h}}{p})=1$,
$1\leq\left|\frac{A_{3}(p,-m^{2},s)}{1+p^{-(2s-\frac{1}{2})}}\right|\leq\frac{\nu_{p}(n)}{2}+2\quad\text{ for }\operatorname{Re}s\in[\tfrac{1}{2},1],$ (7.34)
and
$\left.\frac{A_{3}(p,-m^{2},s)}{1+p^{-(2s-\frac{1}{2})}}\right|_{s=\frac{3}{4}}=\frac{1+p^{-1}-p^{-\frac{h+2}{2}}+p^{-\frac{h+1}{2}-\frac{1}{2}}}{1+p^{-1}}=1.$ (7.35)
When $n\neq 0$ is not a negative square, we still have the rough bound
$0\leq|A_{3}(p,n,s)|\leq\frac{\nu_{p}(n)}{2}+2\quad\text{ for }\operatorname{Re}s\in[\tfrac{1}{2},1].$ (7.36)
A special case is that if $p>2$ and $p\nmid n$, i.e. $\nu_{p}(n)=0$, then
$A_{3}(p,n,s)=1+\left(\frac{-n}{p}\right)p^{-(2s-\frac{1}{2})}=1+\left(\frac{-4n}{p}\right)p^{-(2s-\frac{1}{2})}=\frac{1-p^{-(4s-1)}}{1-(\frac{-4n}{p})p^{-(2s-\frac{1}{2})}}.$ (7.37)
We have the following lemma on $A_{3}(p,n,s)$ at $s=\frac{3}{4}$ for $n\neq 0$.
Lemma 7.6.
If $n$ is not a negative square nor 0, then
$\prod_{p\nmid 2n}A_{3}(p,n,s)\quad\text{ is convergent at }s=\tfrac{3}{4}.$
Proof.
This proof can be traced back to [Bat51, Lemma 4.1 and (4.05)]. The product
$\prod_{p\nmid 2n}A_{3}(p,n,s)=\sum_{\begin{subarray}{c}k\text{ square-free}\\ (k,2n)=1\end{subarray}}\left(\frac{-n}{k}\right)\frac{1}{k^{2s-\frac{1}{2}}}=\sum_{k\text{ square-free}}\left(\frac{-4n}{k}\right)\frac{1}{k^{2s-\frac{1}{2}}}$
is convergent at $s=\frac{3}{4}$ and converge to a positive value
$K(-4n)\sum_{(\ell,4n)=1}\frac{\mu(\ell)}{\ell^{2}},\quad\text{where}\quad K(-4n)=\sum_{m=1}^{\infty}\left(\frac{-4n}{m}\right)\frac{1}{m}.\qed$
We conclude the following proposition based on the above discussion.
Proposition 7.7.
For $n\neq 0$, let $Z_{n}(s)$ be the Kloosterman-Selberg zeta function
$Z_{n}(s)=\sum_{2|c>0}\frac{S(n,0,c,\nu_{\Theta}^{3})}{c^{2s}}\quad\text{or}\quad\sum_{2\nmid\widetilde{c}>0}\frac{S(n,0,\widetilde{c},\nu_{\Theta}^{3})}{{\widetilde{c}}^{2s}}.$
Then for either case, there exist functions $C_{1}(n,s)$ and $C_{2}(n,s)$ entire in $s$, and positive absolute constants $C$, $C^{\prime}$ such that for $s\in[\frac{1}{2},1]$,
$0<C^{\prime}\leq|C_{1}(n,s)|\leq C\log n\quad\text{and}\quad|C_{2}(n,s)|\leq C\log n,$
and we have
$Z_{n}(s)=\left\{\begin{array}[]{ll}\displaystyle C_{1}(n,s)\frac{\zeta(2s-\frac{1}{2})}{\zeta(4s-1)},&\text{if }n=-m^{2}<0;\vskip 10.03749pt\\ \displaystyle C_{2}(n,s)\frac{L(2s-\frac{1}{2},(\frac{-4n}{\cdot}))}{\zeta(4s-1)},&\text{other }n\neq 0.\end{array}\right$.
Specifically,
$\operatorname*{Res}_{s=\frac{3}{4}}\sum_{2|c>0}\frac{S(n,0,c,\nu_{\Theta}^{3})}{c^{2s}}=\left\{\begin{array}[]{ll}{\mathbf{e}}(-\frac{1}{8})\frac{2}{\pi^{2}},&\text{ if }n=-m^{2}<0;\\ 0,&\text{ other }n\neq 0\end{array}\right$.
and
$\operatorname*{Res}_{s=\frac{3}{4}}\sum_{2\nmid\widetilde{c}>0}\frac{S(n,0,\widetilde{c},\nu_{\Theta}^{3})}{{\widetilde{c}}^{2s}}=\left\{\begin{array}[]{ll}{\mathbf{e}}(\frac{3}{8})\frac{2}{\pi^{2}},&\text{ if }n=-m^{2}<0;\\ 0,&\text{ other }n\neq 0\end{array}\right$.
Proof.
The proof is direct by combining (7.16), (7.17), (7.20), (7.21), (7.22), (7.34), (7.35), (7.36), (7.37) and Lemma 7.6. Since for every $n$, there are only finitely many $p|2n$, the factors $\zeta(2s-\frac{1}{2})/\zeta(4s-1)$ and $L(2s-\frac{1}{2},(\frac{-4n}{\cdot}))/\zeta(4s-1)$ are the result of (7.37). The residue in the end is given by (7.21), (7.35) and
$\operatorname*{Res}_{s=\frac{3}{4}}\ 2^{2s}\,\frac{{\mathbf{e}}(-\frac{1}{8})}{3\sqrt{2}}\frac{\zeta(2s-\frac{1}{2})}{\zeta(4s-1)}={\mathbf{e}}(-\tfrac{1}{8})\frac{2}{\pi^{2}},\quad\operatorname*{Res}_{s=\frac{3}{4}}\frac{{\mathbf{e}}(\frac{3}{8})\zeta(2s-\frac{1}{2})}{(1+2^{-1})\zeta(4s-1)}={\mathbf{e}}(\tfrac{3}{8})\frac{2}{\pi^{2}}.\qed$
Using Perron’s formula (see e.g. [Dav80, §17]), we get the following proposition.
Proposition 7.8.
For $n\neq 0$, let
$S_{n}(x)=\sum_{2|c\leq x}\frac{S(n,0,c,\nu_{\Theta}^{3})}{c\cdot{\mathbf{e}}(-\frac{1}{8})}\quad\text{or}\quad\sum_{2\nmid\widetilde{c}\leq x}\frac{S(n,0,\widetilde{c},\nu_{\Theta}^{3})}{\widetilde{c}\cdot{\mathbf{e}}(\frac{3}{8})}.$
Then we have
[…]
The same bound holds if $S(n,0,\cdot,\nu_{\Theta}^{3})$ is changed to $S(0,n,\cdot,\nu_{\Theta}^{3})$ because of (2.17).
Proof outline.
We only prove the case for $2|c$ because the other case follows in the same way. We define the Kloosterman-Selberg zeta function as
$Z(s)\vcentcolon=\sum_{2|c>0}\frac{S(n,0,c,\nu_{\Theta}^{3})}{c^{2s}}.$
The Weil bound (7.2) gives us
$\sum_{2|c\leq x}\frac{|S(n,0,c,\nu_{\Theta}^{3})|}{c^{\frac{3}{2}+\delta}}\ll_{\varepsilon,\delta}|n|^{\varepsilon}\log x,\quad\text{for any }\delta>0.$
Applying Perron’s formula, for $\delta>0$ we get
$\left|\sum_{2|c\leq x}\frac{S(n,0,c,\nu_{\Theta}^{3})}{c}-\frac{1}{2\pi i}\int_{\frac{1}{2}+\delta-iT}^{\frac{1}{2}+\delta+iT}Z(\tfrac{1+s}{2})\frac{x^{s}}{s}ds\right|\ll_{\varepsilon,\delta}\frac{x^{\frac{1}{2}+\delta}}{T}|n|^{\varepsilon}\log x.$
We then shift the path of integral to $\frac{1}{2}+\delta-iT\rightarrow\delta-iT\rightarrow\delta+iT\rightarrow\frac{1}{2}+\delta+iT$.
Estimates on the integral along this path follow from the uniform bounds of $\zeta(s)$ and $L(s,(\frac{-4n}{\cdot}))$ in the critical strip. By [MV06, Corollary 1.17], we get the convexity bound
$\zeta(\sigma+it)\ll(1+\tau^{1-\sigma})\min\left(\frac{1}{|\sigma-1|},\log t\right)\quad\text{for }\delta\leq\sigma\leq 2,\ |t|\geq 1.$
By Burgess’s bound [HB78, Theorem 1] [Bur63] we have the subconvexity bound
$L\left(\sigma+it,\left(\frac{-4n}{\cdot}\right)\right)\ll|n|^{\frac{3}{8}(1-\sigma)+\varepsilon}|s|^{\frac{1-\sigma}{4}+\varepsilon}\quad\text{for }\frac{1}{2}\leq\sigma\leq 1,\ |t|\geq 1.$
Therefore, by combining above and Proposition 7.7 we get
$\frac{1}{2\pi i}$ $\left(\int_{\frac{1}{2}+\delta-iT}^{\delta-iT}+\int_{\delta-iT}^{\delta+iT}+\int_{\delta+iT}^{\frac{1}{2}+\delta+iT}\right)Z(\tfrac{1+s}{2})\frac{x^{s}}{s}dx$
$\ll_{\varepsilon,\delta}\left\{\begin{array}[]{ll}|xnT|^{\varepsilon}\left(x^{\delta}T^{\frac{1}{2}}+x^{\frac{1}{2}+\delta}T^{-1}\right),&\ n=-m^{2}<0;\\ |n|^{\frac{3}{16}+\varepsilon}|xnT|^{\varepsilon}\left(x^{\delta}T^{\frac{1}{2}}+x^{\frac{1}{2}+\delta}T^{-1}\right),&\ \text{other }n\neq 0.\end{array}\right$.
In the first case of the proposition where $n$ is a negative square, $Z(\frac{1+s}{2})$ has a pole at $s=\frac{1}{2}$ with residue
$\operatorname*{Res}_{s=\frac{1}{2}}Z(\tfrac{1+s}{2})=2\operatorname*{Res}_{s=\frac{3}{4}}Z(s)={\mathbf{e}}(-\tfrac{1}{8})\frac{4}{\pi^{2}}.$
So the residue of $Z(\tfrac{1+s}{2})\frac{x^{s}}{s}$ at $s=\frac{1}{2}$ is ${\mathbf{e}}(-\tfrac{1}{8})\frac{8}{\pi^{2}}x^{\frac{1}{2}}$. We finish the proof by taking $T=x^{\frac{1}{3}}$, $\delta=\varepsilon$ and combining all of the above calculations. ∎
Remark.
Burgess’s bound has been improved by recent research in subconvexity bounds for $L$-functions, but it is enough for our estimates here due to the worse bound on Kloosterman sums with $mn\neq 0$ in Proposition 7.1 in the $mn$-aspect. Comparing with the proposition, the second author wishes to conclude better “subconvexity bound” for sum of Kloosterman sums in general ($mn\neq 0$). Such estimates will also improve the conclusions in estimates in $n$-aspect in this paper.
There is a remaining case $m=n=0$, which we quickly deal with here. For $2|c>0$,
$S(0,0,c,\nu_{\Theta}^{3})=\left\{\begin{array}[]{ll}\sqrt{2}{\mathbf{e}}(-\frac{1}{8})n\varphi(2n),&\ \frac{\sqrt{2c}}{2}=n\in{\mathbb{Z}}_{+};\\ 0,&\ \text{otherwise}.\end{array}\right$. (7.39)
Therefore, we get
$\sum_{2|c>0}\frac{S(0,0,c,\nu_{\Theta}^{3})}{c^{2s}}$ $=\frac{\sqrt{2}{\mathbf{e}}(-\frac{1}{8})}{2^{1-2s}}\sum_{n=1}^{\infty}\frac{\varphi(2n)}{(2n)^{4s-1}}$
$=\frac{\sqrt{2}{\mathbf{e}}(-\frac{1}{8})}{2^{1-2s}}\cdot\frac{1}{2^{4s-1}-1}\cdot\frac{\zeta(4s-2)}{\zeta(4s-1)},$
$\operatorname*{Res}_{s=\frac{3}{4}}\sum_{2|c>0}\frac{S(0,0,2c,\nu_{\Theta}^{3})}{c^{2s}}=\frac{2{\mathbf{e}}(-\frac{1}{8})}{3}\times\frac{1}{4}\times\frac{6}{\pi^{2}}=\frac{{\mathbf{e}}(-\frac{1}{8})}{\pi^{2}},$ (7.40)
and by Perron’s formula,
$\sum_{2|c\leq x}\frac{S(0,0,c,\nu_{\Theta}^{3})}{c}={\mathbf{e}}(-\tfrac{1}{8})\frac{4}{\pi^{2}}x^{\frac{1}{2}}+O_{\varepsilon}(x^{\frac{1}{6}+\varepsilon}).$ (7.41)
For $2\nmid\widetilde{c}>0$ we also have
[…]
Then
$\sum_{2\nmid\widetilde{c}>0}\frac{S(0,0,\widetilde{c},\nu_{\Theta}^{3})}{{\widetilde{c}}^{2s}}$ $={\mathbf{e}}(\tfrac{3}{8})\sum_{\begin{subarray}{c}n=1\\ n\text{ odd}\end{subarray}}^{\infty}\frac{\varphi(n)}{n^{4s-1}}$
$={\mathbf{e}}(\tfrac{3}{8})\frac{2^{4s-1}-2}{2^{4s-1}-1}\cdot\frac{\zeta(4s-2)}{\zeta(4s-1)}.$
We have
$\operatorname*{Res}_{s=\frac{3}{4}}\sum_{2\nmid\widetilde{c}>0}\frac{S(0,0,\widetilde{c},\nu_{\Theta}^{3})}{{\widetilde{c}}^{2s}}=\frac{2{\mathbf{e}}(\frac{3}{8})}{3}\cdot\frac{1}{4}\cdot\frac{6}{\pi^{2}}=\frac{{\mathbf{e}}(\frac{3}{8})}{\pi^{2}}$ (7.43)
and apply Perron’s formula to get
$\sum_{2\nmid\widetilde{c}\leq x}\frac{S(0,0,\widetilde{c},\nu_{\Theta}^{3})}{\widetilde{c}}={\mathbf{e}}(\tfrac{3}{8})\frac{4}{\pi^{2}}x^{\frac{1}{2}}+O_{\varepsilon}(x^{\frac{1}{6}+\varepsilon}).$ (7.44)
8. Real-variable Kloosterman sums in weight $3/2$
Recall the real-variable Kloosterman sums defined in (3.1). In this section we assume $r^{2}\in{\mathbb{R}}_{+}$. We would like to analyze the convergence of
$\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\quad\text{for }n>0$
and find properties of
$\sum_{2|c>0}\frac{K(r^{2},0,c,\nu_{\Theta}^{3})}{c^{2s}}\quad\text{and}\quad\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\quad\text{for }n<0$
when $s\rightarrow\frac{3}{4}^{+}$, as well as the corresponding cases for $2\nmid\widetilde{c}>0$.
First we prove a Weil-type bound for $K(r^{2},n,c,\nu_{\Theta}^{3})$ and $\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})$ for $2|c$ and $2\nmid{\widetilde{c}}$.
Lemma 8.1.
For $r^{2}\in{\mathbb{R}}_{+}$, $n\neq 0$, $2|c$ and $2\nmid{\widetilde{c}}$, we have
$K(r^{2},n,c,\nu_{\Theta}^{3})\ll_{\varepsilon}(n,c)^{\frac{1}{2}+\varepsilon}c^{\frac{1}{2}+\varepsilon}\quad\text{and}\quad\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})\ll_{\varepsilon}(n,{\widetilde{c}})^{\frac{1}{2}+\varepsilon}{\widetilde{c}}^{\frac{1}{2}+\varepsilon}.$
Proof.
The proof is the same as Lemma 5.1, taking the Weil bound for half-integral weight Kloosterman sums (7.2) into account. For the $2\nmid{\widetilde{c}}$ case, we combine (2.19) and (7.2) to get the desired estimate. ∎
Proposition 8.2.
For $r^{2}\in{\mathbb{R}}_{+}$, $x\geq 1$, $n\in{\mathbb{Z}}\setminus\{0\}$, $2|c$ and $2\nmid{\widetilde{c}}$, we have
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c\leq x}\frac{|K(r^{2},n,c,\nu_{\Theta}^{3})|}{c}\\ \displaystyle\sum_{2\nmid{\widetilde{c}}\leq x}\frac{|\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})|}{\widetilde{c}}\\ \end{array}\right\}\ll_{\varepsilon}|n|^{\varepsilon}x^{\frac{1}{2}+\varepsilon}.$
Proof.
The proof follows directly from Lemma 8.1 and the Weil bound of Kloosterman sums (when $r^{2}\in{\mathbb{Z}}$). ∎
8.1. Case $n>0$
We begin with an estimate for sums of real-variable Kloosterman sums.
Proposition 8.3.
For $r^{2}\in{\mathbb{R}}_{+}$, $x\geq 1$, $n\in{\mathbb{Z}}_{+}$, and $X=\max(x,r^{2})$, we have
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c\leq x}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}\\ \displaystyle\sum_{2\nmid{\widetilde{c}}\leq x}\frac{\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})}{\widetilde{c}}\end{array}\right\}\ll_{\varepsilon}X^{\frac{1}{4}}n^{\frac{3}{16}}\left(n^{\frac{1}{16}-\delta}+u_{n}^{\frac{1}{8}}\right)|Xn|^{\varepsilon},$
where $\delta=\frac{1}{147}$ and $u_{n}=2^{\left\lfloor\nu_{2}(n)\right\rfloor/2}\ll n^{1/2}$.
Proof.
When $r^{2}\in{\mathbb{Z}}_{+}$, the proposition follows from the better bounds in Proposition 7.1 and Proposition 7.2. We prove the bound for $r^{2}\in{\mathbb{R}}_{+}\setminus{\mathbb{Z}}$ and the sum on $2|c$ and the other case $2\nmid{\widetilde{c}}$ follows similarly. By Lemma 2.8, we have
$\begin{split}\sum_{2|c\leq x}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}&=\sum_{2|c\leq x}\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ |r^{2}-k|<c\end{subarray}}\frac{i\sin(\pi r^{2})(-1)^{k}}{{\mathbf{e}}(\frac{r^{2}-k}{2c})-1}\frac{S(k,n,c,\nu_{\Theta}^{3})}{c^{2}}\\ &=\frac{i\sin(\pi r^{2})}{\pi}\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ |r^{2}-k|\leq x\end{subarray}}(-1)^{k}\sum_{\begin{subarray}{c}2|c\\ |r^{2}-k|<c\leq x\end{subarray}}\frac{S(k,n,c,\nu_{\Theta}^{3})}{\left({\mathbf{e}}(\frac{r^{2}-k}{2c})-1\right)c^{2}}\end{split}$ (8.1)
For $k\in{\mathbb{Z}}$ and $n>0$, we apply Proposition 7.1 and Proposition 7.8 to get
$S(x)\vcentcolon=\sum_{2|c\leq x}\frac{S(k,n,c,\nu_{\Theta}^{3})}{c}\ll_{\varepsilon}\left(x^{\frac{1}{6}}+A_{u}(k,n)\right)|xkn|^{\varepsilon}.$ (8.2)
By partial summation and (5.1) we get
$\sum_{\begin{subarray}{c}2|c\\ |r^{2}-k|<c\leq x\end{subarray}}$ $\frac{S(k,n,c,\nu_{\Theta}^{3})}{\left({\mathbf{e}}(\frac{r^{2}-k}{2c})-1\right)c^{2}}=\int_{|r^{2}-k|}^{x}\frac{t^{-1}dS(t)}{{\mathbf{e}}(\frac{r^{2}-k}{2t})-1}$
$=\frac{t^{-1}S(t)}{{\mathbf{e}}(\frac{r^{2}-k}{2t})-1}\Big|_{|r^{2}-k|}^{x}+\int_{|r^{2}-k|}^{x}S(t)\frac{{\mathbf{e}}(\frac{r^{2}-k}{2t})-1+\pi i{\mathbf{e}}(\frac{r^{2}-k}{2t})\frac{r^{2}-k}{t}}{\left({\mathbf{e}}(\frac{r^{2}-k}{2t})-1\right)^{2}t^{2}}dt$
$\ll\left(x^{\frac{1}{6}}+A_{u}(k,n)\right)\frac{|xkn|^{\varepsilon}}{|r^{2}-k|}.$
By Proposition 7.1,
$A_{u}(k,n)\ll|k|^{\frac{1}{4}}n^{\frac{3}{16}}\left(n^{\frac{1}{16}-\delta}+u_{n}^{\frac{1}{8}}\right),$
where $\delta=\frac{1}{147}$ and $u_{n}=2^{\left\lfloor\nu_{2}(n)/2\right\rfloor}\leq n^{1/2}$.
If $x<r^{2}$, summing $k$ in $|r^{2}-k|\leq x$ with the help of Lemma 2.10 and $\frac{\sin(\pi r^{2})}{\|r^{2}\|}\in[2,\pi]$, we get the desired bound. If $x>r^{2}$, we have an extra contribution from $k=0$. By Proposition 7.8, a similar partial summation process as above gives a better bound:
$\sum_{\begin{subarray}{c}2|c\\ r^{2}<c\leq x\end{subarray}}\frac{S(0,n,c,\nu_{\Theta}^{3})}{({\mathbf{e}}(\frac{r^{2}}{2c})-1)c^{2}}\ll_{\varepsilon}x^{\frac{1}{6}+\varepsilon}n^{\frac{3}{16}+\varepsilon}.$
In either case we get the desired bound and the proposition is proved. ∎
Comparing Proposition 8.2 and Proposition 8.3, we find that Proposition 8.2 is better when $x$ is small, while Proposition 8.3 is better when $x$ is large. In particular, Proposition 8.3 allows us to establish the convergence of sums in the following result, since the exponent of $x$ is smaller than $\frac{1}{2}$. To estimate the growth rate of the following sums as $n$ becomes large, both of the propositions above should be applied to refine the estimate.
Proposition 8.4.
For $r^{2}\in{\mathbb{R}}_{+}$ and $n\in{\mathbb{Z}}_{+}$, the following sums are convergent for $s\geq\frac{3}{4}$, and both have the same estimate depending on $r^{2}n\leq 1$ or $r^{2}n\geq 1$:
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ \displaystyle\sum_{2\nmid{\widetilde{c}}>0}\frac{\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})}{\widetilde{c}}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{{\widetilde{c}}}\right)\end{array}\right\}\ll_{\varepsilon}\left\{\begin{array}[]{ll}|r|^{\frac{1}{2}}\max\left(n^{\frac{145}{294}},n^{\frac{7+\kappa}{16}}\right)n^{\varepsilon}&\text{if }r^{2}\leq\frac{1}{n},\\ (1+|r|^{\frac{1}{2}+\varepsilon})n^{\frac{1}{4}+\varepsilon}&\text{if }r^{2}\geq\frac{1}{n}.\end{array}\right$.
Here $\kappa\in[0,1]$ is determined by $2^{\nu_{2}(n)}\ll|n|^{\kappa}$.
Remark.
For simplicity, the exponent of $n$ in the last bound is $\frac{145}{294}\approx 0.493\ldots$ if $\nu_{2}(n)$ is absolutely bounded, e.g. if $n$ is odd, and the exponent is $\frac{1}{2}=0.5$ if $n$ is a power of $2$.
Proof.
We prove the case for $2|c>0$ and the other case $2\nmid{\widetilde{c}}>0$ can be proved in the same way. The convergence for $\operatorname{Re}(s)>1$ is given by the definition in (3.8), so we only need to focus on $s\in[\frac{3}{4},1+\delta]$ for any fixed $\delta>0$.
By [DLMF, (10.7.3), (10.7.8), (10.14.1), (10.14.4)], for $x\in{\mathbb{R}}$ and $\nu\geq-\frac{1}{2}$,
$J_{\nu}(x)\ll\min\left(x^{-\frac{1}{2}},x^{\nu}\right),\quad\text{hence }J_{2s-1}(x)\ll\min\left(x^{-\frac{1}{2}},x^{\frac{1}{2}}\right).$ (8.3)
By [DLMF, (10.6.2)],
$\begin{split}\frac{d}{dt}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{t}\right)&=\frac{2\pi|r|\sqrt{n}}{t^{2}}J_{2s}\left(\frac{2\pi|r|\sqrt{n}}{t}\right)-\frac{(2s-1)}{t}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{t}\right)\\ &\ll\frac{|r^{2}n|^{s-\frac{1}{2}}}{t^{2s}}\ll\frac{|r^{2}n|^{\frac{1}{4}}}{t^{\frac{3}{2}}}\quad\text{ if }t\gg|r|n^{\frac{1}{2}}.\end{split}$ (8.4)
The convergence in the proposition is by partial summation and Cauchy’s criterion: for $y>x>\max(1,r^{2},2\pi|r|\sqrt{n})$, Proposition 8.3 and (8.4) gives
$\sum_{\begin{subarray}{c}2|c\\ x<c\leq y\end{subarray}}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\ll_{\varepsilon}|r|^{\frac{1}{2}}n^{\frac{1}{2}+\varepsilon}x^{-\frac{1}{4}+\varepsilon}.$ (8.5)
Let $\beta>\frac{1}{2}$, we will choose it later. We have the following two cases:
(1) when $r^{2}\geq\frac{1}{n}$, by Lemma 8.1 and (8.3), we have
$\begin{split}\sum_{2|c\leq|r^{2}n|^{\beta}}&\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ &\ll_{\varepsilon}|r^{2}n|^{-\frac{1}{4}}\sum_{2|c\leq|r|\sqrt{n}}(n,c)^{\frac{1}{2}+\varepsilon}c^{\varepsilon}+|r^{2}n|^{\frac{1}{4}}\sum_{\begin{subarray}{c}2|c\\ |r^{2}n|^{1/2}\leq c\leq|r^{2}n|^{\beta}\end{subarray}}(n,c)^{\frac{1}{2}+\varepsilon}c^{-1+\varepsilon}\\ &\ll_{\varepsilon}|r^{2}n|^{\frac{1}{4}+\beta\varepsilon+\varepsilon}.\end{split}$ (8.6)
By partial sum with Proposition 8.2 and (8.4), we have
$\sum_{2|c\geq|r^{2}n|^{\beta}}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}J_{\frac{1}{2}}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\ll_{\varepsilon}\max(|r|^{\beta},|r|)^{\frac{1}{2}}n^{\frac{7}{16}-\frac{\beta}{4}+\varepsilon}\left(n^{\frac{1}{16}-\delta}+u_{n}^{\frac{1}{8}}\right).$ (8.7)
Whenever $\beta\geq 1$, the exponent of $n$ in (8.7) is dominated by (8.6).
(2) when $r^{2}\leq\frac{1}{n}$, by partial sum with Proposition 8.2 and (8.4), we also have
$\sum_{2|c}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}J_{\frac{1}{2}}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\ll_{\varepsilon}|r|^{\frac{1}{2}}n^{\frac{7}{16}+\varepsilon}\left(n^{\frac{1}{16}-\delta}+u_{n}^{\frac{1}{8}}\right).$ (8.8)
The proof for the case $2|c$ is done by combining (8.8) and (8.6) with $\beta=1$. ∎
8.2. Case $n=0$
Proposition 8.5.
For $r\in{\mathbb{R}}_{+}$, we have
$\lim_{s\rightarrow\frac{3}{4}^{+}}\frac{1}{\Gamma(s-\frac{3}{4})}\sum_{2|c>0}\frac{K(r^{2},0,c,\nu_{\Theta}^{3})}{c^{2s}}={\mathbf{e}}(-\tfrac{1}{8})\frac{\sin(\pi r^{2})}{\pi^{2}r\sinh(\pi r)},$
$\lim_{s\rightarrow\frac{3}{4}^{+}}\frac{1}{\Gamma(s-\frac{3}{4})}\sum_{2\nmid{\widetilde{c}}>0}\frac{\widetilde{K}(r^{2},0,{\widetilde{c}},\nu_{\Theta}^{3})}{{\widetilde{c}}^{2s}}={\mathbf{e}}(\tfrac{3}{8})\frac{\sin(\pi r^{2})}{\pi^{2}r\sinh(\pi r)}.$
Proof.
The case $r^{2}\in{\mathbb{Z}}_{+}$ follows from Proposition 7.7 and that for any $m\in{\mathbb{Z}}_{+}$,
$\lim_{r\rightarrow\sqrt{m}}\frac{\sin(\pi r^{2})}{r\sinh(\pi r)}=0.$ (8.9)
So we only need to prove the case $r^{2}\in{\mathbb{R}}_{+}\setminus{\mathbb{Z}}$.
By Lemma 2.8 and Lemma 2.9, we have
$\sum_{2|c>0}$ $\frac{K(r^{2},0,c,\nu_{\Theta}^{3})}{c^{2s}}=\sum_{2|c>0}\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ |r^{2}-k|<c\end{subarray}}\frac{i\sin(\pi r^{2})(-1)^{k}}{{\mathbf{e}}(\frac{r^{2}-k}{2c})-1}\cdot\frac{S(k,0,c,\nu_{\Theta}^{3})}{c^{2s+1}}$
$=\frac{\sin(\pi r^{2})}{\pi}\sum_{k\in{\mathbb{Z}}}\frac{(-1)^{k}}{r^{2}-k}\sum_{2|c>|r^{2}-k|}\frac{S(k,0,c,\nu_{\Theta}^{3})}{c^{2s}}$
$+\frac{\sin(\pi r^{2})}{\pi}\sum_{k\in{\mathbb{Z}}}\sum_{\ell=1}^{\infty}B_{\ell}\frac{(-1)^{k}(\pi i)^{\ell}(r^{2}-k)^{\ell}}{(r^{2}-k)\ell!}\sum_{2|c>|r^{2}-k|}\frac{S(k,0,c,\nu_{\Theta}^{3})}{c^{2s+\ell}}$
$=:\Sigma_{0}+\Sigma_{1}.$
We are going to prove that the summands for $k=-m^{2}\leq 0$ in $\Sigma_{0}$ contributes to the result in the proposition, and prove that $\Sigma_{1}$ and the remaining sums in $\Sigma_{0}$ are convergent when $s\rightarrow\frac{3}{4}^{+}$.
We begin by estimating the innermost sum for $\ell\geq 0$. Let
$A(x)\vcentcolon=\sum_{2|c\leq x}\frac{S(k,0,c,\nu_{\Theta}^{3})}{c}\quad\text{and}\quad\mathfrak{R}\vcentcolon={\mathbf{e}}(-\tfrac{1}{8})\frac{8}{\pi^{2}}.$
By Proposition 7.8, when $k=-m^{2}<0$ and $s>\frac{3}{4}$, we have
$\sum_{2|c>r^{2}+m^{2}}$ $\frac{S(-m^{2},0,c,\nu_{\Theta}^{3})}{c^{2s+\ell}}=\int_{r^{2}+m^{2}}^{\infty}\frac{dA(t)}{t^{2s+\ell-1}}$
$=-\frac{\mathfrak{R}(r^{2}+m^{2})^{\frac{1}{2}}+O_{\varepsilon}((r^{2}+m^{2})^{\frac{1}{6}+\varepsilon}(m^{2})^{\frac{1}{4}+\varepsilon})}{(r^{2}+m^{2})^{2s+\ell-1}}$
$-(1-2s-\ell)\int_{r^{2}+m^{2}}^{\infty}\frac{\mathfrak{R}t^{\frac{1}{2}}+O(t^{\frac{1}{6}+\varepsilon}m^{\frac{1}{4}+\varepsilon})}{t^{2s+\ell}}dt$
$=\frac{1}{4s-3+2\ell}\cdot\frac{\mathfrak{R}}{(r^{2}+m^{2})^{2s+\ell-\frac{3}{2}}}+O_{\varepsilon}\left(\frac{1}{(r^{2}+m^{2})^{2s+\ell-\frac{17}{12}-\varepsilon}}\right),$
It is important to note that the last implied constant does not depend on $\ell\geq 0$, because it is uniformly bounded from above and below for $s\in[\frac{3}{4},1]$ and $\ell\geq 0$.
Similarly, we also have
$\sum_{2|c>r^{2}}\frac{S(0,0,c,\nu_{\Theta}^{3})}{c^{2s+\ell}}$ $=\frac{1}{4s-3+2\ell}\cdot\frac{\mathfrak{R}/2}{(r^{2})^{2s+\ell-\frac{3}{2}}}+O_{\varepsilon}\left((r^{2})^{-2s-\ell+\frac{7}{6}+\varepsilon}+1\right).$
Additionally, when $k\neq 0$ nor a negative square,
$\sum_{2|c>|r^{2}-k|}\frac{S(k,0,c,\nu_{\Theta}^{3})}{c^{2s+\ell}}$ $\ll_{\varepsilon}|r^{2}-k|^{-2s-\ell+\frac{17}{12}+\varepsilon}.$
Therefore, we can apply the $\ell=0$ case in $\Sigma_{0}$ to get
$\Sigma_{0}$ $=\frac{\sin(\pi r^{2})\mathfrak{R}}{2\pi(4s-3)}\left(\frac{1}{(r^{2})^{2s-\frac{1}{2}}}+\sum_{m=1}^{\infty}\frac{2(-1)^{m}}{(r^{2}+m^{2})^{2s-\frac{1}{2}}}\right)$
$+\frac{\sin(\pi r^{2})}{\pi}O_{\varepsilon}\left(\sum_{k\in{\mathbb{Z}}}|r^{2}-k|^{-2s+\frac{5}{12}+\varepsilon}\right)$
$=\frac{\sin(\pi r^{2})\mathfrak{R}}{2\pi(4s-3)}\left(\frac{1}{(r^{2})^{2s-\frac{1}{2}}}+\sum_{m=1}^{\infty}\frac{2(-1)^{m}}{(r^{2}+m^{2})^{2s-\frac{1}{2}}}\right)+O(1).$
The last $O(1)$ part is bounded by an absolute constant because $2s-\frac{5}{12}\geq\frac{13}{12}$ and $\frac{\sin(\pi r^{2})}{\|r^{2}\|}\in[2,\pi]$ for all $r\in{\mathbb{R}}$. Since $r>0$ and
$\frac{1}{r^{2}}+\sum_{m=1}^{\infty}\frac{2(-1)^{m}}{r^{2}+m^{2}}=\frac{\pi}{r\sinh(\pi r)},$ (8.10)
we get
$\lim_{s\rightarrow\frac{3}{4}^{+}}\frac{\Sigma_{0}}{\Gamma(s-\frac{3}{4})}=\frac{\sin(\pi r^{2})\mathfrak{R}}{8\pi}\cdot\frac{\pi}{r\sinh(\pi r)}={\mathbf{e}}(-\tfrac{1}{8})\frac{\sin(\pi r^{2})}{\pi^{2}r\sinh(\pi r)}.$
For $\Sigma_{1}$, we have
$\Sigma_{1}$ $=\frac{\sin(\pi r^{2})\mathfrak{R}}{2\pi}\left(\frac{1}{(r^{2})^{2s-\frac{1}{2}}}+\sum_{m=1}^{\infty}\frac{2(-1)^{m}}{(r^{2}+m^{2})^{2s-\frac{1}{2}}}\right)\sum_{\ell=1}^{\infty}\frac{B_{\ell}(\pi i)^{\ell}}{\ell!(4s-3+2\ell)}$
$+\frac{\sin(\pi r^{2})}{\pi}O_{\varepsilon}\left(\sum_{k\in{\mathbb{Z}}}|r^{2}-k|^{-2s+\frac{5}{12}+\varepsilon}\sum_{\ell=1}^{\infty}\frac{B_{\ell}(\pi i)^{\ell}}{\ell!(4s-3+2\ell)}\right).$
It is then direct to conclude $\Sigma_{1}$ converges to a finite value for $s\rightarrow\frac{3}{4}^{+}$ because
$\sum_{\ell=1}^{\infty}\frac{B_{\ell}(\pi i)^{\ell}}{\ell!(4s-3+2\ell)}\quad\text{converges for }s\in[\tfrac{3}{4},1].$
This is the reason why we separate the $\ell=0$ and $\ell\geq 1$ terms in Lemma 2.9. We have finished the proof for the case $2|c$.
The other case $2\nmid{\widetilde{c}}$ follows by the same process applying Proposition 7.8 and (7.44). ∎
8.3. Case $n<0$
Similarly to the proof of Proposition 5.3, the following proposition follows from Lemma 8.1, Proposition 8.2, partial summation, Proposition 8.3, and by [DLMF, (10.29.1)]
[…]
. We omit the proof here.
Proposition 8.6.
Let $r^{2}\in{\mathbb{R}}_{+}$, $n\in{\mathbb{Z}}$, $n<0$ and $s\in[\frac{3}{4},1.001]$.
(1) If $n$ is not a negative square, the following sums are convergent and we have
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\\ \displaystyle\sum_{2\nmid{\widetilde{c}}>0}\frac{\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})}{{\widetilde{c}}}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{{\widetilde{c}}}\right)\end{array}\right\}\ll(1+|r|^{2})|n|e^{2\pi|r||n|^{1/2}}.$
(2) If $r^{2}=q\in{\mathbb{Z}}_{+}$, $m\in{\mathbb{Z}}_{+}$ and $n=-m^{2}$, the following sums are convergent and we have
$\begin{array}[]{r}\displaystyle\sum_{2|c>0}\frac{S(q,-m^{2},c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi m\sqrt{q}}{c}\right)\\ \displaystyle\sum_{2\nmid{\widetilde{c}}>0}\frac{S(q,-m^{2},{\widetilde{c}},\nu_{\Theta}^{3})}{{\widetilde{c}}}I_{2s-1}\left(\frac{2\pi m\sqrt{q}}{{\widetilde{c}}}\right)\end{array}\ll|qn|e^{2\pi|qn|^{1/2}}.$
Remark.
In case (1), we need the estimates for
$\sum_{2|c\leq x}\frac{S(k,n,c,\nu_{\Theta}^{3})}{c}\quad\text{and}\quad\sum_{2\nmid{\widetilde{c}}\leq x}\frac{S(k,n,{\widetilde{c}},\nu_{\Theta}^{3})}{{\widetilde{c}}}$
for all $k\in{\mathbb{Z}}$; in case (2), we keep $q>0$. In both cases, we avoid to have both $k,n\in\{-t^{2}:t\in{\mathbb{Z}}\}$, which corresponds to the “other $mn\neq 0$” cases in Proposition 7.1 and Proposition 7.3.
Proposition 8.7.
For $r\in{\mathbb{R}}_{+}$ and $m,n\in{\mathbb{Z}}$, $n<0$, let
$\mathscr{F}(r,n)=\left\{\begin{array}[]{ll}\displaystyle{\mathbf{e}}(-\tfrac{1}{8})\frac{4(rm)^{\frac{1}{2}}\sin(\pi r^{2})}{\pi^{2}r\sinh(\pi r)},&\text{if }n=-m^{2}<0,\\ 0,&\text{if }n\text{ is not a negative square,}\end{array}\right$.
and $\widetilde{\mathscr{F}}(r,n)=-\mathscr{F}(r,n)$. Then we have the following limits:
$\lim_{s\rightarrow\frac{3}{4}^{+}}\frac{1}{\Gamma(s-\frac{3}{4})}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi r|n|^{\frac{1}{2}}}{c}\right)$ $=\mathscr{F}(r,n),$
[…]
Remark.
Heuristically, this proposition is similar to Proposition 8.5, by noticing that the coefficient of $x^{\frac{1}{2}}$ in Proposition 7.1 is twice the value in Proposition 7.8 and four times the value in (7.41), and by [DLMF, (10.30.1)]:
$I_{2s-1}\left(\frac{2\pi r|n|^{\frac{1}{2}}}{c}\right)\sim\frac{1}{\Gamma(2s)}\left(\frac{\pi r|n|^{\frac{1}{2}}}{c}\right)^{2s-1}\quad\text{for }c\rightarrow\infty.$
We give a detailed argument below.
Proof.
Recall (8.9), hence the case for $n<0$ not a negative square or for $r^{2}\in{\mathbb{Z}}_{+}$ has been covered by Proposition 8.6. Here we only need to consider the case $n=-m^{2}<0$ and $r^{2}\in{\mathbb{R}}_{+}\setminus{\mathbb{Z}}$. We write $m>0$ for simplicity.
By applying Lemma 2.8 and Lemma 2.9, we get
$\sum_{2|c>0}\frac{K(r^{2},-m^{2},c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi rm}{c}\right)$
$=\frac{\sin(\pi r^{2})}{\pi}\sum_{k\in{\mathbb{Z}}}\frac{(-1)^{k}}{r^{2}-k}\sum_{2|c>|r^{2}-k|}\frac{S(k,-m^{2},c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi rm}{c}\right)$
$+\frac{\sin(\pi r^{2})}{\pi}\sum_{k\in{\mathbb{Z}}}\sum_{\ell=1}^{\infty}B_{\ell}\frac{(-1)^{k}(\pi i)^{\ell}(r^{2}-k)^{\ell}}{(r^{2}-k)\ell!}\sum_{2|c>|r^{2}-k|}\frac{S(k,-m^{2},c,\nu_{\Theta}^{3})}{c^{1+\ell}}I_{2s-1}\left(\frac{2\pi rm}{c}\right)$
$=:\Sigma_{0}+\Sigma_{1},$
where $\Sigma_{0}$ is given by
$\Sigma_{0}=\frac{\sin(\pi r^{2})}{\pi}\sum_{k\in{\mathbb{Z}}}\frac{(-1)^{k}}{r^{2}-k}\sum_{2|c>|r^{2}-k|}\frac{S(k,-m^{2},c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi rm}{c}\right).$ (8.12)
We estimate the latter sum with the help of results in §7. Let
$A(x)\vcentcolon=\sum_{2|c\leq x}\frac{S(k,-m^{2},c,\nu_{\Theta}^{3})}{c}\quad\text{and}\quad\mathfrak{R}\vcentcolon={\mathbf{e}}(-\tfrac{1}{8})\frac{16}{\pi^{2}}.$
By [DLMF, (10.29.2)], for $\alpha=2\pi rm$,
$\frac{d}{dt}I_{2s-1}\left(\frac{\alpha}{t}\right)=-\frac{\alpha}{t^{2}}I_{2s}\left(\frac{\alpha}{t}\right)-\frac{2s-1}{t}I_{2s-1}\left(\frac{\alpha}{t}\right).$ (8.13)
By [DLMF, (10.25.2)],
$I_{\nu}(x)=\frac{(x/2)^{\nu}}{\Gamma(\nu+1)}+O(x^{\nu+2})\quad\text{for }\nu\in[0,3]\text{ and }x\in[0,1].$ (8.14)
By [DLMF, (10.30.4)] and (8.14),
$I_{\nu}(x)\ll e^{x}\quad\text{for }\nu\in[0,3]\text{ and }x\geq 0.$ (8.15)
When $k=-q^{2}<0$ with $q>0$, note that $r^{2}+q^{2}\geq 2rq$ and $q\geq\pi m$ implies $r^{2}+q^{2}\geq 2\pi rm$. By applying Proposition 7.1, (8.14), and (8.15), we have that for $s\in(\frac{3}{4},1.001]$, $2s-1>\frac{1}{2}$ and
$\sum_{2|c>r^{2}+q^{2}}\frac{S(-q^{2},-m^{2},c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi rm}{c}\right)$
$=A(t)I_{2s-1}\left(\frac{2\pi rm}{t}\right)\Bigg|_{t=r^{2}+q^{2}}^{\infty}-\int_{r^{2}+q^{2}}^{\infty}A(t)\frac{d}{dt}I_{2s-1}\left(\frac{2\pi rm}{t}\right)dt$
$=-\frac{\mathfrak{R}(r^{2}+q^{2})^{\frac{1}{2}}}{\Gamma(2s)}\left(\frac{\pi rm}{r^{2}+q^{2}}\right)^{2s-1}+\frac{\mathfrak{R}(2s-1)}{\Gamma(2s)}\int_{r^{2}+q^{2}}^{\infty}{t^{-\frac{1}{2}}}\left(\frac{\pi rm}{t}\right)^{2s-1}dt$
$\ +\left\{\begin{array}[]{ll}\displaystyle O_{\varepsilon}\left((1+r^{3})m^{3}(r^{2}+q^{2})^{\frac{5}{4}-2s+\varepsilon}\right),&\ q\geq\pi m\\ \displaystyle O_{\varepsilon}\left((1+r^{3})m^{3}e^{2\pi rm}\right),&\ q\leq\pi m\end{array}\right$.
$=\frac{\mathfrak{R}(\pi rm)^{2s-1}}{(4s-3)\Gamma(2s)}(r^{2}+q^{2})^{\frac{3}{2}-2s}+\left\{\begin{array}[]{ll}\displaystyle O_{\varepsilon}\left((1+r^{3})m^{3}(r^{2}+q^{2})^{\frac{5}{4}-2s+\varepsilon}\right),&\ q\geq\pi m\\ \displaystyle O\left((1+r^{3})m^{3}e^{2\pi rm}\right),&\ q\leq\pi m\end{array}\right$.
with crude bounds in the big-$O$ term. Similarly, with Proposition 7.8, we have
$\sum_{2|c>r^{2}}$ $\frac{S(0,-m^{2},c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi rm}{c}\right)$
$=\frac{(\pi rm)^{2s-1}\mathfrak{R}/2}{(4s-3)\Gamma(2s)}(r^{2})^{\frac{3}{2}-2s}+O\left((1+r^{3})m^{3}e^{2\pi rm}\right).$
For $k\neq 0$ nor negative square, we also have the crude bound
$\sum_{2|c>|r^{2}-k|}\frac{S(k,-m^{2},c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi rm}{c}\right)=O_{\varepsilon}\left(\frac{(1+r^{3})m^{3}e^{2\pi rm}}{|r^{2}-k|^{2s-\frac{5}{4}-\varepsilon}}\right).$
Recall (8.12). Combining the three equations above and noting that there are at most $O(m)$ of $q$’s such that $q\leq\pi m$, we conclude for $\Sigma_{0}$ that
$\lim_{s\rightarrow\frac{3}{4}^{+}}\frac{\Sigma_{0}}{\Gamma(s-\frac{3}{4})}$ $=\lim_{s\rightarrow\frac{3}{4}^{+}}\frac{(\pi rm)^{2s-1}\mathfrak{R}/2}{\Gamma(s-\frac{3}{4})(4s-3)}\cdot\frac{\sin(\pi r^{2})}{\pi\Gamma(2s)}\left(\frac{1}{r^{2}}+\sum_{q=1}^{\infty}\frac{2(-1)^{q}}{r^{2}+q^{2}}\right)$
$\ +\lim_{s\rightarrow\frac{3}{4}^{+}}\frac{\sin(\pi r^{2})}{\Gamma(s-\frac{3}{4})}O_{\varepsilon}\left(\sum_{k\in{\mathbb{Z}}}\frac{(1+r^{3})m^{4}e^{2\pi rm}}{|r^{2}-k|^{2s-\frac{9}{4}-\varepsilon}}\right)$
$={\mathbf{e}}(-\tfrac{1}{8})(rm)^{\frac{1}{2}}\frac{4\sin(\pi r^{2})}{\pi^{2}r\sinh(\pi r)}.$
The similar method on $\Sigma_{1}$ in the proof of Proposition 8.5 allows us to conclude that
$\Sigma_{1}=O((1+r^{3})m^{4}e^{2\pi rm})\quad\text{and}\quad\lim_{s\rightarrow\frac{3}{4}^{+}}\frac{\Sigma_{1}}{\Gamma(s-\frac{3}{4})}=0.$
The case for $2\nmid{\widetilde{c}}$ follows similarly. We have finished the proof. ∎
Combining Proposition 8.6 and the proof of Proposition 8.7, we have the following crude estimate.
Proposition 8.8.
For $r^{2}\in{\mathbb{R}}_{+}$ and $m,n\in{\mathbb{Z}}$, $n<0$, let $\mathscr{F}(r,n)$ and $\widetilde{\mathscr{F}}(r,n)$ be the same as in Proposition 8.7. Then
$\left.\begin{array}[]{r}\displaystyle\frac{1}{\Gamma(s-\frac{3}{4})}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)-\mathscr{F}(r,n)\\ \displaystyle\frac{1}{\Gamma(s-\frac{3}{4})}\sum_{2\nmid{\widetilde{c}}>0}\frac{\widetilde{K}(r^{2},n,{\widetilde{c}},\nu_{\Theta}^{3})}{{\widetilde{c}}}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{{\widetilde{c}}}\right)-\widetilde{\mathscr{F}}(r,n)\end{array}\right\}\ll(1+|r|^{3})|n|^{2}e^{2\pi|r||n|^{1/2}}.$
9. Proof of Theorem 3.6 and Theorem 1.3
Based on the discussion of weight $3/2$ real-variable Kloosterman sums from the previous section, we are now ready to prove Theorem 3.6.
Proof of Theorem 3.6.
Recall (3.6) and the bound (6.1) of the Whittaker function. By combining Lemma 3.4, Proposition 8.4, Proposition 8.5, and Proposition 8.8, we find that $B_{3,n}(r;s,y)$ and $\widetilde{B}_{3,n}(r;s,y)$ have the claimed bounds for $n\neq 0$ and the limits of $B_{3,0}(r;s,y)$ and $\widetilde{B}_{3,0}(r;s,y)$ for $s\rightarrow\frac{3}{4}^{+}$ exist.
By dominated convergence theorem, the exponential decay $e^{-\pi|n|y}$ ensures that the Fourier expansions of $F_{\frac{3}{2}}(\tau;r,s)$ (3.13) and of $\widetilde{F}_{\frac{3}{2}}(\tau;r,s)$ (3.14) are absolutely convergent for $s\in[\frac{3}{4},1.001]$. By (3.6), Proposition 8.5, Proposition 8.7 and by analytic continuation, we conclude
$F_{\frac{3}{2}}(\tau;r,\tfrac{3}{4})=\lim_{s\rightarrow\frac{3}{4}^{+}}F_{\frac{3}{2}}(\tau;r,s)\quad\text{and}\quad\widetilde{F}_{\frac{3}{2}}(\tau;r,\tfrac{3}{4})=\lim_{s\rightarrow\frac{3}{4}^{+}}\widetilde{F}_{\frac{3}{2}}(\tau;r,s)$
and Theorem 3.6 follows. ∎
The rest of this section is devoted to the proof of Theorem 1.3. We use Zagier’s famous weight $\frac{3}{2}$ mock modular form on $\Gamma_{0}(4)$ in [Zag75, Théorème 2]. Here $q=e^{2\pi i\tau}$ and $\tau=x+iy\in{\mathbb{H}}$.
Theorem 9.1 (Zagier).
Let $H(n)$ be the Hurwitz class number. Then the function
$\mathcal{H}(\tau)\vcentcolon=-\frac{1}{12}+\sum_{n=1}^{\infty}H(n)q^{n}+\frac{1}{8\pi\sqrt{y}}+\frac{1}{4\sqrt{\pi}}\sum_{m=1}^{\infty}m\Gamma(-\tfrac{1}{2},4\pi m^{2}y)q^{-m^{2}}$ (9.1)
transform like a weight $3/2$ modular form on $\Gamma_{0}(4)$.
More precisely, we have the following transformation formula for $\mathcal{H}$:
$\mathcal{H}(\gamma\tau)=\nu_{\theta}(\gamma)^{3}(c\tau+d)^{\frac{3}{2}}\mathcal{H}(\tau)\quad\text{for }\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}(4).$ (9.2)
By [HZ76, §2.2], we write
$\mathscr{H}_{1}(\tau)=-\frac{1}{12}+\sum_{n=1}^{\infty}H(n)q^{n}\quad\text{and}\quad\mathcal{H}(\tau)-\mathscr{H}_{1}(\tau)=\frac{1+i}{16\pi}\psi(\tau)$ (9.3)
where
$\psi(\tau)=\int_{-\overline{\tau}}^{i\infty}\frac{\theta(v)}{(\tau+v)^{\frac{3}{2}}}dv$ (9.4)
and the integral is taken along the vertical path $v=2iuy-\tau$, $u\in[-1,\infty)$. Then the discussion after [HZ76, Theorem 2, Corollary], together with $\theta(-\frac{1}{4\tau})=\sqrt{-2i\tau}\theta(\tau)$, shows that
$\begin{split}(-2i\tau)^{-\frac{3}{2}}\mathscr{H}_{1}(-\tfrac{1}{4\tau})+\mathscr{H}_{1}(\tau)&=-\frac{1}{24}\theta(\tau)^{3}-\sqrt{\frac{\tau}{8i}}\int_{{\mathbb{R}}}{\mathbf{e}}(\xi^{2}\tau)\frac{1+{\mathbf{e}}(2\xi\tau)}{1-{\mathbf{e}}(2\xi\tau)}\xi d\xi,\\ (-2i\tau)^{-\frac{3}{2}}\psi(-\tfrac{1}{4\tau})+\psi(\tau)&=\int_{0}^{i\infty}\frac{\theta(u)}{(\tau+u)^{\frac{3}{2}}}du.\end{split}$ (9.5)
Lemma 9.2.
For $\tau\in{\mathbb{H}}$, we have
$\frac{1+i}{16\pi}\int_{0}^{i\infty}\frac{\theta(u)}{(\tau+u)^{\frac{3}{2}}}du=\sqrt{\frac{\tau}{8i}}\int_{\mathbb{R}}{\mathbf{e}}(\xi^{2}\tau)\frac{1+{\mathbf{e}}(2\xi\tau)}{1-{\mathbf{e}}(2\xi\tau)}\xi d\xi.$
Proof.
We start with the left hand side. Since $\theta(u)=1+2\sum_{n=1}^{\infty}{\mathbf{e}}(n^{2}u)$, we have
$\int_{0}^{i\infty}(\tau+u)^{-\frac{3}{2}}du=2\tau^{-\frac{1}{2}},\quad\int_{0}^{i\infty}\frac{2e^{2\pi in^{2}u}}{(\tau+u)^{\frac{3}{2}}}du=2n\sqrt{2\pi}{\mathbf{e}}(-\tfrac{1}{8}){\mathbf{e}}(-n^{2}\tau)\Gamma(-\tfrac{1}{2},-2\pi in^{2}\tau).$
For the right side, note that the integrand is an even function of $\xi$, so we only need to deal with the part $\xi>0$. Since $\tau\in{\mathbb{H}}$, in this range we have $|{\mathbf{e}}(2\xi\tau)|<1$ and
$\frac{1+{\mathbf{e}}(2\xi\tau)}{1-{\mathbf{e}}(2\xi\tau)}=1+2\sum_{n=1}^{\infty}{\mathbf{e}}(2n\xi\tau).$
Hence we get
$\int_{{\mathbb{R}}}{\mathbf{e}}(\xi^{2}\tau)\xi d\xi=2\int_{0}^{\infty}{\mathbf{e}}(\xi^{2}\tau)\xi d\xi=\frac{i}{2\pi\tau}$
and
$4\int_{0}^{\infty}e^{2\pi i\xi^{2}\tau+4\pi in\xi\tau}\xi d\xi=ne^{-2\pi in^{2}\tau}(-2\pi i\tau)^{-\frac{1}{2}}\Gamma(-\tfrac{1}{2},-2\pi in^{2}\tau).$
The lemma is proved by comparing the two expressions term by term. ∎
Combining (9.3), (9.4) and Lemma 9.2, we conclude that for $\tau\in{\mathbb{H}}$,
$(-2i\tau)^{-\frac{3}{2}}\mathcal{H}(-\tfrac{1}{4\tau})+\mathcal{H}(\tau)=-\frac{1}{24}\theta(\tau)^{3}.$
If we define
$\begin{split}\mathcal{H}^{*}(\tau)\vcentcolon=\mathcal{H}(\tfrac{\tau}{2})+\frac{1}{48}\Theta(\tau)^{3}=&-\frac{1}{16}+\sum_{n=1}^{\infty}\left(H(n)+\frac{r_{3}(n)}{48}\right)e^{\pi in\tau}\\ &+\frac{1}{8\pi}\left(\sqrt{\frac{2}{y}}+2\sqrt{\pi}\sum_{m=1}^{\infty}m\Gamma(-\tfrac{1}{2},2\pi m^{2}y)e^{-\pi im^{2}\tau}\right),\end{split}$ (9.6)
then $\mathcal{H}^{*}$ satisfies
$(-i\tau)^{-\frac{3}{2}}\mathcal{H}^{*}(-\tfrac{1}{\tau})+\mathcal{H^{*}}(\tau)=0.$ (9.7)
Therefore, we define
$\begin{split}G_{\frac{3}{2}}(\tau;r)&\vcentcolon=F_{\frac{3}{2}}(\tau;r,\tfrac{3}{4})-\frac{8\sin(\pi r^{2})}{r\sinh(\pi r)}\mathcal{H}^{*}(\tau)\\ &=\frac{\sin(\pi r^{2})}{2r\sinh(\pi r)}-\frac{\sin(\pi r^{2})}{r\sinh(\pi r)}\sum_{n=1}^{\infty}e^{\pi in\tau}\left(8H(n)+\frac{r_{3}(n)}{6}\right)\\ &+\pi{\mathbf{e}}(\tfrac{1}{8})\sum_{n=1}^{\infty}e^{\pi in\tau}\left(\frac{n}{r^{2}}\right)^{\frac{1}{4}}\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c}J_{\frac{1}{2}}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\end{split}$ (9.8)
and
$\begin{split}\widetilde{G}_{\frac{3}{2}}(\tau;r)&\vcentcolon=\widetilde{F}_{\frac{3}{2}}(\tau;r,\tfrac{3}{4})-\frac{8\sin(\pi r^{2})}{r\sinh(\pi r)}\mathcal{H}^{*}(\tau)\\ &=\frac{\sin(\pi r^{2})}{2r\sinh(\pi r)}-\frac{\sin(\pi r^{2})}{r\sinh(\pi r)}\sum_{n=1}^{\infty}e^{\pi in\tau}\left(8H(n)+\frac{r_{3}(n)}{6}\right)\\ &+\pi{\mathbf{e}}(-\tfrac{3}{8})\sum_{n=1}^{\infty}e^{\pi in\tau}\left(\frac{n}{r^{2}}\right)^{\frac{1}{4}}\sum_{2\nmid d>0}\frac{\widetilde{K}(r^{2},n,d,\nu_{\Theta}^{3})}{d}J_{\frac{1}{2}}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{d}\right).\end{split}$ (9.9)
These functions satisfy the functional equation
$G_{\frac{3}{2}}(\tau;r)+(-i\tau)^{-\frac{3}{2}}\widetilde{G}_{\frac{3}{2}}(-\tfrac{1}{\tau};r)=e^{\pi ir^{2}\tau}.$ (9.10)
Here we write the linear combination $8H(n)+\frac{r_{3}(n)}{6}$ because it is always an integer.
Proof of Theorem 1.3.
The functions $b_{3,n}(r)$ and $\widetilde{b}_{3,n}(r)$ can be read from (9.8) and (9.9), respectively, with the help of [DLMF, (10.16.1)]:
$J_{\frac{1}{2}}(z)=\left(\frac{2}{\pi z}\right)^{\frac{1}{2}}\sin(z).$
It remains to compute the values
$B_{3,n}(0)=\lim_{r\rightarrow 0^{+}}B_{3,n}(r)\quad\text{and}\quad\widetilde{B}_{3,n}(0)=\lim_{r\rightarrow 0^{+}}\widetilde{B}_{3,n}(r)\quad\text{for }n\geq 1$ (9.11)
because by Proposition 8.4 and the dominated convergence theorem,
$G_{\frac{3}{2}}(\tau;0)$ $=\lim_{r\rightarrow 0}G_{\frac{3}{2}}(\tau;r)=\sum_{n=0}^{\infty}e^{\pi in\tau}\lim_{r\rightarrow 0}b_{3,n}(r),$
$\widetilde{G}_{\frac{3}{2}}(\tau;0)$ $=\lim_{r\rightarrow 0}{G}_{\frac{3}{2}}(\tau;r)=\sum_{n=0}^{\infty}e^{\pi in\tau}\lim_{r\rightarrow 0}\widetilde{b}_{3,n}(r)$
Similarly to (6.9), by $\sin(z)=z-z^{3}/3!+\cdots$, we claim that
[…]
This can be directly proved by applying Proposition 8.3 and Proposition 8.4, as well as (8.5) and Proposition 7.8 in order to show
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c>x}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c^{3/2}}\\ \displaystyle\sum_{2|c>x}\frac{K(0,n,c,\nu_{\Theta}^{3})}{c^{3/2}}\end{array}\right\}\ll x^{-\frac{1}{4}+\varepsilon}n^{\frac{1}{4}+\varepsilon}\quad\text{for }x>\max(1,r^{2},2\pi|r|\sqrt{n})$
which implies
$\lim_{r\rightarrow 0}\sum_{2|c}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c^{3/2}}=\sum_{2|c}\frac{S(0,n,c,\nu_{\Theta}^{3})}{c^{3/2}}$
and similarly for the case $2\nmid\widetilde{c}$.
We can evaluate the sums of Kloosterman sums in (9.12) by recalling the calculations in §7.2. Comparing with the quantity $B_{k}(n)$ defined by Bateman (taking $s=3$) [Bat51, (1.02)], one can find that
$B_{2^{\nu}}(n)=2^{-\frac{3}{2}\nu}{\mathbf{e}}(\tfrac{3}{8})\overline{\alpha_{3}(2^{\nu+1},n)},\quad B_{p^{\nu}}(n)=p^{-\frac{3}{2}\nu}\overline{\alpha_{3}(p^{\nu},n)},\ p>2.$
Moreover, for $n\geq 1$ and $4^{\alpha}\|n$ for $\alpha=\left\lfloor\nu_{2}(n)/2\right\rfloor$, the quantity $\chi_{2}(n)$ defined and calculated at [Bat51, under (4.06), also Theorem B] gives
$1-\chi_{2}(n)=2^{\frac{3}{2}}{\mathbf{e}}(\tfrac{1}{8})A_{3}(2,n,\tfrac{3}{4})=\left\{\begin{array}[]{ll}1,&\text{if }n/4^{\alpha}\equiv 7\ (\mathrm{mod}\ 8),\\ 1-2^{-\alpha},&\text{if }n/4^{\alpha}\equiv 3\ (\mathrm{mod}\ 8),\\ 1-3\times 2^{-\alpha-1},&\text{if }n/4^{\alpha}\equiv 1,2,5,6\ (\mathrm{mod}\ 8).\end{array}\right$. (9.13)
Additionally, [Bat51, Theorem B] gives that
$r_{3}(n)$ $=2\pi\sqrt{n}\chi_{2}(n)\cdot\frac{8K(-4n)}{\pi^{2}}\prod_{p^{2}|n}\left(\sum_{j=1}^{\left\lfloor\frac{\nu_{p}(n)}{2}\right\rfloor-1}\frac{1}{p^{j}}+\frac{p^{-\left\lfloor\frac{\nu_{p}(n)}{2}\right\rfloor}}{1-\left(\frac{-n/p^{2\left\lfloor\nu_{p}(n)/2\right\rfloor}}{p}\right)\frac{1}{p}}\right)$ (9.14)
$=2\pi\sqrt{n}\chi_{2}(n)\prod_{p>2}A_{3}(p,n,\tfrac{3}{4}),$ (9.15)
where $K(-4n)=\sum_{m=1}^{\infty}(\frac{-4n}{m})\frac{1}{m}$ is convergent.
Therefore, by the discussion above, as well as (7.14)-(7.17), for $n\geq 1$, if $n\neq 4^{\alpha}(8\beta+7)$ for any $\alpha,\beta\in{\mathbb{Z}}$, we have
$B_{3,n}(0)=r_{3}(n)\cdot\frac{1-\chi_{2}(n)}{\chi_{2}(n)}\quad\text{and}\quad\widetilde{B}_{3,n}(0)=\frac{r_{3}(n)}{\chi_{2}(n)}=B_{3,n}(0)+r_{3}(n);$ (9.16)
if $n=4^{\alpha}(8\beta+7)$ for some $\alpha,\beta\in{\mathbb{Z}}$, we have
$\begin{split}B_{3,n}(0)&=\frac{16}{\pi}\sqrt{n}\left(1-\chi_{2}(n)\right)K(-4n)\prod_{p^{2}|n}\left(\sum_{j=1}^{\left\lfloor\frac{\nu_{p}(n)}{2}\right\rfloor-1}\frac{1}{p^{j}}+\frac{p^{-\left\lfloor\frac{\nu_{p}(n)}{2}\right\rfloor}}{1-\left(\frac{-n/p^{2\left\lfloor\nu_{p}(n)/2\right\rfloor}}{p}\right)\frac{1}{p}}\right)\\ \widetilde{B}_{3,n}(0)&=B_{3,n}(0)/\left(1-\chi_{2}(n)\right)=B_{3,n}(0).\end{split}$ (9.17)
The formulas in the theorem about $B_{3,n}(0)$, $\widetilde{B}_{3,n}(0)$, $b_{3,n}(0)$, and $\widetilde{b}_{3,n}(0)$ can then be obtained from the relations between $r_{3}(n)$ and the Hurwitz class number $H(n)$ (see e.g. [Gro85, §4.8]).
This finishes the proof of Theorem 1.3. ∎
10. Concluding remarks
One of the corollaries of Theorem 1.2 and Theorem 1.3 is that the functions $B_{d,n}(r)$ satisfy the following:
$\begin{array}[]{l}|B_{4,n}(r)|,|\widetilde{B}_{4,n}(r)|\ll_{\varepsilon,c}n^{3/4+\varepsilon}\\ |B_{3,n}(r)|,|\widetilde{B}_{3,n}(r)|\ll_{\varepsilon,c}n^{1/2+\varepsilon}\end{array}\quad\text{if }r^{2}\geq c>0,\quad\varepsilon>0.$
In particular, these bounds give improved control over the coefficients $a_{d,n}(r)$, $\widetilde{a}_{d,n}(r)$ of the radial Fourier interpolation formulas (1.1) for $d=3,4$. Optimistically, we conjecture the following stronger bounds to hold.
Conjecture 10.1.
The functions $B_{d,n}(r)$ from Theorem 1.3 and Theorem 1.2 satisfy the following bounds:
$\begin{array}[]{l}|B_{4,n}(r)|,|\widetilde{B}_{4,n}(r)|\ll_{\varepsilon,c}n^{1/2+\varepsilon}\\ |B_{3,n}(r)|,|\widetilde{B}_{3,n}(r)|\ll_{\varepsilon,c}n^{1/4+\varepsilon}\end{array}\quad\text{if }r^{2}\geq c>0,\quad\varepsilon>0.$
The (hopeful) expectation is that, at least for fixed $r>0$, the coefficients $B_{d,n}(r)$ and $\widetilde{B}_{d,n}(r)$ should grow like coefficients of a weight $d/2$ cusp form, so the above conjecture corresponds roughly to the bound from the Ramanujan-Petersson conjecture. They are consistent with the numerics as far as we can check, and they can also be deduced assuming the Linnik-Selberg conjecture about sums of Kloosterman sums (see [ST09, (7)] for $d=4$, $k=2$ and [Sun25, Conjecture 1.4] for $d=3$, $k=\frac{3}{2}$).
We plan to generalize Theorem 1.2 and Theorem 1.3 to dimensions $d=1,2$ in a future work. In particular, we expect to recover the interpolation formula for even Schwartz functions in dimension 1 [RV19, Theorem 1], obtaining an explicit formula for the interpolation basis functions $a_{n}(x)$ in terms of real-variable Kloosterman sums.